### v2

# Giovanni Compiani Ilya Morozov Stephan Seiler

Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Ayush Kanodia, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments. We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance. Contact information: Giovanni Compiani (gio1compiani@gmail.com), Ilya Morozov (ilya.morozov@kellogg.northwestern.edu), Stephan Seiler (stephan.a.seiler@gmail.com). Giovanni Compiani gratefully acknowledges support from the Center for Applied AI at Booth.
Giovanni Compiani Ilya Morozov Stephan Seiler
University of Chicago Northwestern University Imperial College London
& CEPR
We propose a demand estimation method that leverages unstructured text and image data to infer substitution patterns. Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a random coefficients logit model. This approach enables researchers to estimate demand even when they lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design or functional benefits. Using data from a choice experiment, we show that our approach outperforms standard attribute-based models in counterfactual predictions of consumers’ second choices. We also apply it across 40 product categories on Amazon.com and consistently find that text and image data help identify close substitutes within each category.
Keywords: Demand Estimation, Unstructured Data, Deep Learning
1 Introduction
Many problems in economics and marketing—such as merger analysis, tariff evaluation, and optimal pricing—require researchers to estimate demand for differentiated products. A standard approach to these problems has been to estimate demand models that capture substitution through the similarity of product attributes. While common, this approach faces two practical challenges. First, researchers rarely observe all attributes that differentiate products. Instead, they rely on third-party data, where attributes are chosen based on unknown criteria, or gather their own data, subjectively selecting which attributes to collect. The collected attributes may not align with those most relevant to consumer choices. Second, consumers often consider visual design and functional benefits of products—dimensions that are difficult to capture through observed attributes. [^1] This criticism of attribute-based models has a long history in the economics literature. For example, Hausman (1994, p.229) skeptically remarks that applying such models to French champagne choices would require researchers to somehow quantify the bubble content.
In this paper, we show how researchers can incorporate text and image data in demand estimation to recover substitution patterns. Our approach uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging. Using pre-trained deep learning models, we extract low-dimensional features from images and texts and include them in a standard logit demand model, interacting them with random coefficients to capture substitution patterns. This approach enables researchers to incorporate hard-to-quantify product attributes, such as visual design from images and functional benefits from text, while also circumventing the need to subjectively choose which observed attributes to collect.
To validate our approach, we run an experiment that elicits consumers’ first and second choices. We estimate our model on first choices and show that it outperforms standard attribute-based demand models at counterfactually predicting second choices—a key empirical measure of substitution. Our findings suggest that researchers can use text and images instead of traditional attribute data and that these unstructured data may even capture substitution beyond what is reflected in observed product attributes. Lastly, we apply our approach to e-commerce data and consistently find that text and images effectively identify close substitutes across many product categories.
Proposed Approach and Validation
To extract information from unstructured data, we use pre-trained deep learning models, transforming images and texts into vector representations—embeddings. We then apply principal component analysis (PCA) to these embeddings to further reduce dimensionality and capture the main dimensions of product differentiation. This process generates separate embeddings and principal components for each data source: product images, titles, descriptions, and customer reviews. We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation. To identify the principal components that best capture substitution in the data, we select the specification that minimizes the Akaike Information Criterion (AIC).
Just as standard demand models capture substitution through the similarity of observed attributes, our approach captures it by looking at how close products are in the space of texts and images. This proximity will reflect substitution if consumers consider how a product looks and how it is described by sellers and other consumers. Alternatively, visual and textual descriptions may serve as proxies for product features and functional benefits. Take tablets, for example: product titles may reveal brand and screen size, descriptions may highlight whether the tablet is suitable for drawing or gaming, reviews may mention that a tablet is durable and child-friendly, and photos may showcase design features like color and casing style. By extracting embeddings from texts and images, we identify these dimensions of differentiation.
To evaluate our approach, we run an experiment where participants choose a book from a list of options. We randomize both the prices and rankings of books to generate the variation needed to identify substitution patterns (Berry and Haile, 2014). We elicit participants’ first and second choices, estimate each demand model using only first choices, and then counterfactually predict second choices. Because second choices reveal which books consumers substitute to when their first choice is unavailable, predicting them well requires a demand model that accurately captures substitution patterns.
We find that incorporating either text or images in demand estimation improves both in-sample fit and counterfactual predictions relative to the plain logit model without random coefficients. Images capture substitution because book covers visually convey genre and distinguish between fiction and non-fiction. Text performs even better, especially when we use more advanced models and richer text data. This is likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre. Crucially, our best-fitting specification—based on customer reviews—outperforms standard attribute-based demand models in counterfactual predictions. These findings suggest that text and images can replace standard attribute data and may even capture substitution patterns beyond what is reflected in observed attributes.
Beyond improving fit and counterfactual second-choice predictions, our approach accurately identifies the closest substitutes among products. We show that for some books our method correctly recovers the five closest substitutes, whereas alternative models misidentify them. Additionally, we simulate hypothetical mergers of book publishers and find that demand models diverge in their predictions of expected price increases. This divergence has significant implications for downstream decisions. For instance, if our model captures substitution accurately, an antitrust agency relying on a plain logit model or an attribute-based mixed logit model will incorrectly approve mergers that should be challenged and block mergers that should be approved.
Lastly, we apply our approach to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing. We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages. Across all categories, we consistently find that text and images contain useful information about substitution patterns. This additional information helps identify close substitutes within each category, leading to substantial deviations from restrictive substitution patterns in the plain logit model. These results confirm that our approach performs consistently well across a wide range of products, even in observational data that has less price variation than our experiment.
Contributions
Our main innovation is to show how researchers can incorporate unstructured text and image data in demand estimation to better recover substitution patterns. Standard demand estimation methods assume a pre-determined nesting structure (McFadden, 1974a), account for preference heterogeneity over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Berry et al., 2004), or parameterize utility covariances based on distances between products in attribute space (Bresnahan et al., 1997; Pinkse and Slade, 2004; Dotson et al., 2018; Chintagunta and Li, 2024). [^2] The idea of treating products as bundles of characteristics dates back to Lancaster (1966) and McFadden (1974b). Building on this body of work, we show how researchers can incorporate product images and textual product descriptions into standard demand models. [^3] In concurrent work, Sisodia et al. (2024) extract interpretable product attributes from images and use them as inputs to a conjoint analysis.
Given the prevalence of unstructured data, we see our approach as a valuable addition to the toolbox of empirical researchers. To make it easier for others to apply our approach, we provide a publicly available Python package, DeepLogit, alongside this paper. [^4] The package is available on PyPI and in the public GitHub repository: github.com/deep-logit-demand/deeplogit. Our package heavily borrows from the existing package Xlogit for GPU-accelerated estimation of mixed logit models, developed by Arteaga et al. (2022).
Extracting information from images can be especially valuable in markets where visual design and style matter, such as shoes, clothing, luxury watches, handbags, and art. Similarly, textual descriptions and reviews can capture hard-to-quantify attributes, such as network reliability and customer service for telecom services or usability, customizability, and developer support for software products. Even when product attributes are available, including information from texts and images may still improve demand estimates by capturing additional dimensions of product differentiation.
A natural consequence of using text and images is that our approach scales well across product categories. In contrast, capturing substitution through product attributes across many categories can be challenging. For example, Döpper et al. (2024) remark that doing so “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach circumvents the need to collect distinct attributes for each category, making it easier to estimate demand across many categories. This scalability may also be valuable for retailers and platforms seeking to estimate demand across diverse categories and optimize prices at scale.
We validate our approach using both experimental and observational data. Our experiment design lets us directly measure second-choice diversion ratios, allowing us to assess how well an estimated demand model predicts them counterfactually. Diversion ratios reflect substitutability and thus the intensity of price competition between two products. They are crucial for antitrust authorities trying to predict price increases from horizontal mergers (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2022), and for managers optimizing assortments and prices. We show that incorporating texts and images into demand models improves diversion ratio estimates, highlighting the value of our approach. Additionally, we use observational data to show that our method performs well across various product categories. We consider this extensive validation with both experimental and observational data an important contribution of the paper.
Several other papers have proposed leveraging text and image data in demand estimation. [^5] Netzer et al. (2012) use data on the co-occurrence of products mentioned in online discussion forums to generate a visual representation of competing products, but they do not incorporate these measures into demand estimation. Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities in a demand model, showing their approach produces reasonable price elasticity estimates. Similarly, Lee (2024) employs large-language models to predict utility intercepts of new products based on their textual descriptions. These methods use texts or images to predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution. [^6] For example, Lee (2024) inverts a plain logit model without random coefficients to extract product fixed effects for the prediction part of his analysis. Thus, his model restricts substitution patterns due to the IIA property. In contrast, our approach uses the same types of unstructured data to model utility covariances. This enables us to capture flexible substitution patterns—a key input for many empirical applications of interest.
A few papers propose estimating substitution patterns using survey data, a possible alternative to our use of unstructured data. Dotson et al. (2019) have survey participants rate each product image and incorporate rating correlations as utility correlation shifters in their demand model. Similarly, Magnolfi et al. (2022) ask survey participants to rank products relative to each other (e.g., “Product A is closer to B than to C”). They generate low-dimensional product embeddings from these rankings and use them as attributes in a random coefficients logit model. Like our paper, Magnolfi et al.’s approach provides a way to crowdsource product attributes when collecting such data is challenging. While our approach can complement these survey-based methods, it has the advantage of using only widely available text and image data, eliminating the need to conduct costly category-specific surveys. [^7] Dew (2024) also uses survey data to estimate consumer preferences for text and image embeddings. However, his focus is on predicting individual choices rather than estimating demand and modeling substitution.
Our approach is well-suited for several empirical applications. Researchers can use it to address a wide range of economics and marketing questions that require accurate and flexible estimates of product substitution. This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through changing assortments and prices. Additionally, marketers can apply this approach to help multi-product firms optimize prices and promotions (Hoch et al., 1995; Hitsch et al., 2021).
2 Proposed Approach
Our approach involves three steps: (1) extracting embeddings from images and texts, (2) reducing the dimensionality of these embeddings using PCA, and (3) including the resulting principal components in a random coefficients logit model. In this section, we focus on the first two steps, as the mixed logit model is standard and does not require a detailed explanation. Section 3.2 describes the exact model specification used in our empirical applications.
2.1 Extracting Embeddings from Texts and Images
We extract embeddings from texts and images using pre-trained deep-learning models. This enables us to leverage models trained on large, general-purpose datasets, which can efficiently capture differences in product images and descriptions. Further, using pre-trained models reduces the computational burden of estimation.
2.1.1 Image Embeddings
We employ deep learning models that were originally trained for object detection and classification tasks. Specifically, we use four pre-trained convolutional neural networks available in the Keras Deep Learning Library: VGG19, ResNet50, InceptionV3, Xception. [^8] VGG19, a deep convolutional neural network with 19 layers (Simonyan and Zisserman, 2015). ResNet50, is a 50-layer convolutional neural network (He et al., 2016). InceptionV3 and Xception are convolutional neural networks with 48 and 71 layers (Szegedy et al., 2016; Chollet, 2017). All four models achieve high predictive accuracy on the ImageNet validation dataset, and are widely known for their strong performance in image classification tasks. [^9] The ImageNet data (https://www.image-net.org/) is used for the ImageNet Large Scale Visual Recognition Challenge (ILSVRC), which evaluates algorithms for object detection and image classification, a benchmark in object category classification (Russakovsky et al., 2015).
We extract image embeddings from product photos, which researchers can easily obtain via online search or by scraping product pages from e-commerce platforms. Each model transforms the original image into a vector representation – an embedding – and classifies an image from its embedding by predicting the object it contains. These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”). However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings. Because these models perform well at distinguishing similar objects, we expect the embeddings to capture key visual features that differentiate products.
We do not commit to a single model and intentionally use multiple deep learning models with different architectures. As we explain in Section 3.2, we select the one that best explains substitution patterns in the data.
2.1.2 Text Embeddings
Next, we extract information from text data, including product titles, descriptions, and customer reviews. We use two simple but widely used “count” models as well as pre-trained models with varying levels of sophistication, generating a separate set of embeddings for each model and type of text data. Before extracting embeddings, we pre-process text data as described in Appendix A.
First, we use a bag-of-words count model, which transforms text into fixed-length vectors by counting word occurrences. This model does not consider word order and only measures how often words appear in a document. Second, we apply a bag-of-words model with a TF-IDF vectorizer. Although similar to the first method, TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words. This makes it more likely that text embeddings capture distinctive words that differentiate products from one another.
Next, we use the Universal Sentence Encoder (USE), a transformer-based language model (Cer et al., 2018), and the Sentence Transformer (ST) model, a pre-trained Sentence-BERT model designed to produce semantically meaningful sentence embeddings (Reimers and Gurevych, 2019). [^10] The model trained by Reimers and Gurevych (2019) is a more efficient version of the widely used BERT network (Devlin et al., 2018). Unlike the bag-of-words models, both USE and ST account for word order and the context in which words appear within sentences. Embeddings from these models are commonly used for text classification, semantic similarity analysis, clustering, and other natural language processing tasks. Both models achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017).
While count models are relatively simple, they can detect individual attributes mentioned in titles, descriptions, and reviews. In contrast, the more advanced ST and USE models are better at capturing nuanced similarities in how sellers and consumers describe products. For example, they can identify when sellers or consumers describe the same functional benefits using similar language, even if the words are not exactly the same. Thus, we view count models as a useful benchmark that tells us whether our method merely extracts attributes mentioned in texts or captures more subtle cues about product substitution.
2.2 Generating Principal Components
Embeddings provide a lower-dimensional representation of the original texts and images. However, they remain high-dimensional compared to the number of attributes researchers typically include in demand models. For example, it would be impractical to incorporate 512-dimensional VGG19 embeddings into a random coefficients logit model, as it would require computationally intensive numerical integration over all dimensions. For this reason, we further reduce the dimensionality by applying Principal Component Analysis (PCA) to the embeddings.
For a given category, we extract principal components from each set of embeddings corresponding to a given model and data type (e.g., USE embeddings from customer reviews) and incorporate them into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each component. Principal components are particularly appealing because they are orthogonal to each other, which mitigates multicollinearity and simplifies the estimation of random coefficients in the mixed logit model (Backus et al., 2021). We perform model selection by choosing the combination of principal components that minimizes the AIC (see Section 3.2.1).
Because we apply PCA separately within each category, the extracted principal components capture the key dimensions that differentiate products within that category. By contrast, raw embeddings come from models trained on large, general-purpose datasets. For instance, image embeddings come from models trained to classify images into broad categories, such as “tablet,” “laptop,” or “cellphone,” and thus contain features that distinguish images across these categories. However, to estimate demand, we need to analyze consumer choices within a given category, where products often share visual features due to similar designs. PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution patterns within each category.
2.3 Interpretation of Principal Components
Standard mixed logit models capture substitution by measuring how similar products are in terms of their observed attributes. Similarly, our approach captures substitution by measuring how similar products are in terms of their visual designs and textual descriptions.
There are several reasons why this proximity may be linked to substitution. Consumers may choose based on how a product looks and how it is described by sellers or other consumers. For instance, a consumer who finds one product visually appealing may also be drawn to similar-looking products, and they might expect products with similar descriptions to meet their needs in a similar way.
Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect key product features and benefits. Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly. Such texts may convey functional benefits that are hard to quantify. Similarly, photos may showcase design features like color, casing style, or the presence of a keyboard. By extracting principal components from texts and images, our approach identifies these dimensions of product differentiation and uses them to capture which products are close substitutes.
Given this interpretation, researchers should not treat the coefficients on the principal components as causal in the sense that changing product images or textual descriptions does not necessarily cause a change in substitution patterns. For example, if the principal components serve as proxies for choice-relevant product attributes, these attributes need not change if the seller modifies the product’s description or visual design. Hence, the principal components should be held fixed in counterfactual analyses. For this reason, our approach is not suited for studying optimal product design and positioning, or any other questions that require causal estimates of the effects of non-price characteristics on choices.
In some markets, texts and images might not convey information relevant to consumer choices, in which case the principal components would fail to predict substitution patterns. However, we consider this scenario unlikely, as sellers have strong incentives to highlight important product attributes in titles and descriptions, and online platforms may encourage them to do so. Similarly, when writing reviews, consumers often mention product aspects they think others will find useful. Consistent with this idea, we find that in all product categories we study, at least some image- or text-based principal components predict substitution patterns.
Similar to observed product attributes in standard mixed logit demand models, we treat principal components as fixed over time. This assumption is reasonable because principal components capture time-invariant product attributes, functional benefits, and visual design features. While sellers could theoretically modify images and textual descriptions to reposition their products, we find that in practice, product images, titles, and descriptions rarely change in our empirical context (see Appendix C for details).
2.4 Advantages of Our Approach
To understand how our approach relates to existing methods, we compare it to the standard approach commonly used in demand estimation, which relies on observed product attributes. These attributes are usually incorporated into a discrete choice demand model, such as a random coefficients logit model, where substitution patterns are captured through the interaction between product attributes and preference heterogeneity.
While common, this approach presents two practical challenges. First, researchers rarely observe all attributes that differentiate products. Instead, they often rely on third-party datasets, where data providers report a limited set of attributes selected based on unknown criteria. Others gather their own data using a combination of web scraping and human coders—–a costly and time-consuming process that typically limits researchers to collecting only easily quantifiable attributes. Since attributes are selected subjectively in both cases, they may not align with those most relevant to consumer choices. Our approach avoids collecting data on product attributes and instead extracts information about product substitution from text and image data.
The second challenge is that consumers often consider visual design and functional benefits—dimensions that are difficult to capture through observed attributes. For instance, in categories like clothing or home decor, product aesthetics can strongly influence choices. A shopper selecting a dress might care about overall style, design pattern, or neckline type—details visible in photos but difficult to quantify without training a custom classification model. Similarly, textual descriptions might reveal attributes that are hard to quantify. Customer reviews, for instance, might highlight that certain headphones are ideal for listening to audiobooks, revealing a benefit not captured by observed attributes. Our approach allows researchers to estimate substitution patterns along these otherwise unobserved dimensions.
A natural consequence of using text and images is that our approach scales more easily across categories than standard attribute-based models. For instance, Döpper et al. (2024) estimate demand estimation across a large set of categories and note that incorporating product attributes would be difficult because it would require them to decide which attributes are relevant in each category and collect data on those attributes. Our approach circumvents the need to collect category-specific attributes, making it easier to estimate demand across many categories. This scalability may also be valuable for online retailers like Amazon or Alibaba looking to optimize prices across hundreds of categories. In Section 4, we further illustrate this point by applying our approach across 40 Amazon categories.
3 Choice Experiment
We begin by applying our method to data from a choice experiment. In this experiment, we randomize product prices and positions, generating clean variation that helps estimate substitution patterns. We also collect data on consumers’ second choices, which are rarely available in observational datasets. These data allow us to assess how well our model recovers substitution patterns by testing its ability to predict second choices counterfactually relative to alternative models.
3.1 Experiment Design
We recruited participants from the online platform Prolific and asked each participant to complete two choice tasks via a Qualtrics survey.
In the first choice task, we asked each participant to choose one book from a set of ten alternatives (Figure 1). These ten books were chosen from Amazon’s bestseller lists (Figure 2). We chose them from three different genres—Mystery, Fantasy, and Self-Help—anticipating that genre would predict substitution patterns and could thus be incorporated in attribute-based demand models. We instructed each participant to choose the book they would purchase if faced with this selection in a real bookstore.
Figure 1 shows an example of a choice task, displaying the top portion of the page as it appeared to participants. For each book, we showed participants the cover image, title, author, publication year, genre, number of pages, description, and five customer reviews, all collected from Amazon product pages. We presented books in a random order to each participant and drew each book’s price from a discrete uniform distribution ranging from \$3 to \$7.
Figure 1: Example of a choice task in our experiment. The screenshot displays the top portion of the page as it appeared to participants.
After a participant selected a book from the list of ten options, we removed this book from the list. The participant then proceeded to a second choice task, where they chose from the remaining nine books. The positions and prices of the books remained the same as in the first choice task, with the only difference being that the participant’s first choice was no longer available.
Based on power calculations from a pilot experiment, we estimated that we needed around 10,000 participants to determine whether our model outperforms alternative models in predicting second choices. We pre-registered this target sample size. [^11] The pre-registration document for this study can be accessed at https://tinyurl.com/ekjp4pfp. In total, we recruited 10,775 participants between June 14 and June 27, 2024. Following the pre-registered sample selection criteria, we excluded around 14% participants who failed comprehension questions, did not complete the survey, or spent less than one minute on the entire study. After applying these criteria, our final sample consists of 9,265 participants.
Mystery Books Refer to caption “The Inmate” Refer to caption “Please Tell Me” Refer to caption “The Ritual” Refer to caption “The Housemaid”
Fantasy Books Refer to caption “The Ashes & The Star Cursed King” Refer to caption “Court of Ravens and Ruin” Refer to caption “The Serpent & The Wings of Night”
Self-Help Books Refer to caption “The Art of Letting Go” Refer to caption “Don’t Believe Everything You Think” Refer to caption “Dopamine Detox”
Figure 2: Ten books used in our experiment.
Because choice tasks in our experiment were hypothetical and not incentivized, it is important to verify that participants made meaningful selections. In Appendix B, we show that participants did not rush through the survey, responded to changes in book rankings and prices, and made choices consistent with their self-reported genre preferences.
3.2 Model Specifications
We compare our approach against two benchmark models: the plain logit model and a standard random coefficients model that leverages observed product attributes but uses no texts or images. All three models fall into the following framework, where the indirect utility of consumer $i$ from purchasing book $j$ is given by
$u_{ij}=\beta_{i}^{\prime}x_{j}+\theta_{i}^{\prime}PC_{j}-\gamma\cdot\text{rank}_{ij}-\alpha_{i}\cdot\text{price}_{ij}+\delta_{j}+\varepsilon_{ij}.$ (1)
Here $x_{j}$ is a vector of observed book attributes, $PC_{j}$ are principal components extracted from text and image embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in consumer $i$’s choice tasks, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d. taste shocks following a Type I Extreme Value distribution. We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books. This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does. When specifying observed attributes $x_{j}$, we include genre dummies for Mystery and Sci-Fi, publication year, and length in pages—all attributes participants observe in the choice tasks.
We estimate the following three models within this framework: [^12] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns related to product features, captured through observed characteristics or principal components from text and image data.
  1. 1.
    Mixed Logit with Principal Components. We include principal components $PC_{j}$, interacting them with random coefficients $\theta_{i}\sim N(0,\Sigma_{\theta})$, where $\Sigma_{\theta}$ is diagonal. Since the principal components do not vary for a given product over time, the mean of the random coefficients is absorbed by the product fixed effects $\delta_{j}$. In some specifications, we also estimate a random coefficient on price, assuming $\alpha_{i}\sim N(\bar{\alpha},\sigma_{\alpha})$. We do not include observed attributes, setting $\beta_{i}=0$. This model mimics a scenario where a researcher has access to unstructured data but does not observe attributes other than price.
  2. 2.
    Mixed Logit with Observed Attributes. We include observed attributes $x_{j}$, interacting them with random coefficients $\beta_{i}\sim N(0,\Sigma_{\beta})$, where $\Sigma_{\beta}$ is diagonal, and again, the mean is absorbed by product fixed effects. In some specifications, we also estimate a random coefficient on price, assuming $\alpha_{i}\sim N(\bar{\alpha},\sigma_{\alpha})$. We do not include principal components $PC_{j}$, setting $\theta_{i}=0$. This model reflects the standard practice of estimating substitution patterns through an interaction of observed attributes with consumer heterogeneity.
  3. 3.
    Plain Logit. We include neither attributes $x_{j}$ not principal components $PC_{j}$, estimating only the uniform price coefficient $\alpha_{i}=\alpha$, the rank coefficient $\gamma$ and product fixed effects $\delta_{j}$. This model provides a baseline where product substitution is solely driven by the average utilities of books and not their correlations.
Comparing plain logit to mixed logit with observed attributes tells us whether adding random coefficients on book attributes improves the estimation of substitution patterns. Similarly, comparing mixed logit with attributes to mixed logit with principal components tells us whether incorporating texts and images into demand estimation enables us to recover substitution patterns as well as, or better than, using product attributes. [^13] To ensure a fair comparison between models, for each model we start optimization from 100 different initial points covering a wide range of parameter values. This helps us maximize the probability of finding the global optimum in each model.
3.2.1 Model Selection
Within each of the three model classes listed in Section 3.2, we perform model selection based on $AIC$ [^14] An advantage of using AIC is that it can be interpreted as an estimator of the expected relative Kullback-Leibler (KL) Information based on the maximized log-likelihood function, corrected for asymptotic bias (Akaike, 1998; Burnham and Anderson, 2004). In the mixed logit model with observed attributes, we consider all possible combinations of random coefficients across four book attributes: price, publication year, length in pages, and genre. Genre is represented by two dummy variables, one each for Mystery and one for Self-Help, making Sci-Fi the baseline genre. We estimate all 16 specifications and select the one with the lowest $AIC$.
In the mixed logit model with principal components, we take the first three principal components from each set of embeddings and evaluate all possible combinations of random coefficients on the three components and price, selecting the specification with the lowest $AIC$ [^15] Increasing the number of principal components beyond three has only a marginal impact on our results and does not affect the in-sample fit or counterfactual performance of our selected model. We repeat this selection process for every text and image model and data type (e.g., USE embeddings from reviews). Lastly, among the selected specifications, we choose the one that minimizes $AIC$.
We do not combine principal components extracted from different embeddings (e.g, mixing texts with images) in our main analysis. As we discuss later in Section 3.4.4, we tested mixing data types in our experimental data but found no improvement in model fit.
3.3 Validation Approach
Prior research often evaluated demand models based on how intuitive the estimated cross-price elasticities look, given researchers’ prior knowledge of the market (Berry et al., 1995; Nevo, 2001). We follow a different approach by leveraging second choice data to compare our approach against alternatives. We estimate each model using only first choices and evaluate how well it predicts second choices. [^16] Conceptually, our validation approach reverses the logic of Berry et al. (2004), who show that second choice data are useful for estimating substitution patterns. By contrast, we omit second choice data from estimation and use them only for model validation.
Since the second-choice data are not used in estimation, our validation approach forces alternative models to confront a difficult counterfactual prediction problem. Second choices reveal which books consumers substitute to when their most preferred option becomes unavailable. Therefore, a model can only predict second choices well if it has accurately captured the true substitution patterns. Predicting second choices is also directly relevant in some applications. For instance, antitrust authorities can use them in merger analysis to calculate diversion ratios—a measure reflecting substitutability and thus the intensity of price competition between products (Conlon and Mortimer, 2021). We discuss this further in Section 3.6.
We evaluate each model as follows. First, we obtain MLE estimates by maximizing the total likelihood of participants’ first choices. Given these point estimates $\hat{\theta}$, we then compute the counterfactual second choice probabilities for each participant conditional on their first choice. Averaging these probabilities across participants, we construct the matrix $\mathbb{P}$, where each element $(j,k)$ represents the probability that a participant whose first choice is $j$ selects $k\neq j$ as their second choice. To assess each model, we compute the $RMSE$ based on element-wise differences between the matrix $\hat{\mathbb{P}}$ predicted by the model and $\mathbb{P}$ computed from empirical probabilities in the experimental data. A lower $RMSE$ indicates better model performance in predicting counterfactual second choices.
3.4 Validation Results
We compare the results of three models described in Section 3.2: (1) the plain logit model, (2) the mixed logit model with observed attributes, and (3) the mixed logit model with principal components. We evaluate these models based on the $RMSE$ of their counterfactual second-choice predictions. Figures 3 and 4 summarize the validation results, whereas Appendix Table A1 reports the $AIC$ and $RMSE$ values for all specifications considered within each model class.
3.4.1 Plain Logit
We first report the results from the plain logit model. As shown in Appendix Table A1, this model achieves a counterfactual $RMSE$ of $0.089$. Because this model assumes independence of irrelevant alternatives (IIA), it cannot capture which close substitutes participants switch to when their first choices become unavailable. As a result, the model predicts relatively uniform second-choice probabilities that do not match the second choice data.
3.4.2 Mixed Logit with Principal Components
Next, Figure 3 presents results from the mixed logit model with principal components. The first panel shows $RMSE$ from the plain logit model, indicated by a vertical dashed line, while other panels display results from our proposed approach, which incorporates principal components obtained from different types of unstructured data and machine learning models. Each marker reflects the second-choice $RMSE$ from the best configuration of random coefficients on the principal components and price chosen by our model selection algorithm (see Section 3.2.1).
Performance varies widely across specifications. While some models perform similarly to the plain logit, others deliver significantly better counterfactual predictions. These differences highlight that researchers should always perform model selection to find the specification that best captures substitution patterns from text and images. In practice, researchers can rely on $AIC$ as the main selection criterion. This works well in our experimental data, as $AIC$ selects the model that also achieves the lowest counterfactual $RMSE$ (see Appendix Table A1).
Figure 3: Counterfactual performance of mixed logit models with principal components in terms of predicting second choices.
Figure 4: Counterfactual performance of attribute-based mixed logit models in terms of predicting second choices.
Images
Figure 3 shows that image-based mixed logit specifications perform well. All models in this class outperform the plain logit model, and the lowest-$AIC$ model, InceptionV3, reduces $RMSE$ by 9.8% relative to the plain logit model (from $0.089$ to $0.080$).
To understand why images predict substitution, consider the book covers used in our study (see Figure 2). Within the same genre, book covers often share similar design elements. For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, feature natural objects like twisted vines and golden roses, and have similar fonts and title placements. Two of these books are by the same author and have nearly identical covers. Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds. Given these visual similarities, we expect image embeddings in our models to capture genre information, which correlates with substitution patterns.
However, we emphasize that image embeddings might be informative about substitution beyond simply revealing the book’s genre. For example, second-choice data shows that participants perceive mystery books as closer substitutes to fantasy books than self-help titles (Appendix Figure A3). Our image embeddings capture this pattern because mystery and fantasy book covers tend to be similar. Both use dark or muted tones with dramatic highlights (e.g., gold, silver, teal, or yellow) to create contrast, which is captured by our image embeddings.
Texts
The remaining panels of Figure 3 show the performance of text-based models. We find that counterfactual second-choice predictions improve when we either use more advanced text models or incorporate richer text data.
First, counterfactual performance improves substantially when we move from simple bag-of-words models to the more advanced USE and ST models. While bag-of-words models provide little to no improvement over the plain logit, both USE and ST consistently outperform the plain logit across all types of text data. This may be because textual descriptions contain detailed information about book plots but do not always use the exact same words or phrases. Therefore, extracting substitution patterns from text requires a natural language model that can accurately measure semantic similarity.
Second, counterfactual performance improves as we include more text data into the demand model. For instance, the USE model achieves a second-choice $RMSE$ of $0.080$ with book titles, $0.075$ with descriptions, and $0.069$ with reviews. This indicates that customer reviews provide more information about substitution patterns than descriptions, which, in turn, are more informative than titles.
These results make intuitive sense. Although book titles are brief, they often contain subtle genre cues. For example, mystery titles frequently hint at secrets (Please Tell Me) or characters in confined situations (Housemaid, Inmate), whereas fantasy titles often include words that evoke a gothic atmosphere (Serpent, Wings, Ravens, Ruin). Book descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals. Reviews go a step further: consumers summarize plots while also expressing opinions. When reviewing mystery novels, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings.
As with images, texts can provide information about substitution beyond simply revealing genres. For example, book descriptions and reviews make it clear that the self-help books Dopamine Detox and Don’t Believe Everything You Think both discuss how modern distractions and mental habits negatively affect well-being. In contrast, another self-help book, The Art of Letting Go, focuses on letting go of painful memories and emotions. These differences may explain why our text-based models correctly identify the first two books as much closer substitutes, something alternative models fail to detect (see Section 3.5 for details).
[…]
3.4.3 Mixed Logit with Observed Attributes
Figure 4 presents the results for the mixed logit model with observed attributes. We compare it against two benchmarks: the plain logit model and the lowest-$AIC$ text-based model identified earlier. Several specifications significantly outperform the plain logit model. The lowest-$AIC$ specification is a mixed logit model with random coefficients on pages, year, and genre. [^17] To be precise, this model achieves the same $AIC$ as two other mixed logit models (i.e., “Price, Pages & Year” and “Pages & Year”), but all three specifications generate identical second-choice predictions, and thus, achieve the same $RMSE$. By capturing similarities in observed attributes, this model breaks the IIA pattern and generates more accurate second-choice predictions. Consequently, it reduces $RMSE$ by 12.9% compared to plain logit (from $0.089$ to $0.077$).
3.4.4 Summary
Taken together, these findings illustrate the value of our approach, showcasing its ability to extract information about substitution patterns from text and images. We deliberately chose books for this study, thinking that observed attributes like genre would strongly predict substitution. Yet, we find that many text-based models achieve comparable $RMSE$ to that of attribute-based models. Moreover, the best-fitting text-based model reduces $RMSE$ by an additional 10% relative to the best attribute-based model. Thus, incorporating text improves counterfactual predictions beyond what is achievable with observed attributes.
One might wonder whether texts, images, and attributes provide distinct information about substitution patterns. If so, the best approach to estimating demand would be to incorporate all three data types. To explore this, we start with our selected review-based specification and examine whether model fit can be improved by adding (1) any combination of observed product attributes or (2) any combination of principal components from ResNet, the lowest-$AIC$ image model. In both cases, we find no fit improvements. This result suggests information across the three data types is highly correlated, with text providing the strongest signal of substitution.
Panel A. Predicted Second-Choice Probabilities when First Choice is Dopamine Detox (S)
Experimental Data Plain Logit Attribute-Based Mixed Logit Review-Based Mixed Logit
Book Prob. Book Prob. Book Prob. Book Prob.
Don’t Believe (S) 0.353 Please Tell Me (M) 0.182 Don’t Believe (S) 0.221 Don’t Believe (S) 0.259
Art of Letting Go (S) 0.249 The Inmate (M) 0.152 The Inmate (M) 0.153 Art of Letting Go (S) 0.176
Please Tell Me (M) 0.112 The Housemaid (M) 0.141 Please Tell Me (M) 0.152 Please Tell Me (M) 0.129
The Inmate (M) 0.094 Don’t Believe (S) 0.137 Art of Letting Go (S) 0.135 The Inmate (M) 0.125
The Housemaid (M) 0.057 Serpent & Wings (F) 0.095 The Housemaid (M) 0.132 The Housemaid (M) 0.100
Serpent & Wings (F) 0.042 Art of Letting Go (S) 0.094 Court of Ravens (F) 0.101 Court of Ravens (F) 0.074
Court of Ravens (F) 0.034 Court of Ravens (F) 0.080 Serpent & Wings (F) 0.057 The Ritual (M) 0.049
The Ritual (M) 0.031 Ashes & Star (F) 0.065 The Ritual (M) 0.028 Serpent & Wings (F) 0.044
Ashes & Star (F) 0.030 The Ritual (M) 0.054 Ashes & Star (F) 0.022 Ashes & Star (F) 0.043
Panel B. Predicted Second-Choice Probabilities when First Choice is Please Tell Me (M)
Experimental Data Plain Logit Attribute-Based Mixed Logit Review-Based Mixed Logit
Book Prob. Book Prob. Book Prob. Book Prob.
The Inmate (M) 0.325 Don’t Believe (S) 0.164 The Inmate (M) 0.163 The Housemaid (M) 0.170
The Housemaid (M) 0.250 Dopamine Detox (S) 0.149 The Housemaid (M) 0.149 The Inmate (M) 0.158
Don’t Believe (S) 0.088 The Inmate (M) 0.147 Don’t Believe (S) 0.145 Don’t Believe (S) 0.119
Art of Letting Go (S) 0.066 The Housemaid (M) 0.145 Dopamine Detox (S) 0.117 Serpent & Wings (F) 0.110
Serpent & Wings (F) 0.066 Art of Letting Go (S) 0.119 Art of Letting Go (S) 0.109 Art of Letting Go (S) 0.108
Dopamine Detox (S) 0.063 Serpent & Wings (F) 0.091 Court of Ravens (F) 0.099 Court of Ravens (F) 0.104
Court of Ravens (F) 0.058 Court of Ravens (F) 0.087 Serpent & Wings (F) 0.097 Dopamine Detox (S) 0.096
The Ritual (M) 0.056 The Ritual (M) 0.057 Ashes & Star (F) 0.070 Ashes & Star (F) 0.072
Ashes & Star (F) 0.029 Ashes & Star (F) 0.041 The Ritual (M) 0.050 The Ritual (M) 0.064
Panel C. Predicted Second-Choice Probabilities when First Choice is Ashes & Star (F)
Experimental Data Plain Logit Attribute-Based Mixed Logit Review-Based Mixed Logit
Book Prob. Book Prob. Book Prob. Book Prob.
Serpent & Wings (F) 0.275 The Housemaid (M) 0.155 Please Tell Me (M) 0.178 The Inmate (M) 0.184
Court of Ravens (F) 0.243 Don’t Believe (S) 0.155 The Inmate (M) 0.157 Please Tell Me (M) 0.166
The Inmate (M) 0.105 The Inmate (M) 0.153 The Housemaid (M) 0.134 The Housemaid (M) 0.123
The Ritual (M) 0.082 Please Tell Me (M) 0.139 Serpent & Wings (F) 0.116 Don’t Believe (S) 0.110
Please Tell Me (M) 0.080 Dopamine Detox (S) 0.116 The Ritual (M) 0.105 Court of Ravens (F) 0.108
The Housemaid (M) 0.070 Art of Letting Go (S) 0.100 Don’t Believe (S) 0.093 Dopamine Detox (S) 0.092
Don’t Believe (S) 0.052 Court of Ravens (F) 0.080 Dopamine Detox (S) 0.075 Serpent & Wings (F) 0.090
Art of Letting Go (S) 0.048 Serpent & Wings (F) 0.062 Art of Letting Go (S) 0.075 The Ritual (M) 0.065
Dopamine Detox (S) 0.045 The Ritual (M) 0.041 Court of Ravens (F) 0.066 Art of Letting Go (S) 0.063
Table 1: Predicted second-choice probabilities and their data counterparts. Letters in parentheses indicate book genres: F=Fantasy, M=Mystery, and S=Self-Help.
3.5 Implications for Product Substitution
So far, we have shown that our approach improves counterfactual predictions of second choices. One might wonder whether these $RMSE$ improvements are substantial and if they fundamentally change our understanding of product substitution in this market. To explore this further, in Table 1 we compare predicted second-choice probabilities with their counterparts observed in the experimental data.
Panel A examines substitution patterns for the self-help book Dopamine Detox. The data shows that the two other self-help books are, by far, its closest substitutes. The plain logit model misidentifies these substitutes, incorrectly predicting that people would switch to books Please Tell Me and The Inmate—two popular books with the largest market shares. The attribute-based mixed logit correctly identifies Don’t Believe Everything You Think as the closest substitute but mispredicts the second one, likely due to its over-reliance on estimated fixed effects. By contrast, our review-based model is the only one that correctly predicts all five closest substitutes in the correct order.
Panel B shows a similar example with the substitutes for the mystery book Please Tell Me. The plain logit model mispredicts second-choice probabilities, incorrectly suggesting that, in the absence of this book, participants would mainly switch to self-help books. By contrast, the attribute-based mixed logit model captures strong within-genre substitution, correctly identifying the top three closest substitutes. Our review-based model also correctly predicts the identities of these substitutes but swaps the first two. However, the review-based model recognizes that The Housemaid and The Inmate have significantly higher second-choice probabilities than the third-closest substitute—an insight the attribute-based model misses.
These examples illustrate that, beyond reducing $RMSE$ and predicting second-choice probabilities more accurately on average, our approach can learn which products are the closest substitutes.
Despite these favorable examples, our approach does not always accurately capture substitution. Panel C shows an example where the review-based model misidentifies the closest substitutes for the fantasy book The Ashes & The Star-Cursed King. In fact, all three models fail, incorrectly predicting mystery and self-help books as the closest substitutes. These deviations from observed second choices suggest that there is not enough variation in the first-choice data to reliably estimate substitution patterns for some alternatives.
3.6 Implications for Pricing: Merger Simulations
To illustrate how estimated substitution patterns can influence counterfactuals of interest, we conduct simulations of horizontal mergers—a natural application of our approach. Antitrust agencies routinely use demand models to assess whether a hypothetical merger would lead to a significant price increase (Federal Trade Commission, 2022). If the merging firms’ products are close substitutes, the merger creates strong upward pricing pressure, as the firm can “recapture” some consumers after raising prices. Therefore, predicting how a merged firm would set prices requires accurate estimates of substitution patterns.
Figure 5: Results of hypothetical merger simulations. For each simulated merger of Dopamine Detox with one other book, the figure shows the average price increase among the two merging books.
For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher, who sets prices by solving a joint first-order condition. In both cases, publishers take the prices of the other eight books as given, fixed at \$5.181818We fix the prices of the other books at \$5, the average value in the experimental dataset. We do not optimize prices of other books because the model does not include an outside option. Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
We recognize that limiting the choice set to ten books makes our analysis somewhat artificial, as major publishers typically manage vast assortments with hundreds of thousands of titles. However, our simulations offer a clear way to evaluate how estimated substitution patterns translate into pricing decisions. This application is also policy-relevant given the high-profile mergers among major book publishers in the past few decades, which have drawn attention and regulatory scrutiny. [^19] In 2022, a federal judge blocked the proposed merger between Penguin Random House and Simon & Schuster, citing concerns that it would stifle competition (U.S. Department of Justice, 2022). This merger would have reduced the “Big Five” publishers to the “Big Four” and significantly increased the concentration of the publishing market. Prior to that, Penguin Books and Random House merged in 2013 to form Penguin Random House—the largest trade book publisher globally.
Figure 5 illustrates the predicted price increase resulting from the joint ownership of Dopamine Detox with each of the other books. We select Dopamine Detox as an example where our approach outperforms alternative models in capturing substitution patterns (see Table 1). For each simulated merger, Figure 5 reports the average price increase across the two merging books.
As discussed in Section 3.5, the plain logit model fails to capture strong within-genre substitution. Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones. This misclassification affects the predicted price changes in merger simulations. Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think. These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 9%—approximately 2 to 2.5 times higher.
The attribute-based mixed logit model produces predictions more aligned with the review-based model. In most cases, both models deviate from the plain logit in the same direction, yielding similar or even identical predicted price increases. However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to our selected text-based model.
These divergent predictions could lead decision-makers to different conclusions. As an example, consider an antitrust agency that applies a heuristic rule, challenging all mergers expected to increase prices by more than 5% (Bhattacharya et al., 2023). Compared to decisions made using the review-based model, which best captures substitution patterns, a decision-maker relying on the plain logit model would approve two mergers that should be challenged (Don’t Believe Everything You Think, The Art of Letting Go) and challenge one that should be approved (The Housemaid). Similarly, a decision-maker using the attribute-based mixed logit model would approve a merger that should be challenged (The Art of Letting Go) and challenge one that should be approved (The Housemaid).
4 Application to Online Retail Data
Next, we apply our approach to choice data from several online markets. We pursue two goals with this application. First, we aim to assess the external validity of our experimental results. Although our estimation approach works for books, this does not mean it would capture substitution patterns well in other product categories. By applying this approach across multiple categories, we also show that it can be applied broadly without being tailored to specific markets. Second, by exploring which text or image data best predict substitution patterns in various categories, we can offer practical guidance to empirical researchers on what unstructured data to collect for demand estimation.
4.1 Data
Online Purchases.
We use purchase data from the 2019-2020 Comscore Web Behavior Panel, which includes a sample of around two million U.S. households. We use the dataset constructed by Greminger et al. (2023), who classify over 12 million unique products from Amazon.com—browsed or purchased by Comscore panelists—into narrowly defined categories. This dataset also matches purchases with daily product price histories obtained from the third-party database Keepa.com. [^20] We do not observe product rankings, so we do not include them in our demand models. We focus on Amazon.com due to the high volume of Amazon purchases in the Comscore dataset.
We apply our method to 40 product categories. To construct this sample, we first select the 15 most-purchased products in each category and drop all other products. We then retain only the categories where these 15 products collectively account for at least 2,000 purchases. This selection criterion ensures that we observe a sufficient number of purchases to estimate both product-fixed effects and substitution patterns. Within each category, we treat each purchase as a choice instance and assume that consumers’ choice sets consist of the 15 selected products. [^21] We omit products purchased fewer than 10 times and those with missing price data, which means some categories in our sample include fewer than 15 products.
The chosen product categories cover a diverse range of items such as clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation). The first column of Table 2 lists all included categories. On average, each category contains 662 purchases, with an average product price of \$43.
Figure 6: Example: image and text data collected from product detail pages.
Images and Texts.
We collect image and text data directly from product detail pages on Amazon.com. Figure 6 shows a sample product page. For image embeddings, we use the default product photo displayed at the top of the product detail page. For text embeddings, we gather information from several sections: product titles at the top, bullet points describing the product’s attributes (which we refer to as the “product description”), and the texts of the 100 most recent reviews for each product.
4.2 Estimation and Results
In each product category, we follow the same approach as in the experiment, extracting embeddings, reducing their dimensionality through PCA, and including them in the random coefficients logit model.
We do not observe product attributes in these categories. Collecting such data is challenging because each category has its own set of choice-relevant attributes. Further, we would need to subjectively decide which attributes to collect in each category—a challenge that originally motivated us to look at images and texts. For these reasons, we focus on comparing our estimates to those from the plain logit model. In other words, we focus on showing that texts and images contain information about product substitution and that our approach can reliably extract this information across many categories.
4.2.1 $AIC$ Improvements
Table 2 summarizes the estimation results, showing the selected model and data type for each category, along with the corresponding $AIC$ improvement relative to the plain logit model. Figure 7 further plots the distribution of $AIC$ improvements across all categories.
As a rule of thumb, we consider that a model has strong support compared to a simpler model if it lowers $AIC$ by at least $2.0$, and very strong support if it lowers $AIC$ by at least $5.0$. When a simpler model is nested within a more complex one with only one additional parameter, applying these $AIC$ thresholds is approximately equivalent to conducting a likelihood ratio test at the 5% and 1% significance levels. [^22] When nested models differ by one parameter, the reduction in $AIC$ is given by $\Delta AIC=2-\lambda_{LR}$, where $\lambda_{LR}$ is the likelihood ratio statistic. Therefore, our $\Delta AIC$ thresholds of $2.0$ and $5.0$ correspond to the likelihood ratio thresholds $4.0$ and $7.0$ (p-values $0.0455$ and $0.008$).
As Figure 7 shows, we achieve, on average, an AIC reduction of $28.0$, with improvements reaching as high as $121.8$ in some categories. These results confirm that our approach performs consistently well across a wide range of categories, which span both durable and perishable goods.
Figure 7: Fit improvements relative to plain logit across 40 product categories. For each category, we compute $\Delta AIC$ as the $AIC$ of the lowest-$AIC$ specification that includes texts or images minus that of the plain logit model.
Category Data Type Model Type $\Delta$ AIC $\Delta$ HHI
1. Clothing Active Descriptions ST -30.4 129.0%
2. Clothing Shirts Reviews USE -19.1 116.7%
3. Clothing Underwear Images Resnet50 -18.5 88.2%
4. Clothing Sleep Reviews ST -18.5 93.6%
5. Clothing Tops & Blouses Descriptions USE -18.5 94.6%
6. Electronics Cables Titles USE -18.4 240.0%
7. Electronics Accessories Reviews TFIDF -19.8 10.7%
8. Electronics Keyboards Images VGG19 -20.0 88.2%
9. Electronics Memory Cards Images VGG16 -109.7 103.5%
10. Electronics Tablets Images Resnet50 -121.8 120.9%
11. Electronics Monitors Reviews USE -31.6 197.3%
12. Electronics Headphones Descriptions COUNT -26.0 1.2%
13. Electronics Media Players Reviews TFIDF -66.9 42.7%
14. Groceries Water Descriptions COUNT -22.5 123.0%
15. Groceries Coffee Images VGG16 -17.2 45.6%
16. Groceries Tea Titles ST -20.0 145.2%
17. Groceries Chips Reviews ST -13.5 61.9%
18. Household Aromatherapy Titles COUNT -19.7 190.7%
19. Household Batteries Reviews TFIDF -20.1 4.8%
20. Household Trash Bags Titles ST -24.6 93.5%
21. Household Paper Towels Titles TFIDF -28.9 188.4%
22. Home Sheets & Pillowcases Descriptions USE -22.3 121.2%
23. Bedroom Beds Reviews USE -17.1 69.4%
24. Bedroom Mattresses Titles USE -17.4 66.3%
25. Kitchen Food Storage Images Resnet50 -27.7 133.2%
26. Office Folders Titles USE -29.7 199.7%
27. Office Paper Images Resnet50 -19.4 11.4%
28. Office Markers Images VGG19 -22.3 228.7%
29. Office Pens Titles ST -14.4 241.0%
30. Office Printer Supplies Titles TFIDF -23.2 247.7%
31. Pet Cat Food Titles ST -16.9 130.6%
32. Pet Cat Litter Titles ST -21.9 126.9%
33. Pet Cat Snacks Images Xception -19.5 142.3%
34. Pet Dog Food Titles COUNT -19.2 65.6%
35. Pet Dog Treats Reviews ST -22.9 163.9%
36. Game Consoles Nintendo Titles COUNT -13.1 118.3%
37. Video Games Nintendo Images Resnet50 -36.7 26.2%
38. Video Games PC Images VGG19 -18.6 43.9%
39. Video Games PS4 Titles TFIDF -32.1 159.6%
40. Video Games Xbox Reviews USE -40.5 120.6%
Table 2: Estimation results across 40 categories in Comscore data. For each category, the table shows the selected model and data type yielding the lowest $AIC$, the corresponding $AIC$ improvement relative to the plain logit model, and the relative change in the $HHI$ of the predicted diversion ratios relative to the plain logit model.
4.2.2 Implications for Substitution
Next, we examine how incorporating unstructured data into demand estimation affects our understanding of product substitution. While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how diversion ratios estimated by our approach deviate from the restrictive patterns imposed by the plain logit model.
The assumption of independent utilities in the plain logit model implies that diversion ratios depend only on market shares and not on product similarities. Consequently, this model often fails to identify close substitutes, leading to overly flat diversion ratios (Conlon et al., 2023). By contrast, if our approach better identifies close substitutes, it should produce more unequal diversion ratios. To quantify this, we compute diversion ratios using the definition from Section 3 and calculate the Herfindahl-Hirschman Index ($HHI$) of predicted diversion ratios. [^23] Specifically, we compute the $HHI$ across $(J-1)$ conditional second choice probabilities for each first choice and then average the resulting $HHI$ measure across all $J$ possible first choices. We expect our approach to yield a higher $HHI$ than the plain logit model.
The last column in Table 2 presents the results. [^24] In some categories, our estimates of the price coefficient $\alpha$ are positive, likely due to correlation between prices and unobserved demand shocks, even after accounting for product fixed effects. While instrumental variables could address this, we do not pursue this approach, as addressing price endogeneity across many product categories is orthogonal to the contribution of our paper. Instead, we focus on counterfactuals, such as the diversion ratios from removing a product, which remain well-defined even when the price coefficient is positive. Across categories, our approach more than doubles the $HHI$ of diversion ratios, with an average increase of 115%. Additionally, $HHI$ rises in all 40 product categories, with some increases approaching 250%. These findings confirm that our approach produces significantly more variable diversion ratios across the board.
4.2.3 Relevance of Different Data Types
Lastly, we analyze which types of unstructured data yield the largest improvement in fit in each category. It would be misleading to report only the best-performing model in each category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit. We therefore follow the statistical literature on model selection and report Akaike weights for each data type (Burnham and Anderson, 2004). Formally, we define $\Delta_{i}=AIC_{i}-AIC_{min}$ for each estimated specification $i$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category. We then compute the Akaike weight for a data type $d\in\{\text{images},\text{titles},\text{descriptions},\text{reviews}\}$ as follows:
$w_{d}=\frac{\sum_{\{r:d(r)=d\}}\exp(-\Delta_{r}/2)}{\sum_{i}\exp(-\Delta_{i}/2)}$ (2)
The numerator sums over specifications using data type $d$, while the denominator sums over all estimated specifications. We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data. [^25] More precisely, in large samples, $w_{d}$ reflects the probability that this class of models is, in fact, the best model for the data in the sense of Kullback-Leibler information (Burnham and Anderson, 2004, p.272).
Figure 8: Akaike Weights of different types of unstructured data across 40 product categories. These weights reflect the relative importance of data types for predicting substitution patterns in the data.
Figure 8 shows the estimated Akaike weights by category and data type. We find considerable variation across categories in terms of which types of data are most important for predicting substitution patterns.
Importantly, many of these results are difficult to predict ex ante. For example, while we might expect visual product features to be most relevant in clothing categories, where consumers care about visual design, only one of the four clothing categories (“Clothing Underwear”) has an image weight exceeding those of texts. By contrast, estimating demand for “Tops & Blouses” requires product descriptions, for “Shirts” relies on customer reviews, and for “Activewear” it works best with either descriptions or reviews. Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in categories of Nintendo and PC video games, images contain significantly more information about substitution than text.
These results highlight an important lesson: researchers may not be able to predict ex ante which data types will best capture substitution.
5 Discussion
5.1 A Practitioner’s Guide for Using Text and Images
In this section, we summarize our key results and offer guidelines for researchers on how to incorporate unstructured data into demand estimation.
The main takeaway from our analyses is that text and image data contain information about substitution patterns, meaning these unstructured data can be useful for demand estimation. At the same time, our cross-category analysis in Section 4 highlights that researchers may not be able to predict ex ante which data type will best capture substitution in a given product category. Therefore, we recommend that researchers collect various types of unstructured data and perform model selection as we did in Sections 3 and 4. Specifically, researchers can first identify the lowest-$AIC$ set of principal components for each combination of data type and embedding model, and then determine which of these selected specifications minimizes $AIC$ across all data-model combinations. This selection rule performed well in our experiment, where it identified the model with the best counterfactual second-choice predictions.
In our experimental data in Section 3, we found that a text-based model produces the best counterfactual predictions, and that extending this model with observed attributes or image data did not significantly improve model fit. Nevertheless, in other contexts, researchers may still want to check whether incorporating observed product attributes can capture additional dimensions of product differentiation. This can be done by searching for a set of observed attributes that, when added to the selected model, further reduce $AIC$. A similar search can be conducted over additional text- and image-based principal components to determine whether including them improves model fit.
To conduct inference based on the estimates from our approach, researchers must account for model selection to obtain valid confidence intervals. One potential solution is to perform sample-splitting: researchers can use one randomly selected subset of the data for model selection and the other for estimating the covariance matrix of the estimates. [^26] For a detailed discussion of sample-splitting, see Wasserman and Roeder (2009). Taylor and Tibshirani (2015) provide examples of how researchers can construct corrected p-values and confidence intervals for specific estimation problems.
We mostly abstracted away from the issue of price endogeneity. Endogeneity concerns do not arise in our experiment because we randomize prices, whereas in the analysis of Amazon purchases, we rely on product fixed effects to capture product-level unobservables (e.g., unobserved quality) that may correlate with prices. In practice, researchers are often concerned that prices may be correlated with unobserved demand shocks that vary across markets and over time. To address this, researchers can combine our approach with existing methods for handling price endogeneity based on instrumental variables.
5.2 Future Research Directions
While we are not the first to use text and image data in demand estimation, our paper proposes and validates a concrete method for leveraging such data to flexibly estimate substitution patterns – a critical input for answering many empirical questions of interest. We do not claim that ours is the only or the best way to use unstructured data for this purpose. In fact, several questions remain open.
First, other demand modeling approaches might be able to better take advantage of text and image data than the random coefficients logit model. In an earlier version of this paper, we tested a pairwise combinatorial logit model from Koppelman and Wen (2000), which generalizes the logit model by allowing each product pair to have its own correlation in utilities. We allowed these correlations to depend on the distance between products in the space of text and image embeddings. This model performed significantly worse than our current approach of incorporating principal components into the standard random coefficients logit model, which highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance. [^27] Researchers can also try relaxing the parametric assumptions of standard demand models. For example, with small number of products, one could incorporate principal components into a nonparametric demand model (Compiani, 2022).
Second, our approach can be easily extended to newer, more advanced ML models that might better extract substitution patterns from texts and images. The approach we propose in this paper is modular in the sense that researchers can easily try alternative embeddings and test whether they improve counterfactual second-choice predictions. To facilitate such tests, we make our experimental dataset and estimation code publicly available. [^28] Replication codes and experimental data are available in our public repository: github.com/ilyamorozov/DeepLogitReplication. Separately, the Python package for implementing our method is available on PyPI and at: github.com/deep-logit-demand/deeplogit.
Third, researchers could improve upon our approach by constructing “optimal” embeddings specifically designed to improve the counterfactual predictions of interest. [^29] For example, Lee (2024) shows that fine-tuning LLMs helps researchers to better capture brand intercepts. Our current method relies on text and image models that were trained on datasets unrelated to our specific application, meaning they may not capture the most relevant dimensions of product differentiation. This suggests potential for improvement: researchers could fine-tune existing deep learning models or train custom ones to optimize for a specific objective, such as improving counterfactual predictions. It would be interesting to assess how much these tailored demand models improve upon pre-trained models and whether they can do so without significantly increasing the computational burden.
Lastly, our approach could extend to other forms of unstructured data, such as music or videos, potentially opening doors to studying product differentiation and competition in markets where this would otherwise be challenging. One could also leverage tools from the explainable AI literature to construct interpretable embeddings, which would help tackle questions such as optimal product design. [^30] See Sisodia et al. (2024) for a detailed discussion of related ideas.
6 Conclusion
In this paper, we demonstrated how researchers can incorporate unstructured text and image data in demand estimation to better recover substitution patterns. Our proposed estimation approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard discrete choice model. Using experimental data, we showed that our model outperforms standard attribute-based models in counterfactual predictions of second choices. We further validated our method with e-commerce data across dozens of categories and found that text and image data consistently help identify close substitutes within each category.
Appendix A Text Processing Steps
Before applying our text models, listed in Section 2.1.2, we pre-process our text data as follows. Working with product titles is straightforward because each title is a short text that typically includes only 10-15 words. When assessing similarity based on product descriptions, we merge the text from all bullet points, and apply our models to the merged text. For customer reviews, we transform the text of each review into a separate vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
For both bag-of-words approaches, we further pre-process text data by removing stopwords and lemmatizing words. We remove stopwords using the standard dictionary of common English words in the NLTK package. We lemmatize words using the WordNet Lemmatizer from the same package, NLTK. Then, we convert each pre-processed text into a vector of word occurrences (weighted word occurrences for TF-IDF). The other two models, USE and ST, have a built-in text pre-processing step. We therefore apply these models directly to the unprocessed text data.
Appendix B Choice Survey: Sanity Checks
Because choices in our experiment were hypothetical and not incentivized, we want to verify that participants made meaningful choices. Several summary statistics suggest that participants took the choice tasks seriously. First, only 50 of the initially recruited participants (less than 1%) completed the entire study in less than one minute and thus had to be dropped from our sample. The remaining participants spent, on average, 7 minutes on the survey overall and 1.3 minutes on the choice tasks, indicating they took their time to make careful selections and did not mindlessly click through the survey (Figure A1).
Second, in the choice tasks, participants were disproportionately more likely to select books of the genre they reported to be their favorite in a questionnaire before the choice tasks (Figure A2), suggesting they considered the book attributes when making their choices.
Finally, participants’ choices were consistent across the two choice tasks (Figure A3). For example, over 60% of participants who selected a mystery book in the first task chose another mystery book in the second choice task. This observation suggests participants considered their genre preferences when making both choices.
Appendix C Changes in Text and Images Over Time
As discussed in Section 2.3, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product. To validate this assumption, we examine whether text and images change over time.
Since we lack data on these changes for 2019-2020, we construct a separate sample by repeatedly collecting unstructured data from Amazon’s product detail pages daily from January 23 to March 4, 2025. To keep data collection manageable, we do not gather customer reviews and select a subset of 11 out of 40 categories, ensuring they cover all of Amazon’s departments (e.g., “Clothing,” “Food,” “Electronics”) represented in the full dataset. [^31] Recall that we average over embeddings extracted from the 100 most recent reviews. Even though customer reviews accumulate over time as consumers continue to write them, the average embeddings extracted from these reviews may remain approximately constant if consumers consistently discuss the same attributes. While we do not have review data to verify this claim, future research should examine whether review embeddings are stable in online markets. The selected categories are: “Shirts,” “Coffee,” “Aromatherapy,” “Mattresses,” “Markers,” “Pet Litters,” “Nintendo Games,” “Tablets,” “Memory Cards,” “Monitors,” and “Earbuds.” In each category, we collect data for the products used in our estimation in Section 4 that were not discontinued, totaling 136 products.
We find that product images do not change over time. Titles change for only six products (4%), mostly by adding or removing specific attributes or functional benefits. Similarly, descriptions change for just 21 products (15%). Most changes do not affect which product features are revealed, but they alter which ones are immediately visible versus being revealed only after clicking on “See more product details.” Thus, we conclude that changes in text and images are not a significant concern for the products in our empirical application.
Figure A1: Time spent by participants on choice tasks in the experiment.
Figure A2: Selected genres and self-reported genre preferences in the experiment.
Figure A3: Genres of participants’ first and second choices in the experiment.
First Choices (Data) Second Choices (Counterf.)
Model $\log L$ AIC RMSE Rel. to Plain Logit
$\Delta$
Panel A. Benchmark Model
Plain Logit -20488.4 41006.7 0.089
Panel B. Mixed Logit Models with Principal Components
Images: InceptionV3 -20475.8 40985.6 0.080 -0.009 -9.8%
Images: ResNet50 -20479.9 40991.8 0.080 -0.009 -9.8%
Images: VGG19 -20481.2 40994.4 0.083 -0.006 -6.8%
Images: Xception -20479.1 40990.3 0.086 -0.003 -3.1%
Product Titles: Bag-of-Words Count -20479.3 40992.7 0.085 -0.004 -4.6%
Product Titles: Bag-of-Words TF-IDF -20479.3 40992.7 0.085 -0.004 -4.6%
Product Titles: Sentence Encoder (USE) -20479.1 40992.2 0.080 -0.009 -10.5%
Product Titles: Sentence Transformer (ST) -20477.2 40986.5 0.082 -0.006 -7.1%
Descriptions: Bag-of-Words Count -20480.1 40992.2 0.089 -$0$0.2%
Descriptions: Bag-of-Words TF-IDF -20480.2 40992.3 0.089 -$0$0.2%
Descriptions: Sentence Encoder (USE) -20476.2 40984.4 0.075 -0.014 -15.5%
Descriptions: Sentence Transformer (ST) -20475.8 40983.6 0.073 -0.016 -18.0%
Reviews: Bag-of-Words Count -20481.0 40995.9 0.084 -0.005 -5.2%
Reviews: Bag-of-Words TF-IDF -20480.9 40995.7 0.084 -0.005 -5.1%
Reviews: Sentence Encoder (USE) -20472.0 40975.9 0.069 -0.020 -22.3%
Reviews: Sentence Transformer (ST) -20473.3 40978.6 0.072 -0.017 -18.8%
Panel C. Attribute-Based Mixed Logit Models
Price -20485.9 41003.9 0.088 -0.001 -0.8%
Pages -20484.9 41001.7 0.083 -0.006 -6.8%
Year -20484.7 41001.4 0.087 -0.002 -2.0%
Genre -20483.3 41000.7 0.081 -0.008 -8.9%
Price & Pages -20478.9 40991.9 0.077 -0.012 -13.0%
Price & Year -20484.1 41002.2 0.088 -0.001 -1.4%
Price & Genre -20483.2 41002.5 0.080 -0.008 -9.4%
Pages & Year -20478.3 40990.7 0.077 -0.011 -12.9%
Pages & Genre -20481.6 40999.2 0.076 -0.013 -14.9%
Year & Genre -20483.3 41000.7 0.081 -0.008 -8.9%
Price, Pages, & Year -20478.3 40990.7 0.077 -0.011 -12.9%
Price, Pages, & Genre -20478.9 40991.9 0.077 -0.012 -13.0%
Price, Year, & Genre -20482.2 41002.4 0.079 -0.010 -11.2%
Pages, Year, & Genre -20478.3 40990.7 0.077 -0.011 -12.9%
Price, Pages, Year, & Genre (All Attr.) -20476.9 40993.9 0.076 -0.013 -14.4%
Table A1: Model validation results. The table shows in-sample fit on first-choice data and counterfactual performance on second-choice data for all specifications considered in Figures 3 and 4 (see Section 3.2 for detailed description of these models).

### v3

@@
 Giovanni Compiani Ilya Morozov Stephan Seiler
-Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Ayush Kanodia, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments.
+Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Jon McClure, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments.
 We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance.
@@
 & CEPR
-We propose a demand estimation method that leverages unstructured text and image data to infer substitution patterns.
-Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a random coefficients logit model.
-This approach enables researchers to estimate demand even when they lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design or functional benefits.
-Using data from a choice experiment, we show that our approach outperforms standard attribute-based models in counterfactual predictions of consumers’ second choices.
-We also apply it across 40 product categories on Amazon.com and consistently find that text and image data help identify close substitutes within each category.
+We propose a demand estimation method that leverages unstructured data to infer substitution patterns.
+Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a mixed logit demand model.
+This approach enables demand estimation even when researchers lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design.
+Using a choice experiment, we show this approach outperforms standard attribute-based models at counterfactual predictions of second choices.
+We also apply it to 40 product categories on Amazon.com and consistently find that unstructured data identifies close substitutes within each category.
 Keywords: Demand Estimation, Unstructured Data, Deep Learning
@@
 In this paper, we show how researchers can incorporate text and image data in demand estimation to recover substitution patterns.
-Our approach uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging.
-Using pre-trained deep learning models, we extract low-dimensional features from images and texts and include them in a standard logit demand model, interacting them with random coefficients to capture substitution patterns.
+We propose an approach that uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging.
+Using pre-trained deep learning models, we transform images and texts into vector representations—embeddings.
+We then apply Principal Component Analysis (PCA) to reduce dimensionality and capture the main dimensions of product differentiation.
+We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation.
 This approach enables researchers to incorporate hard-to-quantify product attributes, such as visual design from images and functional benefits from text, while also circumventing the need to subjectively choose which observed attributes to collect.
-To validate our approach, we run an experiment that elicits consumers’ first and second choices.
-We estimate our model on first choices and show that it outperforms standard attribute-based demand models at counterfactually predicting second choices—a key empirical measure of substitution.
-Our findings suggest that researchers can use text and images instead of traditional attribute data and that these unstructured data may even capture substitution beyond what is reflected in observed product attributes.
-Lastly, we apply our approach to e-commerce data and consistently find that text and images effectively identify close substitutes across many product categories.
-Proposed Approach and Validation
-To extract information from unstructured data, we use pre-trained deep learning models, transforming images and texts into vector representations—embeddings.
-We then apply principal component analysis (PCA) to these embeddings to further reduce dimensionality and capture the main dimensions of product differentiation.
-This process generates separate embeddings and principal components for each data source: product images, titles, descriptions, and customer reviews.
-We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation.
-To identify the principal components that best capture substitution in the data, we select the specification that minimizes the Akaike Information Criterion (AIC).
-Just as standard demand models capture substitution through the similarity of observed attributes, our approach captures it by looking at how close products are in the space of texts and images.
-This proximity will reflect substitution if consumers consider how a product looks and how it is described by sellers and other consumers.
-Alternatively, visual and textual descriptions may serve as proxies for product features and functional benefits.
-Take tablets, for example: product titles may reveal brand and screen size, descriptions may highlight whether the tablet is suitable for drawing or gaming, reviews may mention that a tablet is durable and child-friendly, and photos may showcase design features like color and casing style.
-By extracting embeddings from texts and images, we identify these dimensions of differentiation.
-To evaluate our approach, we run an experiment where participants choose a book from a list of options.
-We randomize both the prices and rankings of books to generate the variation needed to identify substitution patterns (Berry and Haile, 2014).
-We elicit participants’ first and second choices, estimate each demand model using only first choices, and then counterfactually predict second choices.
-Because second choices reveal which books consumers substitute to when their first choice is unavailable, predicting them well requires a demand model that accurately captures substitution patterns.
-We find that incorporating either text or images in demand estimation improves both in-sample fit and counterfactual predictions relative to the plain logit model without random coefficients.
-Images capture substitution because book covers visually convey genre and distinguish between fiction and non-fiction.
-Text performs even better, especially when we use more advanced models and richer text data.
-This is likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre.
-Crucially, our best-fitting specification—based on customer reviews—outperforms standard attribute-based demand models in counterfactual predictions.
-These findings suggest that text and images can replace standard attribute data and may even capture substitution patterns beyond what is reflected in observed attributes.
-Beyond improving fit and counterfactual second-choice predictions, our approach accurately identifies the closest substitutes among products.
-We show that for some books our method correctly recovers the five closest substitutes, whereas alternative models misidentify them.
-Additionally, we simulate hypothetical mergers of book publishers and find that demand models diverge in their predictions of expected price increases.
-This divergence has significant implications for downstream decisions.
-For instance, if our model captures substitution accurately, an antitrust agency relying on a plain logit model or an attribute-based mixed logit model will incorrectly approve mergers that should be challenged and block mergers that should be approved.
-Lastly, we apply our approach to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing.
-We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages.
-Across all categories, we consistently find that text and images contain useful information about substitution patterns.
-This additional information helps identify close substitutes within each category, leading to substantial deviations from restrictive substitution patterns in the plain logit model.
-These results confirm that our approach performs consistently well across a wide range of products, even in observational data that has less price variation than our experiment.
-Contributions
-Our main innovation is to show how researchers can incorporate unstructured text and image data in demand estimation to better recover substitution patterns.
-Standard demand estimation methods assume a pre-determined nesting structure (McFadden, 1974a), account for preference heterogeneity over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Berry et al., 2004), or parameterize utility covariances based on distances between products in attribute space (Bresnahan et al., 1997; Pinkse and Slade, 2004; Dotson et al., 2018; Chintagunta and Li, 2024).
-[^2] The idea of treating products as bundles of characteristics dates back to Lancaster (1966) and McFadden (1974b).
-Building on this body of work, we show how researchers can incorporate product images and textual product descriptions into standard demand models.
-[^3] In concurrent work, Sisodia et al.
-(2024) extract interpretable product attributes from images and use them as inputs to a conjoint analysis.
 Given the prevalence of unstructured data, we see our approach as a valuable addition to the toolbox of empirical researchers.
@@
 (2022).
-Extracting information from images can be especially valuable in markets where visual design and style matter, such as shoes, clothing, luxury watches, handbags, and art.
-Similarly, textual descriptions and reviews can capture hard-to-quantify attributes, such as network reliability and customer service for telecom services or usability, customizability, and developer support for software products.
-Even when product attributes are available, including information from texts and images may still improve demand estimates by capturing additional dimensions of product differentiation.
-A natural consequence of using text and images is that our approach scales well across product categories.
-In contrast, capturing substitution through product attributes across many categories can be challenging.
-For example, Döpper et al.
-(2024) remark that doing so “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach circumvents the need to collect distinct attributes for each category, making it easier to estimate demand across many categories.
-This scalability may also be valuable for retailers and platforms seeking to estimate demand across diverse categories and optimize prices at scale.
-We validate our approach using both experimental and observational data.
-Our experiment design lets us directly measure second-choice diversion ratios, allowing us to assess how well an estimated demand model predicts them counterfactually.
-Diversion ratios reflect substitutability and thus the intensity of price competition between two products.
-They are crucial for antitrust authorities trying to predict price increases from horizontal mergers (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2022), and for managers optimizing assortments and prices.
-We show that incorporating texts and images into demand models improves diversion ratio estimates, highlighting the value of our approach.
-Additionally, we use observational data to show that our method performs well across various product categories.
-We consider this extensive validation with both experimental and observational data an important contribution of the paper.
+To show that the proposed method performs well at recovering substitution patterns, we validate it extensively on several datasets.
+Validation in demand estimation is challenging because substitution patterns are usually unobserved, leaving no clear “ground truth” to evaluate model predictions.
+To overcome this challenge, we design a choice experiment where participants choose a book from a list of options.
+We randomize both prices and rankings to identify substitution patterns.
+Crucially, we elicit both first and second choices, estimate different demand models using only first choices, and evaluate their performance based on how well they predict second choices counterfactually.
+[^3] We refer to second choice predictions as “counterfactual predictions” in the sense that second choices are not used during estimation, which is based solely on first choices.
+Conditional second choices are informative about substitution because they reveal which product a consumer would substitute to if their preferred option was unavailable.
+Recording second choices thus enables us to assess how well each demand model predicts substitution patterns.
+[^4] Second choices are related to the notion of diversion ratios, which reflect the degree of substitutability and thus the intensity of price competition between two products (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2023).
+Recovering them is crucial for antitrust authorities trying to predict price increases from horizontal mergers and for managers optimizing assortments and prices.
+Using these experimental data, we show that both text and images contain useful information about substitution patterns.
+Intuitively, images capture substitution because book covers visually convey information about genre.
+Text performs even better, likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre.
+We then show that the best performing model—based on customer reviews—improves counterfactual predictions of second-choices relative to the standard mixed logit with observed attributes.
+In fact, the improvement relative to that model is on par with the improvement that the mixed logit achieves over the plain logit model without random coefficients.
+Thus, in this application, we can recover substitution from unstructured data alone, and in fact can achieve better performance than is possible even with attribute data.
+To assess the economic importance of these improvements, we simulate hypothetical mergers of book publishers.
+We find that differences in estimated substitution patterns are large enough such that an antitrust agency relying on the standard mixed logit with attributes might incorrectly approve mergers that should be challenged and block mergers that should be approved.
+Thus, our results show that incorporating text and image data yields improvements in estimated substitution patterns that are both statistically and economically significant.
+We also validate our approach on observational data.
+We apply it to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing.
+We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages.
+Here, our contribution is twofold.
+First, we show that text and images contain useful signals about substitution in all studied categories, highlighting that our approach is broadly applicable across markets.
+Second, we find—perhaps unexpectedly—that text performs best in some categories where images might be expected to be more useful, such as clothing, and vice versa.
+A practical takeaway is that researchers may benefit from collecting both text and images, even when only one seems relevant a priori.
+Our paper contributes to the extensive literature on demand estimation.
+Specifically, we build on papers that model substitution through the heterogeneity of preferences over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Nevo, 2001; Berry et al., 2004).
+We show that incorporating the principal components of text and image embeddings into these models captures substitution patterns well.
+A key contribution of our paper is that we validate this approach extensively using both experimental and observational data.
+This paper is also related to a vast literature in computer science that transforms unstructured data into lower-dimensional representations (Krizhevsky et al., 2012; Mikolov et al., 2013; Goodfellow et al., 2016).
+There is a growing literature on applications of pre-trained deep learning models in economics and marketing (see Battaglia et al., 2024 for a review).
+While most prior work focuses on predicting observed outcomes, our goal is fundamentally different: we aim to predict counterfactual quantities, such as substitution patterns.
+Predicting these quantities well is crucial for applications like merger simulations and optimal pricing.
+Yet, they are usually unobserved in choice data.
+As a result, we cannot apply standard cross-validation techniques and instead need to design a custom experiment that allows us to measure these otherwise unobserved quantities.
+To our knowledge, we are the first to use such an experiment for validation of demand models.
+[^5] Berry et al.
+(2004) show that second choice data are useful for estimating substitution patterns, and Conlon et al.
+(2022) estimate demand from aggregate data on first-choice probabilities and a subset of second-choice probabilities.
+Conceptually, our validation approach reverses the logic of these papers: we only use first-choice data for estimation and treat second choices as “holdout” data for model validation.
+Additionally, Raval et al.
+(2022) use hospital closures induced by natural disasters to compare the ability of various models to predict diversions.
 Several other papers have proposed leveraging text and image data in demand estimation.
@@
 (2012) use data on the co-occurrence of products mentioned in online discussion forums to generate a visual representation of competing products, but they do not incorporate these measures into demand estimation.
-Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities in a demand model, showing their approach produces reasonable price elasticity estimates.
+Dew (2024) uses image embeddings from pre-trained models to predict consumers’ ratings for products based on past ratings elicited in a survey.
+Sisodia et al.
+(2024) extract interpretable product attributes from images and use them to estimate preferences.
+Our approach is simpler and scales better across product categories, and thus, it may be preferable in applications where interpretability of extracted attributes is not essential.
+Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities.
 Similarly, Lee (2024) employs large-language models to predict utility intercepts of new products based on their textual descriptions.
-These methods use texts or images to predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution.
+These methods predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution.
 [^7] For example, Lee (2024) inverts a plain logit model without random coefficients to extract product fixed effects for the prediction part of his analysis.
 Thus, his model restricts substitution patterns due to the IIA property.
-In contrast, our approach uses the same types of unstructured data to model utility covariances.
-This enables us to capture flexible substitution patterns—a key input for many empirical applications of interest.
-A few papers propose estimating substitution patterns using survey data, a possible alternative to our use of unstructured data.
-Dotson et al.
-(2019) have survey participants rate each product image and incorporate rating correlations as utility correlation shifters in their demand model.
+In contrast, our approach models utility covariances and can thus capture flexible substitution patterns—a key input for many empirical applications of interest.
+Further, while utility intercepts can be recovered from market shares, our goal is to capture substitution patterns that are unobserved.
+We measure them in an experiment by eliciting second choices, and we use these data for validation.
+[^8] A recent paper by Han and Lee (2025) uses unstructured data to estimate substitution patterns among fonts in order to study the effects of copyright protection policies.
+A possible alternative to our use of unstructured data is to estimate substitution patterns using survey data.
+For example, Dotson et al.
+(2019) ask survey participants to rate each product image and then incorporate rating correlations as shifters of utility correlations into a demand model.
 Similarly, Magnolfi et al.
-(2022) ask survey participants to rank products relative to each other (e.g., “Product A is closer to B than to C”).
-They generate low-dimensional product embeddings from these rankings and use them as attributes in a random coefficients logit model.
-Like our paper, Magnolfi et al.’s approach provides a way to crowdsource product attributes when collecting such data is challenging.
-While our approach can complement these survey-based methods, it has the advantage of using only widely available text and image data, eliminating the need to conduct costly category-specific surveys.
-[^7] Dew (2024) also uses survey data to estimate consumer preferences for text and image embeddings.
-However, his focus is on predicting individual choices rather than estimating demand and modeling substitution.
+(2022) elicit relative rankings from survey participants (e.g., “Product A is closer to B than to C”), generate embeddings from these rankings, and use them in a random coefficients logit model.
+Our approach complements these survey-based methods but has the advantage of using only widely available text and image data, thereby avoiding the need for costly category-specific surveys.
+Another complementary approach is proposed by McClure (2025), who shows how publicly available recommendations data from online platforms can also be used to estimate substitution patterns.
 Our approach is well-suited for several empirical applications.
 Researchers can use it to address a wide range of economics and marketing questions that require accurate and flexible estimates of product substitution.
-This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through changing assortments and prices.
-Additionally, marketers can apply this approach to help multi-product firms optimize prices and promotions (Hoch et al., 1995; Hitsch et al., 2021).
+This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through altering assortments and prices.
+Additionally, researchers can apply this approach to study how multi-product firms optimize prices and promotions (Hoch et al., 1995; Vilcassim and Chintagunta, 1995).
+Finally, because our approach circumvents the need to collect distinct attributes for each category, it may be valuable for researchers trying to estimate demand across many categories (Döpper et al., 2024).
+[^9] Döpper et al.
+(2024) remark that estimating demand across multiple categories “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach can address this problem if researchers have access to unstructured data.
 2 Proposed Approach
-Our approach involves three steps: (1) extracting embeddings from images and texts, (2) reducing the dimensionality of these embeddings using PCA, and (3) including the resulting principal components in a random coefficients logit model.
-In this section, we focus on the first two steps, as the mixed logit model is standard and does not require a detailed explanation.
-Section 3.2 describes the exact model specification used in our empirical applications.
-2.1 Extracting Embeddings from Texts and Images
-We extract embeddings from texts and images using pre-trained deep-learning models.
-This enables us to leverage models trained on large, general-purpose datasets, which can efficiently capture differences in product images and descriptions.
-Further, using pre-trained models reduces the computational burden of estimation.
-2.1.1 Image Embeddings
-We employ deep learning models that were originally trained for object detection and classification tasks.
-Specifically, we use four pre-trained convolutional neural networks available in the Keras Deep Learning Library: VGG19, ResNet50, InceptionV3, Xception.
-[^8] VGG19, a deep convolutional neural network with 19 layers (Simonyan and Zisserman, 2015).
-ResNet50, is a 50-layer convolutional neural network (He et al., 2016).
-InceptionV3 and Xception are convolutional neural networks with 48 and 71 layers (Szegedy et al., 2016; Chollet, 2017).
-All four models achieve high predictive accuracy on the ImageNet validation dataset, and are widely known for their strong performance in image classification tasks.
-[^9] The ImageNet data (https://www.image-net.org/) is used for the ImageNet Large Scale Visual Recognition Challenge (ILSVRC), which evaluates algorithms for object detection and image classification, a benchmark in object category classification (Russakovsky et al., 2015).
-We extract image embeddings from product photos, which researchers can easily obtain via online search or by scraping product pages from e-commerce platforms.
-Each model transforms the original image into a vector representation – an embedding – and classifies an image from its embedding by predicting the object it contains.
-These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”).
+The proposed approach involves three steps: (1) extracting embeddings from images and texts, (2) reducing the dimensionality of these embeddings using PCA, and (3) including the resulting principal components into a standard attribute-based logit model with random coefficients.
+Step 1. Extracting Embeddings from Texts and Images
+To extract information from product images, we extract their low-dimensional representations—embeddings—using pre-trained deep learning models.
+We use four convolutional neural networks: VGG19 (Simonyan and Zisserman, 2015), ResNet50 (He et al., 2016), InceptionV3 (Szegedy et al., 2016), and Xception (Chollet, 2017).
+[^10] These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”).
 However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings.
-Because these models perform well at distinguishing similar objects, we expect the embeddings to capture key visual features that differentiate products.
-We do not commit to a single model and intentionally use multiple deep learning models with different architectures.
-As we explain in Section 3.2, we select the one that best explains substitution patterns in the data.
-2.1.2 Text Embeddings
-Next, we extract information from text data, including product titles, descriptions, and customer reviews.
-We use two simple but widely used “count” models as well as pre-trained models with varying levels of sophistication, generating a separate set of embeddings for each model and type of text data.
-Before extracting embeddings, we pre-process text data as described in Appendix A.
-First, we use a bag-of-words count model, which transforms text into fixed-length vectors by counting word occurrences.
-This model does not consider word order and only measures how often words appear in a document.
-Second, we apply a bag-of-words model with a TF-IDF vectorizer.
-Although similar to the first method, TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words.
+The key advantage of using pre-trained models is that it reduces the computational burden of estimation while leveraging models trained on large-scale datasets, known for their strong performance in image classification tasks (Russakovsky et al., 2015).
+Because these models perform well at distinguishing visually similar objects, we expect embeddings to capture the key visual features that differentiate products.
+Rather than committing to a single model, we perform model selection to determine which one best captures substitution patterns, as described below.
+We also extract information from texts, including product titles, descriptions, and customer reviews.
+We use a bag-of-words model, which represents text as fixed-length vectors based on word counts, as well as a variation with a TF-IDF vectorizer.
+[^11] TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words.
 This makes it more likely that text embeddings capture distinctive words that differentiate products from one another.
-Next, we use the Universal Sentence Encoder (USE), a transformer-based language model (Cer et al., 2018), and the Sentence Transformer (ST) model, a pre-trained Sentence-BERT model designed to produce semantically meaningful sentence embeddings (Reimers and Gurevych, 2019).
+While these count models are relatively simple, we view them as useful benchmarks because they can still detect attributes mentioned in titles, descriptions, and reviews.
+In addition, we employ two pre-trained deep learning models: Universal Sentence Encoder (USE) (Cer et al., 2018), and BERT Sentence Transformer (ST) (Reimers and Gurevych, 2019).
 [^12] The model trained by Reimers and Gurevych (2019) is a more efficient version of the widely used BERT network (Devlin et al., 2018).
-Unlike the bag-of-words models, both USE and ST account for word order and the context in which words appear within sentences.
-Embeddings from these models are commonly used for text classification, semantic similarity analysis, clustering, and other natural language processing tasks.
-Both models achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017).
-While count models are relatively simple, they can detect individual attributes mentioned in titles, descriptions, and reviews.
-In contrast, the more advanced ST and USE models are better at capturing nuanced similarities in how sellers and consumers describe products.
-For example, they can identify when sellers or consumers describe the same functional benefits using similar language, even if the words are not exactly the same.
-Thus, we view count models as a useful benchmark that tells us whether our method merely extracts attributes mentioned in texts or captures more subtle cues about product substitution.
-2.2 Generating Principal Components
-Embeddings provide a lower-dimensional representation of the original texts and images.
-However, they remain high-dimensional compared to the number of attributes researchers typically include in demand models.
-For example, it would be impractical to incorporate 512-dimensional VGG19 embeddings into a random coefficients logit model, as it would require computationally intensive numerical integration over all dimensions.
-For this reason, we further reduce the dimensionality by applying Principal Component Analysis (PCA) to the embeddings.
-For a given category, we extract principal components from each set of embeddings corresponding to a given model and data type (e.g., USE embeddings from customer reviews) and incorporate them into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each component.
-Principal components are particularly appealing because they are orthogonal to each other, which mitigates multicollinearity and simplifies the estimation of random coefficients in the mixed logit model (Backus et al., 2021).
-We perform model selection by choosing the combination of principal components that minimizes the AIC (see Section 3.2.1).
-Because we apply PCA separately within each category, the extracted principal components capture the key dimensions that differentiate products within that category.
-By contrast, raw embeddings come from models trained on large, general-purpose datasets.
-For instance, image embeddings come from models trained to classify images into broad categories, such as “tablet,” “laptop,” or “cellphone,” and thus contain features that distinguish images across these categories.
-However, to estimate demand, we need to analyze consumer choices within a given category, where products often share visual features due to similar designs.
-PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution patterns within each category.
-2.3 Interpretation of Principal Components
-Standard mixed logit models capture substitution by measuring how similar products are in terms of their observed attributes.
-Similarly, our approach captures substitution by measuring how similar products are in terms of their visual designs and textual descriptions.
-There are several reasons why this proximity may be linked to substitution.
+Both USE and ST produce semantically meaningful sentence embeddings and achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017).
+As a result, they can identify when sellers or consumers describe the same product attributes and functional benefits using similar language, even if the exact words differ.
+Step 2. Generating Principal Components
+Although embeddings compress texts and images into lower-dimensional representations, they remain high-dimensional compared to the number of attributes typically used in demand models.
+For example, incorporating 512-dimensional VGG19 embeddings into a random coefficients logit model would require prohibitively costly numerical integration over all dimensions.
+We therefore apply PCA to further reduce dimensionality (Backus et al., 2021).
+The dimensionality reduction via PCA is attractive for multiple reasons.
+Raw embeddings are trained on general-purpose datasets to classify images into broad categories, such as “tablet” or “laptop.” However, to estimate demand, we need to analyze consumer choices within a given category.
+PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution within a category.
+Additionally, principal components are appealing because they are orthogonal to each other, which avoids multicollinearity issues and simplifies estimation of random coefficients in demand models.
+Step 3. Including Principal Components into a Demand Model
+We include principal components into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each included component.
+Since each deep learning model and data type produces different principal components, a key question is which to include in the demand model.
+[^13] While we could select the specification that best matches observed second choices in our experiment, this strategy is unavailable to most researchers who usually only have first-choice data.
+We thus propose a model selection algorithm that relies solely on first choices.
+We perform model selection using in-sample $AIC$.
+An advantage of using $AIC$ is that it can be interpreted as a bias-corrected estimator of the expected relative Kullback-Leibler (KL) information based on the maximized log-likelihood function (Akaike, 1998).
+Further, in our experimental results, in-sample $AIC$ strongly correlates with counterfactual performance, justifying its use for model selection (see Section 3.3 where we also implement $BIC$ and cross-validation).
+Our model selection algorithm considers a pool of candidate variables consisting of price and the first $P$ principal components.
+In our application, we choose $P=6$, which collectively account for 70-80% of variance in embeddings (see Appendix Figure A4).
+We first estimate all possible models with a random coefficient on a single variable and retain the best-fitting model.
+We then estimate all possible models with two random coefficients and assess whether fit improves relative to the best single-variable model.
+We continue to increase the number of variables $K$ that have random coefficients until there is no further $AIC$ improvement.
+We repeat this across all specifications—defined as a combination of a text or image model and data type (e.g., USE Reviews)—and choose the one with the lowest overall $AIC$.
+The full procedure is given in Algorithm 1.
+This algorithm avoids costly combinatorial search over all possible specifications.
+At the same time, it ensures that we do not rely too heavily on the ordering of principal components, which reflects the explained variance of embeddings but need not be predictive of preferences.
+For example, the approach might select a model that has random coefficients on the second and fourth principal components, but not the first or third.
+      Algorithm 1: Model Selection
+Input: - Products $j=1,\dots,J$ - Specifications $m=1,\dots,M$ (all combinations of unstructured data and pre-trained text or image models) - Maximum number of principal components $P$ (researcher’s choice) - Demand model with utility: $u_{ij}=\delta_{j}+\alpha_{i}\text{price}_{ij}+\theta_{i}^{\prime}PC_{j}+\varepsilon_{ij}$ ($\delta_{j}$ = product fixed effects)
+1: For each specification $m$ do
+2:   Extract embeddings $e^{(m)}=(e_{1}^{(m)},\dots,e_{J}^{(m)})$
+3:   Extract principal components $PC_{1}^{(m)},\dots,PC_{P}^{(m)}$ from embeddings $e^{(m)}$ using PCA
+4:   Initialize model selection at $K\leftarrow 0$ (zero random coefficients)
+5:   Set $\text{BestSpec}^{(m)}$ as Plain Logit and $\text{BestAIC}^{(m)}$ as the corresponding AICAIC
+6:   while $\text{BestAIC}^{(m)}$ decreases do
+7:     $K\leftarrow K+1$
+8:     Estimate mixed logit models with all possible combinations of random coefficients
+     on subsets of $K$ variables from the candidate set $\mathcal{R}=\{\text{price, P$C_{1}^{(m)}$, $\dots$, P$C_{P}^{(m)}$}\}$
+9:     Find subset $R_{K}^{*}\in\mathcal{R}$ that minimizes $AIC$ at $AIC_{K}^{*}$
+10:     if $AIC_{K}^{*}<\text{BestAIC}^{(m)}$ do
+11:       $\text{BestSpec}^{(m)}\leftarrow R_{K}^{*}$ update the best specification
+12:       $\text{BestAIC}^{(m)}\leftarrow AIC_{K}^{*}$ update the lowest AICAIC
+13:   end while
+14: end for
+15: Choose the best-fitting specification $m^{*}=\arg\min_{m}\text{BestAIC}^{(m)}$
+Why Does This Approach Work?
+There are several reasons why texts and images may predict substitution.
 Consumers may choose based on how a product looks and how it is described by sellers or other consumers.
-For instance, a consumer who finds one product visually appealing may also be drawn to similar-looking products, and they might expect products with similar descriptions to meet their needs in a similar way.
-Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect key product features and benefits.
+Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect attributes that drive substitution.
 Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly.
-Such texts may convey functional benefits that are hard to quantify.
-Similarly, photos may showcase design features like color, casing style, or the presence of a keyboard.
-By extracting principal components from texts and images, our approach identifies these dimensions of product differentiation and uses them to capture which products are close substitutes.
-Given this interpretation, researchers should not treat the coefficients on the principal components as causal in the sense that changing product images or textual descriptions does not necessarily cause a change in substitution patterns.
-For example, if the principal components serve as proxies for choice-relevant product attributes, these attributes need not change if the seller modifies the product’s description or visual design.
-Hence, the principal components should be held fixed in counterfactual analyses.
-For this reason, our approach is not suited for studying optimal product design and positioning, or any other questions that require causal estimates of the effects of non-price characteristics on choices.
-In some markets, texts and images might not convey information relevant to consumer choices, in which case the principal components would fail to predict substitution patterns.
-However, we consider this scenario unlikely, as sellers have strong incentives to highlight important product attributes in titles and descriptions, and online platforms may encourage them to do so.
-Similarly, when writing reviews, consumers often mention product aspects they think others will find useful.
-Consistent with this idea, we find that in all product categories we study, at least some image- or text-based principal components predict substitution patterns.
-Similar to observed product attributes in standard mixed logit demand models, we treat principal components as fixed over time.
-This assumption is reasonable because principal components capture time-invariant product attributes, functional benefits, and visual design features.
-While sellers could theoretically modify images and textual descriptions to reposition their products, we find that in practice, product images, titles, and descriptions rarely change in our empirical context (see Appendix C for details).
-2.4 Advantages of Our Approach
-To understand how our approach relates to existing methods, we compare it to the standard approach commonly used in demand estimation, which relies on observed product attributes.
-These attributes are usually incorporated into a discrete choice demand model, such as a random coefficients logit model, where substitution patterns are captured through the interaction between product attributes and preference heterogeneity.
-While common, this approach presents two practical challenges.
-First, researchers rarely observe all attributes that differentiate products.
-Instead, they often rely on third-party datasets, where data providers report a limited set of attributes selected based on unknown criteria.
-Others gather their own data using a combination of web scraping and human coders—–a costly and time-consuming process that typically limits researchers to collecting only easily quantifiable attributes.
-Since attributes are selected subjectively in both cases, they may not align with those most relevant to consumer choices.
-Our approach avoids collecting data on product attributes and instead extracts information about product substitution from text and image data.
-The second challenge is that consumers often consider visual design and functional benefits—dimensions that are difficult to capture through observed attributes.
-For instance, in categories like clothing or home decor, product aesthetics can strongly influence choices.
-A shopper selecting a dress might care about overall style, design pattern, or neckline type—details visible in photos but difficult to quantify without training a custom classification model.
-Similarly, textual descriptions might reveal attributes that are hard to quantify.
-Customer reviews, for instance, might highlight that certain headphones are ideal for listening to audiobooks, revealing a benefit not captured by observed attributes.
-Our approach allows researchers to estimate substitution patterns along these otherwise unobserved dimensions.
-A natural consequence of using text and images is that our approach scales more easily across categories than standard attribute-based models.
-For instance, Döpper et al.
-(2024) estimate demand estimation across a large set of categories and note that incorporating product attributes would be difficult because it would require them to decide which attributes are relevant in each category and collect data on those attributes.
-Our approach circumvents the need to collect category-specific attributes, making it easier to estimate demand across many categories.
-This scalability may also be valuable for online retailers like Amazon or Alibaba looking to optimize prices across hundreds of categories.
-In Section 4, we further illustrate this point by applying our approach across 40 Amazon categories.
-3 Choice Experiment
-We begin by applying our method to data from a choice experiment.
-In this experiment, we randomize product prices and positions, generating clean variation that helps estimate substitution patterns.
-We also collect data on consumers’ second choices, which are rarely available in observational datasets.
-These data allow us to assess how well our model recovers substitution patterns by testing its ability to predict second choices counterfactually relative to alternative models.
+Similarly, photos may showcase design features like color and casing style.
+Our approach extracts this information and uses it to capture substitution.
+Given this interpretation, embeddings correlate with substitution patterns because they capture choice-relevant attributes, but the relationship need not be causal: changing product images or descriptions will not necessarily alter substitution patterns.
+Consequently, we cannot study questions like optimal product design or positioning unless we further unpack the link between embeddings and substitution, which is outside the scope of this paper.
+By contrast, our method is well-suited for counterfactuals where embeddings can be held fixed, such as optimal pricing or merger simulations.
+We provide a more detailed discussion of the counterfactuals enabled by our approach in Section 5.2.
+Advantages Over Standard Methods
+Our approach has several advantages relative to the standard attribute-based methods.
+First, researchers typically select a limited set of attributes based on their prior knowledge of the market, often those that are easiest to quantify, or rely on attributes supplied by data providers.
+Our approach avoids these subjective choices by automatically extracting information about product substitution from unstructured data.
+Second, this approach captures visual design and functional benefits, which may drive substitution but are difficult to capture through observed attributes.
+Lastly, we circumvent the need to collect category-specific attributes, which makes our approach more scalable.
+This is valuable for researchers seeking to estimate demand and study competition and pricing across many product categories.
+3 Validation with Experimental Data
+We apply our method to data from a choice experiment to show it recovers substitution patterns more effectively than standard attribute-based methods.
+A key challenge is that we usually do not directly observe substitution patterns in the data, and thus do not have a “ground truth” to validate demand models.
+To address this, we design and implement a choice experiment that measures second-choice diversion ratios, which directly reflect substitution.
 3.1 Experiment Design
-We recruited participants from the online platform Prolific and asked each participant to complete two choice tasks via a Qualtrics survey.
-In the first choice task, we asked each participant to choose one book from a set of ten alternatives (Figure 1).
-These ten books were chosen from Amazon’s bestseller lists (Figure 2).
-We chose them from three different genres—Mystery, Fantasy, and Self-Help—anticipating that genre would predict substitution patterns and could thus be incorporated in attribute-based demand models.
-We instructed each participant to choose the book they would purchase if faced with this selection in a real bookstore.
-Figure 1 shows an example of a choice task, displaying the top portion of the page as it appeared to participants.
-For each book, we showed participants the cover image, title, author, publication year, genre, number of pages, description, and five customer reviews, all collected from Amazon product pages.
-We presented books in a random order to each participant and drew each book’s price from a discrete uniform distribution ranging from \$3 to \$7.
+Choice Tasks.
+Each participant completes two choice tasks.
+In the first choice task, a participant chooses one book from a set of ten alternatives.
+We instruct participants to choose books they would purchase if faced with this selection in a real bookstore.
+We limit the number of options to ten to ensure participants pay attention to most options, which allows us to abstract away from limited consideration.
+After a participant selects a book from the list of ten options, we remove this book from the list.
+The participant then proceeds to a second choice task, where they choose among the remaining nine books.
+Figure 1 shows the choice task as presented to participants.
+To give all models a fair chance at capturing substitution, we display attributes (author, year, genre, pages), cover images, and texts (plot description and five reviews), collected from Amazon product pages.
+We take all books from Amazon’s bestseller lists, sampling them in a way that does not bias our model comparisons in favor of any particular specification (see Appendix B for details).
+Lastly, we randomize books’ rankings and prices across participants, keeping them fixed across choice tasks for the same participant.
+[^14] Prices were drawn from a discrete uniform distribution ranging from \$3 to \$7.
+This double randomization generates clean variation for estimating substitution patterns (Berry and Haile, 2024).
+[^15] Although all participants technically see the same set of ten books in the first choice task, they rarely choose books that are displayed near the bottom of the list.
+Therefore, by randomizing the order in which books are displayed to participants, we vary their effective choice sets.
+The book category is well-suited for our analysis because participants are likely to consider both structured attributes (e.g., genre) and unstructured information (e.g., plot descriptions).
+It is also unclear a priori whether substitution patterns are better predicted by images or texts—while texts describe rich plot details, cover images contain clear cues of genres, and whether the book is fiction or non-fiction (see Section 3.3 for a detailed discussion).
 Figure 1: Example of a choice task in our experiment. The screenshot displays the top portion of the page as it appeared to participants.
-After a participant selected a book from the list of ten options, we removed this book from the list.
-The participant then proceeded to a second choice task, where they chose from the remaining nine books.
-The positions and prices of the books remained the same as in the first choice task, with the only difference being that the participant’s first choice was no longer available.
-Based on power calculations from a pilot experiment, we estimated that we needed around 10,000 participants to determine whether our model outperforms alternative models in predicting second choices.
-We pre-registered this target sample size.
-[^11] The pre-registration document for this study can be accessed at https://tinyurl.com/ekjp4pfp.
-In total, we recruited 10,775 participants between June 14 and June 27, 2024.
-Following the pre-registered sample selection criteria, we excluded around 14% participants who failed comprehension questions, did not complete the survey, or spent less than one minute on the entire study.
-After applying these criteria, our final sample consists of 9,265 participants.
+Recruitment and Sample selection.
+We recruited 10,775 participants from the online platform Prolific between June 14 and June 27, 2024.
+This sample size is close to our pre-registered target of 10,000, determined from power calculations in a pilot experiment.
+[^16] The pre-registration document can be accessed at https://tinyurl.com/ekjp4pfp.
+At the time of pre-registration, we planned to estimate a different choice model—the pairwise combinatorial logit (PCL) model from Koppelman and Wen (2000)—which generalizes the logit model by allowing utility correlations across product pairs.
+We switched to a mixed logit model at a later stage because we found it to perform better.
+We adhered to our pre-registered protocol for all key aspects of survey design, including sample size, sample selection, and choice task construction.
+Following the pre-registered sample selection criteria, we excluded 14% participants who failed comprehension questions, did not complete the survey, or spent less than one minute on the study, leaving a final sample of 9,265 participants.
+Because the choice tasks were hypothetical and not incentivized, it is important to verify that participants made meaningful selections.
+In Appendix C, we show that participants did not rush through the survey, responded to changes in book rankings and prices, and made choices consistent with their self-reported genre preferences.
+3.2 Model Specifications and Estimation Results
+We compare our approach against two benchmark models: the plain logit model and a standard logit model with random coefficients model on attributes.
+All three models fall into a framework where the indirect utility of participant $i$ from book $j$ is
+$u_{ij}=\beta_{i}^{\prime}x_{j}+\theta_{i}^{\prime}PC_{j}+\gamma\cdot\text{rank}_{ij}+\alpha_{i}\cdot\text{price}_{ij}+\delta_{j}+\varepsilon_{ij}.$ (1)
+Here $x_{j}$ are observed book attributes: genre dummies, publication year, and length in pages—all attributes participants observe in the choice tasks; $PC_{j}$ are principal components extracted from embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in participant $i$’s choice task, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d.
+taste shocks following a Type I Extreme Value distribution.
+We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books.
+[^17] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns using observed characteristics or principal components.
+This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does.
+We estimate three demand models:
+  1. 1.
+[…]
+  2. 2.
+[…]
+  3. 3.
+    Plain Logit.
+We include neither attributes $x_{j}$ nor principal components $PC_{j}$, estimating a constant price coefficient $\alpha_{i}=\alpha$, rank coefficient $\gamma$ and fixed effects $\delta_{j}$.
+Table 1 compares in-sample $AIC$.
+The best mixed logit with attributes includes random coefficients on the number of pages and year, reducing $AIC$ by $16.0$ relative to plain logit.
+The best model with unstructured data, Reviews USE, puts random coefficients on the first two principal components.
+This model fits better than any other considered specification, reducing $AIC$ by $24.8$ relative to plain logit and by $8.8$ relative to the best attribute-based logit.
+While this superior fit is reassuring, it does not guarantee better counterfactual performance.
+We therefore turn to predicting second choices counterfactually.
+[…]
+Plain Logit None 41006.7 0.0
+Mixed Logit with Attributes Pages and Year 40990.7 -16.0
+Mixed Logit with Images (InceptionV3) Price, PC1, and PC6 40990.5 -16.2
+Mixed Logit with Titles (ST) PC1 and PC5 40992.3 -14.4
+Mixed Logit with Descriptions (USE) PC1 and PC5 40986.4 -20.3
+Mixed Logit with Reviews (USE) PC1 and PC2 40981.9 -24.8
+Table 1: Comparison of models in terms of in-sample AICAIC on first choices.
+The second column shows variables that have random coefficients in the selected specification.
+The last column shows the $AIC$ reduction relative to plain logit.
+3.3 Validation Using Second-Choice Data
+To evaluate demand models, we let them confront a difficult counterfactual prediction problem—predicting second choices, i.e.
+which books participants switch to when their most preferred option becomes unavailable.
+A model can only predict second choices well if it has accurately captured the true substitution patterns.
+Predicting second choices is also directly relevant in antitrust where second-choice diversion ratios are often used to measure substitutability and the intensity of price competition (Conlon and Mortimer, 2021).
+Operationally, we first estimate each model via Maximum Likelihood Estimation (MLE) using first choices only.
+We then use the estimated model to predict counterfactual second-choice diversion ratios, $\hat{s}_{j\rightarrow k}$, defined as the probability that the participant chooses product $k$ in the second choice task conditional on having chosen book $j$ in the first choice task.
+We compare these predictions with diversions observed in the data, $s_{j\rightarrow k}$, computing $RMSE$ as:
+$RMSE=\sqrt{\frac{1}{J(J-1)}\sumop\displaylimits_{j}\sumop\displaylimits_{k\neq j}\left(s_{j\rightarrow k}-\hat{s}_{j\rightarrow k}\right)^{2}}$ (2)
+Since we do not use second-choice data in estimation, lower $RMSE$ indicates that the model performs better at counterfactual predictions of second choices.
+Appendix D provides further details on how we compute diversions $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$.
+Figure 2: Comparison of models in terms of counterfactual RMSERMSE on second choices.
+The two benchmarks in the first panel are the plain logit without random coefficients and the best-fitting mixed logit with random coefficients on observed attributes.
+The remaining specifications correspond to mixed logit models with random coefficients on principal components extracted from image or text embeddings.
+Figure 2 summarizes the validation results, while Appendix Table A1 reports $AIC$ and $RMSE$ values for all estimated specifications.
+The top panel in Figure 2 shows two benchmarks: the plain logit and the lowest-$AIC$ mixed logit with attributes.
+The mixed logit with attributes reduces RMSE by 11.7%, while our best-fitting Review USE model reduces RMSE by 23% relative to plain logit, significantly outperforming the mixed logit with attributes.
+This result illustrates the value of our approach.
+We chose books for this study expecting observed attributes like genre to predict substitution well.
+Yet, by using unstructured data, we can match the counterfactual performance of the mixed logit with attributes and even further reduce $RMSE$ by 14%.
+This shows that our approach recovers substitution patterns better than standard attribute-based methods in this dataset.
+Although the model with product reviews performs best, in general, performance varies widely across specifications.
+Below, we discuss why some specifications recover substitution patterns better than others.
+Images
+All four image models outperform the plain logit model, with the lowest-$AIC$ model, InceptionV3, reducing $RMSE$ by 7.0% relative to plain logit.
+To understand why images predict substitution, consider the book covers used in our study (see Figure 3).
+Within the same genre, book covers often share similar design elements.
+For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, and feature natural objects like twisted vines and golden roses.
+Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds.
+Thus, image embeddings partly encode books’ genres, which correlate with substitution patterns.
 Mystery Books Refer to caption “The Inmate” Refer to caption “Please Tell Me” Refer to caption “The Ritual” Refer to caption “The Housemaid”
@@
 Figure 3: Ten books used in our experiment.
-Because choice tasks in our experiment were hypothetical and not incentivized, it is important to verify that participants made meaningful selections.
-In Appendix B, we show that participants did not rush through the survey, responded to changes in book rankings and prices, and made choices consistent with their self-reported genre preferences.
-3.2 Model Specifications
-We compare our approach against two benchmark models: the plain logit model and a standard random coefficients model that leverages observed product attributes but uses no texts or images.
-All three models fall into the following framework, where the indirect utility of consumer $i$ from purchasing book $j$ is given by
-$u_{ij}=\beta_{i}^{\prime}x_{j}+\theta_{i}^{\prime}PC_{j}-\gamma\cdot\text{rank}_{ij}-\alpha_{i}\cdot\text{price}_{ij}+\delta_{j}+\varepsilon_{ij}.$ (1)
-Here $x_{j}$ is a vector of observed book attributes, $PC_{j}$ are principal components extracted from text and image embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in consumer $i$’s choice tasks, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d.
-taste shocks following a Type I Extreme Value distribution.
-We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books.
-This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does.
-When specifying observed attributes $x_{j}$, we include genre dummies for Mystery and Sci-Fi, publication year, and length in pages—all attributes participants observe in the choice tasks.
-We estimate the following three models within this framework: [^12] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns related to product features, captured through observed characteristics or principal components from text and image data.
-  1. 1.
-    Mixed Logit with Principal Components.
-We include principal components $PC_{j}$, interacting them with random coefficients $\theta_{i}\sim N(0,\Sigma_{\theta})$, where $\Sigma_{\theta}$ is diagonal.
-Since the principal components do not vary for a given product over time, the mean of the random coefficients is absorbed by the product fixed effects $\delta_{j}$.
-In some specifications, we also estimate a random coefficient on price, assuming $\alpha_{i}\sim N(\bar{\alpha},\sigma_{\alpha})$.
-We do not include observed attributes, setting $\beta_{i}=0$.
-This model mimics a scenario where a researcher has access to unstructured data but does not observe attributes other than price.
-  2. 2.
-    Mixed Logit with Observed Attributes.
-We include observed attributes $x_{j}$, interacting them with random coefficients $\beta_{i}\sim N(0,\Sigma_{\beta})$, where $\Sigma_{\beta}$ is diagonal, and again, the mean is absorbed by product fixed effects.
-In some specifications, we also estimate a random coefficient on price, assuming $\alpha_{i}\sim N(\bar{\alpha},\sigma_{\alpha})$.
-We do not include principal components $PC_{j}$, setting $\theta_{i}=0$.
-This model reflects the standard practice of estimating substitution patterns through an interaction of observed attributes with consumer heterogeneity.
-  3. 3.
-    Plain Logit.
-We include neither attributes $x_{j}$ not principal components $PC_{j}$, estimating only the uniform price coefficient $\alpha_{i}=\alpha$, the rank coefficient $\gamma$ and product fixed effects $\delta_{j}$.
-This model provides a baseline where product substitution is solely driven by the average utilities of books and not their correlations.
-Comparing plain logit to mixed logit with observed attributes tells us whether adding random coefficients on book attributes improves the estimation of substitution patterns.
-Similarly, comparing mixed logit with attributes to mixed logit with principal components tells us whether incorporating texts and images into demand estimation enables us to recover substitution patterns as well as, or better than, using product attributes.
-[^13] To ensure a fair comparison between models, for each model we start optimization from 100 different initial points covering a wide range of parameter values.
-This helps us maximize the probability of finding the global optimum in each model.
-3.2.1 Model Selection
-Within each of the three model classes listed in Section 3.2, we perform model selection based on $AIC$ [^14] An advantage of using AIC is that it can be interpreted as an estimator of the expected relative Kullback-Leibler (KL) Information based on the maximized log-likelihood function, corrected for asymptotic bias (Akaike, 1998; Burnham and Anderson, 2004).
-In the mixed logit model with observed attributes, we consider all possible combinations of random coefficients across four book attributes: price, publication year, length in pages, and genre.
-Genre is represented by two dummy variables, one each for Mystery and one for Self-Help, making Sci-Fi the baseline genre.
-We estimate all 16 specifications and select the one with the lowest $AIC$.
-In the mixed logit model with principal components, we take the first three principal components from each set of embeddings and evaluate all possible combinations of random coefficients on the three components and price, selecting the specification with the lowest $AIC$ [^15] Increasing the number of principal components beyond three has only a marginal impact on our results and does not affect the in-sample fit or counterfactual performance of our selected model.
-We repeat this selection process for every text and image model and data type (e.g., USE embeddings from reviews).
-Lastly, among the selected specifications, we choose the one that minimizes $AIC$.
-We do not combine principal components extracted from different embeddings (e.g, mixing texts with images) in our main analysis.
-As we discuss later in Section 3.4.4, we tested mixing data types in our experimental data but found no improvement in model fit.
-3.3 Validation Approach
-Prior research often evaluated demand models based on how intuitive the estimated cross-price elasticities look, given researchers’ prior knowledge of the market (Berry et al., 1995; Nevo, 2001).
-We follow a different approach by leveraging second choice data to compare our approach against alternatives.
-We estimate each model using only first choices and evaluate how well it predicts second choices.
-[^16] Conceptually, our validation approach reverses the logic of Berry et al.
-(2004), who show that second choice data are useful for estimating substitution patterns.
-By contrast, we omit second choice data from estimation and use them only for model validation.
-Since the second-choice data are not used in estimation, our validation approach forces alternative models to confront a difficult counterfactual prediction problem.
-Second choices reveal which books consumers substitute to when their most preferred option becomes unavailable.
-Therefore, a model can only predict second choices well if it has accurately captured the true substitution patterns.
-Predicting second choices is also directly relevant in some applications.
-For instance, antitrust authorities can use them in merger analysis to calculate diversion ratios—a measure reflecting substitutability and thus the intensity of price competition between products (Conlon and Mortimer, 2021).
-We discuss this further in Section 3.6.
-We evaluate each model as follows.
-First, we obtain MLE estimates by maximizing the total likelihood of participants’ first choices.
-Given these point estimates $\hat{\theta}$, we then compute the counterfactual second choice probabilities for each participant conditional on their first choice.
-Averaging these probabilities across participants, we construct the matrix $\mathbb{P}$, where each element $(j,k)$ represents the probability that a participant whose first choice is $j$ selects $k\neq j$ as their second choice.
-To assess each model, we compute the $RMSE$ based on element-wise differences between the matrix $\hat{\mathbb{P}}$ predicted by the model and $\mathbb{P}$ computed from empirical probabilities in the experimental data.
-A lower $RMSE$ indicates better model performance in predicting counterfactual second choices.
-3.4 Validation Results
-We compare the results of three models described in Section 3.2: (1) the plain logit model, (2) the mixed logit model with observed attributes, and (3) the mixed logit model with principal components.
-We evaluate these models based on the $RMSE$ of their counterfactual second-choice predictions.
-Figures 3 and 4 summarize the validation results, whereas Appendix Table A1 reports the $AIC$ and $RMSE$ values for all specifications considered within each model class.
-3.4.1 Plain Logit
-We first report the results from the plain logit model.
-As shown in Appendix Table A1, this model achieves a counterfactual $RMSE$ of $0.089$.
-Because this model assumes independence of irrelevant alternatives (IIA), it cannot capture which close substitutes participants switch to when their first choices become unavailable.
-As a result, the model predicts relatively uniform second-choice probabilities that do not match the second choice data.
-3.4.2 Mixed Logit with Principal Components
-Next, Figure 3 presents results from the mixed logit model with principal components.
-The first panel shows $RMSE$ from the plain logit model, indicated by a vertical dashed line, while other panels display results from our proposed approach, which incorporates principal components obtained from different types of unstructured data and machine learning models.
-Each marker reflects the second-choice $RMSE$ from the best configuration of random coefficients on the principal components and price chosen by our model selection algorithm (see Section 3.2.1).
-Performance varies widely across specifications.
-While some models perform similarly to the plain logit, others deliver significantly better counterfactual predictions.
-These differences highlight that researchers should always perform model selection to find the specification that best captures substitution patterns from text and images.
-In practice, researchers can rely on $AIC$ as the main selection criterion.
-This works well in our experimental data, as $AIC$ selects the model that also achieves the lowest counterfactual $RMSE$ (see Appendix Table A1).
-Figure 3: Counterfactual performance of mixed logit models with principal components in terms of predicting second choices.
-Figure 4: Counterfactual performance of attribute-based mixed logit models in terms of predicting second choices.
-Images
-Figure 3 shows that image-based mixed logit specifications perform well.
-All models in this class outperform the plain logit model, and the lowest-$AIC$ model, InceptionV3, reduces $RMSE$ by 9.8% relative to the plain logit model (from $0.089$ to $0.080$).
-To understand why images predict substitution, consider the book covers used in our study (see Figure 2).
-Within the same genre, book covers often share similar design elements.
-For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, feature natural objects like twisted vines and golden roses, and have similar fonts and title placements.
-Two of these books are by the same author and have nearly identical covers.
-Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds.
-Given these visual similarities, we expect image embeddings in our models to capture genre information, which correlates with substitution patterns.
-However, we emphasize that image embeddings might be informative about substitution beyond simply revealing the book’s genre.
-For example, second-choice data shows that participants perceive mystery books as closer substitutes to fantasy books than self-help titles (Appendix Figure A3).
-Our image embeddings capture this pattern because mystery and fantasy book covers tend to be similar.
-Both use dark or muted tones with dramatic highlights (e.g., gold, silver, teal, or yellow) to create contrast, which is captured by our image embeddings.
 Texts
-The remaining panels of Figure 3 show the performance of text-based models.
-We find that counterfactual second-choice predictions improve when we either use more advanced text models or incorporate richer text data.
-First, counterfactual performance improves substantially when we move from simple bag-of-words models to the more advanced USE and ST models.
-While bag-of-words models provide little to no improvement over the plain logit, both USE and ST consistently outperform the plain logit across all types of text data.
-This may be because textual descriptions contain detailed information about book plots but do not always use the exact same words or phrases.
-Therefore, extracting substitution patterns from text requires a natural language model that can accurately measure semantic similarity.
-Second, counterfactual performance improves as we include more text data into the demand model.
-For instance, the USE model achieves a second-choice $RMSE$ of $0.080$ with book titles, $0.075$ with descriptions, and $0.069$ with reviews.
-This indicates that customer reviews provide more information about substitution patterns than descriptions, which, in turn, are more informative than titles.
-These results make intuitive sense.
+All text models outperform the plain logit, and their relative performance highlights several notable patterns.
+First, performance tends to improve when we move from simple bag-of-words models to the more advanced USE and ST models.
+This is particularly true for descriptions and reviews, which contain detailed information about book plots but do not always describe them in the same words or phrases.
+Therefore, extracting substitution patterns from text requires natural language models that can accurately measure semantic similarity.
+Second, performance improves as we include richer text data.
+For instance, the USE model reduces $RMSE$ by 4.3% with book titles, 17% with descriptions, and 23% with reviews relative to the plain logit.
+This makes intuitive sense.
 Although book titles are brief, they often contain subtle genre cues.
 For example, mystery titles frequently hint at secrets (Please Tell Me) or characters in confined situations (Housemaid, Inmate), whereas fantasy titles often include words that evoke a gothic atmosphere (Serpent, Wings, Ravens, Ruin).
-Book descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals.
-Reviews go a step further: consumers summarize plots while also expressing opinions.
-When reviewing mystery novels, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings.
-As with images, texts can provide information about substitution beyond simply revealing genres.
-For example, book descriptions and reviews make it clear that the self-help books Dopamine Detox and Don’t Believe Everything You Think both discuss how modern distractions and mental habits negatively affect well-being.
-In contrast, another self-help book, The Art of Letting Go, focuses on letting go of painful memories and emotions.
-These differences may explain why our text-based models correctly identify the first two books as much closer substitutes, something alternative models fail to detect (see Section 3.5 for details).
-[…]
-3.4.3 Mixed Logit with Observed Attributes
-Figure 4 presents the results for the mixed logit model with observed attributes.
-We compare it against two benchmarks: the plain logit model and the lowest-$AIC$ text-based model identified earlier.
-Several specifications significantly outperform the plain logit model.
-The lowest-$AIC$ specification is a mixed logit model with random coefficients on pages, year, and genre.
-[^17] To be precise, this model achieves the same $AIC$ as two other mixed logit models (i.e., “Price, Pages & Year” and “Pages & Year”), but all three specifications generate identical second-choice predictions, and thus, achieve the same $RMSE$.
-By capturing similarities in observed attributes, this model breaks the IIA pattern and generates more accurate second-choice predictions.
-Consequently, it reduces $RMSE$ by 12.9% compared to plain logit (from $0.089$ to $0.077$).
-3.4.4 Summary
-Taken together, these findings illustrate the value of our approach, showcasing its ability to extract information about substitution patterns from text and images.
-We deliberately chose books for this study, thinking that observed attributes like genre would strongly predict substitution.
-Yet, we find that many text-based models achieve comparable $RMSE$ to that of attribute-based models.
-Moreover, the best-fitting text-based model reduces $RMSE$ by an additional 10% relative to the best attribute-based model.
-Thus, incorporating text improves counterfactual predictions beyond what is achievable with observed attributes.
-One might wonder whether texts, images, and attributes provide distinct information about substitution patterns.
-If so, the best approach to estimating demand would be to incorporate all three data types.
-To explore this, we start with our selected review-based specification and examine whether model fit can be improved by adding (1) any combination of observed product attributes or (2) any combination of principal components from ResNet, the lowest-$AIC$ image model.
-In both cases, we find no fit improvements.
-This result suggests information across the three data types is highly correlated, with text providing the strongest signal of substitution.
+Descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals.
+Reviews go even further: consumers summarize plots while also expressing opinions.
+When reviewing mystery novels, for instance, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings.
+Thus, it is unsurprising that reviews provide more information about substitution than descriptions, which, in turn, are more informative than titles.
+Extension: Combining Data Types
+If text, images, and attributes provides distinct signals of substitution, combining them might improve fit and counterfactual performance.
+To explore this, we start from the selected Review USE model and attempt to extend it in two ways: (1) by adding a combination of observed product attributes with random coefficients, or (2) by adding a combination of principal components from InceptionV3, the lowest-$AIC$ image model.
+Neither extension improves $AIC$ or $RMSE$.
+This result suggests information across data types is highly correlated, with text providing the strongest signal of substitution.
+What do Principal Components Capture?
+To better understand the variation captured by embeddings, in Figure 4 we show book locations in the space of the two principal components selected into our best-fitting Review USE model.
+These principal components align with intuitive substitution patterns: the first one (horizontal axis) separates non-fiction on the left from fiction on the right, while the second (vertical axis) further distinguishes science fiction from mystery.
+Crucially, the variation in these principal components goes beyond separating genres.
+For example, The Housemaid and The Inmate are by the same author, and The Serpent & The Wings and The Ashes & The Star Cursed King are from the same book series.
+The two principal components detect this similarity from consumers’ reviews of these books.
+As evident from $RMSE$ comparisons in Figure 2, this additional variation helps recover substitution patterns better and improves counterfactual predictions.
+Figure 4: Book locations in the space of selected principal components (Review USE model).
 Panel A. Predicted Second-Choice Probabilities when First Choice is Dopamine Detox (S)
@@
 Book Prob. Book Prob. Book Prob. Book Prob.
-Don’t Believe (S) 0.353 Please Tell Me (M) 0.182 Don’t Believe (S) 0.221 Don’t Believe (S) 0.259
-Art of Letting Go (S) 0.249 The Inmate (M) 0.152 The Inmate (M) 0.153 Art of Letting Go (S) 0.176
-Please Tell Me (M) 0.112 The Housemaid (M) 0.141 Please Tell Me (M) 0.152 Please Tell Me (M) 0.129
-The Inmate (M) 0.094 Don’t Believe (S) 0.137 Art of Letting Go (S) 0.135 The Inmate (M) 0.125
-The Housemaid (M) 0.057 Serpent & Wings (F) 0.095 The Housemaid (M) 0.132 The Housemaid (M) 0.100
-Serpent & Wings (F) 0.042 Art of Letting Go (S) 0.094 Court of Ravens (F) 0.101 Court of Ravens (F) 0.074
-Court of Ravens (F) 0.034 Court of Ravens (F) 0.080 Serpent & Wings (F) 0.057 The Ritual (M) 0.049
-The Ritual (M) 0.031 Ashes & Star (F) 0.065 The Ritual (M) 0.028 Serpent & Wings (F) 0.044
-Ashes & Star (F) 0.030 The Ritual (M) 0.054 Ashes & Star (F) 0.022 Ashes & Star (F) 0.043
+Don’t Believe (S) 0.353 Please Tell Me (M) 0.171 Don’t Believe (S) 0.220 Don’t Believe (S) 0.266
+Art of Letting Go (S) 0.249 The Inmate (M) 0.147 Please Tell Me (M) 0.153 Art of Letting Go (S) 0.169
+Please Tell Me (M) 0.112 Don’t Believe (S) 0.143 Art of Letting Go (S) 0.142 Please Tell Me (M) 0.134
+The Inmate (M) 0.094 The Housemaid (M) 0.139 The Housemaid (M) 0.134 The Inmate (M) 0.123
+The Housemaid (M) 0.057 Art of Letting Go (S) 0.106 The Inmate (M) 0.131 The Housemaid (M) 0.101
+Serpent & Wings (F) 0.042 Court of Ravens (F) 0.096 Court of Ravens (F) 0.101 Court of Ravens (F) 0.070
+Court of Ravens (F) 0.034 Serpent & Wings (F) 0.088 Serpent & Wings (F) 0.060 Serpent & Wings (F) 0.057
+The Ritual (M) 0.031 The Ritual (M) 0.057 The Ritual (M) 0.031 The Ritual (M) 0.043
+Ashes & Star (F) 0.030 Ashes & Star (F) 0.054 Ashes & Star (F) 0.027 Ashes & Star (F) 0.037
 Panel B. Predicted Second-Choice Probabilities when First Choice is Please Tell Me (M)
@@
 Book Prob. Book Prob. Book Prob. Book Prob.
-The Inmate (M) 0.325 Don’t Believe (S) 0.164 The Inmate (M) 0.163 The Housemaid (M) 0.170
-The Housemaid (M) 0.250 Dopamine Detox (S) 0.149 The Housemaid (M) 0.149 The Inmate (M) 0.158
-Don’t Believe (S) 0.088 The Inmate (M) 0.147 Don’t Believe (S) 0.145 Don’t Believe (S) 0.119
-Art of Letting Go (S) 0.066 The Housemaid (M) 0.145 Dopamine Detox (S) 0.117 Serpent & Wings (F) 0.110
-Serpent & Wings (F) 0.066 Art of Letting Go (S) 0.119 Art of Letting Go (S) 0.109 Art of Letting Go (S) 0.108
-Dopamine Detox (S) 0.063 Serpent & Wings (F) 0.091 Court of Ravens (F) 0.099 Court of Ravens (F) 0.104
-Court of Ravens (F) 0.058 Court of Ravens (F) 0.087 Serpent & Wings (F) 0.097 Dopamine Detox (S) 0.096
-The Ritual (M) 0.056 The Ritual (M) 0.057 Ashes & Star (F) 0.070 Ashes & Star (F) 0.072
-Ashes & Star (F) 0.029 Ashes & Star (F) 0.041 The Ritual (M) 0.050 The Ritual (M) 0.064
+The Inmate (M) 0.325 The Inmate (M) 0.153 The Inmate (M) 0.162 The Inmate (M) 0.173
+The Housemaid (M) 0.250 Don’t Believe (S) 0.149 The Housemaid (M) 0.151 The Housemaid (M) 0.169
+Don’t Believe (S) 0.088 The Housemaid (M) 0.145 Don’t Believe (S) 0.136 Don’t Believe (S) 0.115
+Art of Letting Go (S) 0.066 Dopamine Detox (S) 0.137 Dopamine Detox (S) 0.115 Court of Ravens (F) 0.107
+Serpent & Wings (F) 0.066 Art of Letting Go (S) 0.110 Art of Letting Go (S) 0.106 Serpent & Wings (F) 0.107
+Dopamine Detox (S) 0.063 Court of Ravens (F) 0.100 Court of Ravens (F) 0.102 Dopamine Detox (S) 0.101
+Court of Ravens (F) 0.058 Serpent & Wings (F) 0.092 Serpent & Wings (F) 0.100 Art of Letting Go (S) 0.095
+The Ritual (M) 0.056 The Ritual (M) 0.059 The Ritual (M) 0.064 The Ritual (M) 0.068
+Ashes & Star (F) 0.029 Ashes & Star (F) 0.056 Ashes & Star (F) 0.064 Ashes & Star (F) 0.063
 Panel C. Predicted Second-Choice Probabilities when First Choice is Ashes & Star (F)
@@
 Book Prob. Book Prob. Book Prob. Book Prob.
-Serpent & Wings (F) 0.275 The Housemaid (M) 0.155 Please Tell Me (M) 0.178 The Inmate (M) 0.184
-Court of Ravens (F) 0.243 Don’t Believe (S) 0.155 The Inmate (M) 0.157 Please Tell Me (M) 0.166
-The Inmate (M) 0.105 The Inmate (M) 0.153 The Housemaid (M) 0.134 The Housemaid (M) 0.123
-The Ritual (M) 0.082 Please Tell Me (M) 0.139 Serpent & Wings (F) 0.116 Don’t Believe (S) 0.110
-Please Tell Me (M) 0.080 Dopamine Detox (S) 0.116 The Ritual (M) 0.105 Court of Ravens (F) 0.108
-The Housemaid (M) 0.070 Art of Letting Go (S) 0.100 Don’t Believe (S) 0.093 Dopamine Detox (S) 0.092
-Don’t Believe (S) 0.052 Court of Ravens (F) 0.080 Dopamine Detox (S) 0.075 Serpent & Wings (F) 0.090
-Art of Letting Go (S) 0.048 Serpent & Wings (F) 0.062 Art of Letting Go (S) 0.075 The Ritual (M) 0.065
-Dopamine Detox (S) 0.045 The Ritual (M) 0.041 Court of Ravens (F) 0.066 Art of Letting Go (S) 0.063
+Serpent & Wings (F) 0.275 Please Tell Me (M) 0.159 Please Tell Me (M) 0.184 Please Tell Me (M) 0.176
+Court of Ravens (F) 0.243 The Inmate (M) 0.136 The Inmate (M) 0.158 The Housemaid (M) 0.152
+The Inmate (M) 0.105 Don’t Believe (S) 0.133 The Housemaid (M) 0.138 The Inmate (M) 0.150
+The Ritual (M) 0.082 The Housemaid (M) 0.129 Serpent & Wings (F) 0.123 Court of Ravens (F) 0.117
+Please Tell Me (M) 0.080 Dopamine Detox (S) 0.122 The Ritual (M) 0.097 Serpent & Wings (F) 0.104
+The Housemaid (M) 0.070 Art of Letting Go (S) 0.098 Court of Ravens (F) 0.086 Don’t Believe (S) 0.086
+Don’t Believe (S) 0.052 Court of Ravens (F) 0.089 Don’t Believe (S) 0.081 Dopamine Detox (S) 0.080
+Art of Letting Go (S) 0.048 Serpent & Wings (F) 0.082 Art of Letting Go (S) 0.071 Art of Letting Go (S) 0.072
+Dopamine Detox (S) 0.045 The Ritual (M) 0.053 Dopamine Detox (S) 0.061 The Ritual (M) 0.063
 Table 2: Predicted second-choice probabilities and their data counterparts. Letters in parentheses indicate book genres: F=Fantasy, M=Mystery, and S=Self-Help.
-3.5 Implications for Product Substitution
-So far, we have shown that our approach improves counterfactual predictions of second choices.
-One might wonder whether these $RMSE$ improvements are substantial and if they fundamentally change our understanding of product substitution in this market.
-To explore this further, in Table 1 we compare predicted second-choice probabilities with their counterparts observed in the experimental data.
+Implied Substitution Patterns
+To illustrate how our approach improves predictions of substitution patterns, in Table 2 we compare predicted second-choice probabilities with their counterparts observed in the experimental data for three of the books.
 Panel A examines substitution patterns for the self-help book Dopamine Detox.
-The data shows that the two other self-help books are, by far, its closest substitutes.
+The two other self-help books are, by far, its closest substitutes in the data.
 The plain logit model misidentifies these substitutes, incorrectly predicting that people would switch to books Please Tell Me and The Inmate—two popular books with the largest market shares.
@@
 Panel B shows a similar example with the substitutes for the mystery book Please Tell Me.
-The plain logit model mispredicts second-choice probabilities, incorrectly suggesting that, in the absence of this book, participants would mainly switch to self-help books.
-By contrast, the attribute-based mixed logit model captures strong within-genre substitution, correctly identifying the top three closest substitutes.
-Our review-based model also correctly predicts the identities of these substitutes but swaps the first two.
-However, the review-based model recognizes that The Housemaid and The Inmate have significantly higher second-choice probabilities than the third-closest substitute—an insight the attribute-based model misses.
+The plain logit model mispredicts second-choice probabilities, incorrectly suggesting that the second-closest substitute is a self-help book.
+By contrast, both the attribute-based mixed logit and our review-based logit capture strong within-genre substitution, correctly identifying the top three closest substitutes.
+Additionally, the review-based model recognizes that The Inmate and The Housemaid have significantly higher second-choice probabilities than the third-closest substitute—an insight that the attribute-based model misses.
 These examples illustrate that, beyond reducing $RMSE$ and predicting second-choice probabilities more accurately on average, our approach can learn which products are the closest substitutes.
@@
 These deviations from observed second choices suggest that there is not enough variation in the first-choice data to reliably estimate substitution patterns for some alternatives.
+Relation between in-sample and counterfactual performance
+Although we select models based on first-choice $AIC$, our experimental data allows us to verify whether this selection algorithm indeed chooses specifications with the best counterfactual performance.
+Two findings support our choice of $AIC$ as a model selection criterion.
+First, across all specifications with principal components considered by our model selection algorithm, the correlation between first-choice $AIC$ and counterfactual second-choice $RMSE$ is 0.78.
+Further, $AIC$ selects the specification that has the lowest second-choice $RMSE$ across all considered specifications.
+Using $BIC$ instead of $AIC$ selects the same specification, as does a five-fold cross-validation procedure on the first choices.
+This reassures us that the key results are robust to the choice of model selection method.
 3.4 Implications for Pricing: Merger Simulations
@@
 Therefore, predicting how a merged firm would set prices requires accurate estimates of substitution patterns.
-Figure 5: Results of hypothetical merger simulations. For each simulated merger of Dopamine Detox with one other book, the figure shows the average price increase among the two merging books.
-For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher, who sets prices by solving a joint first-order condition.
-In both cases, publishers take the prices of the other eight books as given, fixed at \$5.181818We fix the prices of the other books at \$5, the average value in the experimental dataset.
+Figure 5: Results of hypothetical merger simulations.
+For each simulated merger of Dopamine Detox with one other book (horizontal axis), the figure shows the average price increase among the two merging books.
+The horizontal dashed line represents a hypothetical policy where the decision-maker challenges all mergers expected to raise prices by at least 5%.
+Appendix Table A2 reports the exact price increase estimates used to construct this graph.
+For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher setting prices by solving a joint first-order condition.
+In both cases, publishers take the prices of the other eight books as given, fixed at \$5, and have zero marginal costs.
+[^19] We fix the prices of the other books at \$5, the average value in the experimental dataset.
 We do not optimize prices of other books because the model does not include an outside option.
@@
 We recognize that limiting the choice set to ten books makes our analysis somewhat artificial, as major publishers typically manage vast assortments with hundreds of thousands of titles.
-However, our simulations offer a clear way to evaluate how estimated substitution patterns translate into pricing decisions.
+Nevertheless, our simulations offer a natural way to evaluate how estimated substitution patterns translate into pricing decisions.
 This application is also policy-relevant given the high-profile mergers among major book publishers in the past few decades, which have drawn attention and regulatory scrutiny.
@@
 For each simulated merger, Figure 5 reports the average price increase across the two merging books.
-As discussed in Section 3.5, the plain logit model fails to capture strong within-genre substitution.
+As discussed above, the plain logit model fails to capture strong within-genre substitution.
 Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones.
@@
 Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think.
-These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 9%—approximately 2 to 2.5 times higher.
+These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 8.9%—approximately 2 to 2.5 times higher (see Appendix Table A2).
 The attribute-based mixed logit model produces predictions more aligned with the review-based model.
 In most cases, both models deviate from the plain logit in the same direction, yielding similar or even identical predicted price increases.
-However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to our selected text-based model.
+However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to the review-based model.
 These divergent predictions could lead decision-makers to different conclusions.
@@
 4 Application to Online Retail Data
-Next, we apply our approach to choice data from several online markets.
-We pursue two goals with this application.
-First, we aim to assess the external validity of our experimental results.
-Although our estimation approach works for books, this does not mean it would capture substitution patterns well in other product categories.
-By applying this approach across multiple categories, we also show that it can be applied broadly without being tailored to specific markets.
-Second, by exploring which text or image data best predict substitution patterns in various categories, we can offer practical guidance to empirical researchers on what unstructured data to collect for demand estimation.
+Next, we apply the same approach to choice data from several online markets.
+Our first goal is to show that this approach can be applied broadly across categories without being tailored.
+Our second goal is to identify which text and image data best predict substitution patterns in various categories, thus offering practical guidance on what types of unstructured data researchers should collect for demand estimation.
 4.1 Data
-Online Purchases.
-We use purchase data from the 2019-2020 Comscore Web Behavior Panel, which includes a sample of around two million U.S.
-households.
-We use the dataset constructed by Greminger et al.
+We use purchase data from the 2019-2020 Comscore Web Behavior Panel.
+Specifically, we use the dataset constructed by Greminger et al.
 (2023), who classify over 12 million unique products from Amazon.com—browsed or purchased by Comscore panelists—into narrowly defined categories.
@@
 [^21] We do not observe product rankings, so we do not include them in our demand models.
-We focus on Amazon.com due to the high volume of Amazon purchases in the Comscore dataset.
-We apply our method to 40 product categories.
-To construct this sample, we first select the 15 most-purchased products in each category and drop all other products.
-We then retain only the categories where these 15 products collectively account for at least 2,000 purchases.
-This selection criterion ensures that we observe a sufficient number of purchases to estimate both product-fixed effects and substitution patterns.
-Within each category, we treat each purchase as a choice instance and assume that consumers’ choice sets consist of the 15 selected products.
+We focus on Amazon.com due to the large volume of Amazon purchases in the Comscore dataset.
+Figure 6: Example: image and text data collected from product detail pages.
+We combine purchase data with image and text data we collected from Amazon product detail pages (Figure 6).
+We use the default product photo for image embeddings.
+For text embeddings, we use product titles, product descriptions (i.e., bullet points describing the product), and 100 most recent reviews for each product.
+In estimation, we treat text and image embeddings as fixed over time.
+While sellers could theoretically modify images and textual descriptions to reposition their products in response to unobserved demand shocks, we find that product images, titles, and descriptions very rarely change over time (see Appendix F).
+We apply our method to 40 product categories spanning clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation).
+We select the 15 most-purchased products in each category and retain only those categories where these products collectively account for at least 2,000 purchases.
+This criterion ensures that we observe enough purchases to estimate both product-fixed effects and substitution patterns.
 [^22] We omit products purchased fewer than 10 times and those with missing price data, which means some categories in our sample include fewer than 15 products.
-The chosen product categories cover a diverse range of items such as clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation).
-The first column of Table 2 lists all included categories.
-On average, each category contains 662 purchases, with an average product price of \$43.
-Figure 6: Example: image and text data collected from product detail pages.
-Images and Texts.
-We collect image and text data directly from product detail pages on Amazon.com.
-Figure 6 shows a sample product page.
-For image embeddings, we use the default product photo displayed at the top of the product detail page.
-For text embeddings, we gather information from several sections: product titles at the top, bullet points describing the product’s attributes (which we refer to as the “product description”), and the texts of the 100 most recent reviews for each product.
-4.2 Estimation and Results
-In each product category, we follow the same approach as in the experiment, extracting embeddings, reducing their dimensionality through PCA, and including them in the random coefficients logit model.
-We do not observe product attributes in these categories.
-Collecting such data is challenging because each category has its own set of choice-relevant attributes.
-Further, we would need to subjectively decide which attributes to collect in each category—a challenge that originally motivated us to look at images and texts.
-For these reasons, we focus on comparing our estimates to those from the plain logit model.
-In other words, we focus on showing that texts and images contain information about product substitution and that our approach can reliably extract this information across many categories.
-4.2.1 $AIC$ Improvements
-Table 2 summarizes the estimation results, showing the selected model and data type for each category, along with the corresponding $AIC$ improvement relative to the plain logit model.
-Figure 7 further plots the distribution of $AIC$ improvements across all categories.
-As a rule of thumb, we consider that a model has strong support compared to a simpler model if it lowers $AIC$ by at least $2.0$, and very strong support if it lowers $AIC$ by at least $5.0$.
-When a simpler model is nested within a more complex one with only one additional parameter, applying these $AIC$ thresholds is approximately equivalent to conducting a likelihood ratio test at the 5% and 1% significance levels.
-[^22] When nested models differ by one parameter, the reduction in $AIC$ is given by $\Delta AIC=2-\lambda_{LR}$, where $\lambda_{LR}$ is the likelihood ratio statistic.
-Therefore, our $\Delta AIC$ thresholds of $2.0$ and $5.0$ correspond to the likelihood ratio thresholds $4.0$ and $7.0$ (p-values $0.0455$ and $0.008$).
-As Figure 7 shows, we achieve, on average, an AIC reduction of $28.0$, with improvements reaching as high as $121.8$ in some categories.
-These results confirm that our approach performs consistently well across a wide range of categories, which span both durable and perishable goods.
-Figure 7: Fit improvements relative to plain logit across 40 product categories.
-For each category, we compute $\Delta AIC$ as the $AIC$ of the lowest-$AIC$ specification that includes texts or images minus that of the plain logit model.
-Category Data Type Model Type $\Delta$ AIC $\Delta$ HHI
-1. Clothing Active Descriptions ST -30.4 129.0%
-2. Clothing Shirts Reviews USE -19.1 116.7%
-3. Clothing Underwear Images Resnet50 -18.5 88.2%
-4. Clothing Sleep Reviews ST -18.5 93.6%
-5. Clothing Tops & Blouses Descriptions USE -18.5 94.6%
-6. Electronics Cables Titles USE -18.4 240.0%
-7. Electronics Accessories Reviews TFIDF -19.8 10.7%
-8. Electronics Keyboards Images VGG19 -20.0 88.2%
-9. Electronics Memory Cards Images VGG16 -109.7 103.5%
-10. Electronics Tablets Images Resnet50 -121.8 120.9%
-11. Electronics Monitors Reviews USE -31.6 197.3%
-12. Electronics Headphones Descriptions COUNT -26.0 1.2%
-13. Electronics Media Players Reviews TFIDF -66.9 42.7%
-14. Groceries Water Descriptions COUNT -22.5 123.0%
-15. Groceries Coffee Images VGG16 -17.2 45.6%
-16. Groceries Tea Titles ST -20.0 145.2%
-17. Groceries Chips Reviews ST -13.5 61.9%
-18. Household Aromatherapy Titles COUNT -19.7 190.7%
-19. Household Batteries Reviews TFIDF -20.1 4.8%
-20. Household Trash Bags Titles ST -24.6 93.5%
-21. Household Paper Towels Titles TFIDF -28.9 188.4%
-22. Home Sheets & Pillowcases Descriptions USE -22.3 121.2%
-23. Bedroom Beds Reviews USE -17.1 69.4%
-24. Bedroom Mattresses Titles USE -17.4 66.3%
-25. Kitchen Food Storage Images Resnet50 -27.7 133.2%
-26. Office Folders Titles USE -29.7 199.7%
-27. Office Paper Images Resnet50 -19.4 11.4%
-28. Office Markers Images VGG19 -22.3 228.7%
-29. Office Pens Titles ST -14.4 241.0%
-30. Office Printer Supplies Titles TFIDF -23.2 247.7%
-31. Pet Cat Food Titles ST -16.9 130.6%
-32. Pet Cat Litter Titles ST -21.9 126.9%
-33. Pet Cat Snacks Images Xception -19.5 142.3%
-34. Pet Dog Food Titles COUNT -19.2 65.6%
-35. Pet Dog Treats Reviews ST -22.9 163.9%
-36. Game Consoles Nintendo Titles COUNT -13.1 118.3%
-37. Video Games Nintendo Images Resnet50 -36.7 26.2%
-38. Video Games PC Images VGG19 -18.6 43.9%
-39. Video Games PS4 Titles TFIDF -32.1 159.6%
-40. Video Games Xbox Reviews USE -40.5 120.6%
-Table 2: Estimation results across 40 categories in Comscore data.
-For each category, the table shows the selected model and data type yielding the lowest $AIC$, the corresponding $AIC$ improvement relative to the plain logit model, and the relative change in the $HHI$ of the predicted diversion ratios relative to the plain logit model.
-4.2.2 Implications for Substitution
-Next, we examine how incorporating unstructured data into demand estimation affects our understanding of product substitution.
-While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how diversion ratios estimated by our approach deviate from the restrictive patterns imposed by the plain logit model.
-The assumption of independent utilities in the plain logit model implies that diversion ratios depend only on market shares and not on product similarities.
-Consequently, this model often fails to identify close substitutes, leading to overly flat diversion ratios (Conlon et al., 2023).
-By contrast, if our approach better identifies close substitutes, it should produce more unequal diversion ratios.
-To quantify this, we compute diversion ratios using the definition from Section 3 and calculate the Herfindahl-Hirschman Index ($HHI$) of predicted diversion ratios.
-[^23] Specifically, we compute the $HHI$ across $(J-1)$ conditional second choice probabilities for each first choice and then average the resulting $HHI$ measure across all $J$ possible first choices.
-We expect our approach to yield a higher $HHI$ than the plain logit model.
-The last column in Table 2 presents the results.
-[^24] In some categories, our estimates of the price coefficient $\alpha$ are positive, likely due to correlation between prices and unobserved demand shocks, even after accounting for product fixed effects.
-While instrumental variables could address this, we do not pursue this approach, as addressing price endogeneity across many product categories is orthogonal to the contribution of our paper.
-Instead, we focus on counterfactuals, such as the diversion ratios from removing a product, which remain well-defined even when the price coefficient is positive.
-Across categories, our approach more than doubles the $HHI$ of diversion ratios, with an average increase of 115%.
-Additionally, $HHI$ rises in all 40 product categories, with some increases approaching 250%.
-These findings confirm that our approach produces significantly more variable diversion ratios across the board.
-4.2.3 Relevance of Different Data Types
-Lastly, we analyze which types of unstructured data yield the largest improvement in fit in each category.
+The average product price in selected categories is \$43.
+For four electronics categories, we collect standard attributes from product detail pages to compare our approach with attribute-based mixed logit.
+Specifically, we collect 18 attributes for “Tablets,” 13 attributes for “Monitors,” 20 attributes for “Memory Cards,” and 14 attributes for “Headphones” (see Appendix E for a full list).
+4.2 Estimation Results
+We apply our approach to each product category and compare it to the plain logit model.
+In the four electronics categories for which we collected attribute data, we also compare it to the attribute-based mixed logit model.
+Since we have a relatively large number of attributes in these categories, we reduce their dimensionality via PCA and apply Algorithm 1 for model selection.
+[^23] We normalize each attribute to have mean zero and variance one before applying PCA.
+[…]
+Figure 7: Estimated diversion ratios to closest substitutes.
+This figure plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$ in each category.
+Appendix Table A4 shows results from the electronics categories with attribute data.
+In all four categories, unstructured data meaningfully improves model fit relative to observed attributes, suggesting that our approach captures information about substitution patterns beyond that contained in standard attributes.
+This finding is particularly noteworthy given that one might expect technical specifications of electronics products—captured by our extensive list of attributes—to be highly relevant for consumer choices.
+Next, we examine the estimated substitution patterns.
+While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how estimated diversion ratios deviate from those in the plain logit model.
+A well-known limitation of plain logit is that diversion ratios depend only on market shares and not on attribute similarities, so this model often fails to identify close substitutes, producing overly flat diversions (Conlon et al., 2023).
+If our approach recovers substitution better, it should produce more variable diversion ratios.
+Figure 7 plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$.
+Consistent with the intuition above, the plain logit yields relatively small diversion ratios, about 22% on average, while our approach increases this average substantially to 47%.
+In some categories, our approach estimates diversion ratios as high as 60-80%, which the plain logit does not produce.
+These results confirm that our method generates significantly more variable diversion ratios, suggesting it better identifies which products are close substitutes.
+Apart from showing greater variability, the diversion ratios predicted by our method are also more intuitive.
+For example, Appendix Tables A5 and A6 report predicted diversions for tablets.
+Because the plain logit relies on product fixed effects, it predicts that if any tablet is removed, consumers substitute to the most popular alternatives (e.g., Fire 7 or Fire HD10), producing unrealistically flat diversions.
+By contrast, the Description ST model chosen by our model selection algorithm yields more intuitive diversions: the two kids’ tablets are estimated to be close substitutes (Fire Kids and Fire HD8 Kids), the two iPads are close substitutes (iPad 9.7 and iPad 10.2), and so on.
+4.3 Relevance of Different Data Types
+Lastly, we analyze which types of unstructured data yield the largest fit improvements.
 It would be misleading to report only the best-performing model in each category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit.
-We therefore follow the statistical literature on model selection and report Akaike weights for each data type (Burnham and Anderson, 2004).
-Formally, we define $\Delta_{i}=AIC_{i}-AIC_{min}$ for each estimated specification $i$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category.
-We then compute the Akaike weight for a data type $d\in\{\text{images},\text{titles},\text{descriptions},\text{reviews}\}$ as follows:
-$w_{d}=\frac{\sum_{\{r:d(r)=d\}}\exp(-\Delta_{r}/2)}{\sum_{i}\exp(-\Delta_{i}/2)}$ (2)
-The numerator sums over specifications using data type $d$, while the denominator sums over all estimated specifications.
+We therefore follow the statistical literature on model selection and report Akaike weights (Burnham and Anderson, 2004).
+Formally, for each data type $d$ (images, titles, descriptions, or reviews) we compute the Akaike weight $w_{d}$ as
+$w_{d}=\frac{\sumop\displaylimits_{\{r:d(r)=d\}}\exp(-{}_{r}/2)}{\sumop\displaylimits_{i}\exp(-{}_{i}/2)}$ (3)
+where $i$ indexes estimated specifications, and $_{i}=AIC_{i}-AIC_{min}$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category.
+The numerator in (3) sums over estimated specifications using data type $d$, while the denominator sums over all specifications estimated during model selection.
 We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data.
@@
 These weights reflect the relative importance of data types for predicting substitution patterns in the data.
+We compute Akaike weights using the formula in (3).
 Figure 8 shows the estimated Akaike weights by category and data type.
 We find considerable variation across categories in terms of which types of data are most important for predicting substitution patterns.
-Importantly, many of these results are difficult to predict ex ante.
-For example, while we might expect visual product features to be most relevant in clothing categories, where consumers care about visual design, only one of the four clothing categories (“Clothing Underwear”) has an image weight exceeding those of texts.
-By contrast, estimating demand for “Tops & Blouses” requires product descriptions, for “Shirts” relies on customer reviews, and for “Activewear” it works best with either descriptions or reviews.
-Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in categories of Nintendo and PC video games, images contain significantly more information about substitution than text.
-These results highlight an important lesson: researchers may not be able to predict ex ante which data types will best capture substitution.
-5 Discussion
-5.1 A Practitioner’s Guide for Using Text and Images
-In this section, we summarize our key results and offer guidelines for researchers on how to incorporate unstructured data into demand estimation.
-The main takeaway from our analyses is that text and image data contain information about substitution patterns, meaning these unstructured data can be useful for demand estimation.
-At the same time, our cross-category analysis in Section 4 highlights that researchers may not be able to predict ex ante which data type will best capture substitution in a given product category.
-Therefore, we recommend that researchers collect various types of unstructured data and perform model selection as we did in Sections 3 and 4.
-Specifically, researchers can first identify the lowest-$AIC$ set of principal components for each combination of data type and embedding model, and then determine which of these selected specifications minimizes $AIC$ across all data-model combinations.
-This selection rule performed well in our experiment, where it identified the model with the best counterfactual second-choice predictions.
-In our experimental data in Section 3, we found that a text-based model produces the best counterfactual predictions, and that extending this model with observed attributes or image data did not significantly improve model fit.
-Nevertheless, in other contexts, researchers may still want to check whether incorporating observed product attributes can capture additional dimensions of product differentiation.
-This can be done by searching for a set of observed attributes that, when added to the selected model, further reduce $AIC$.
-A similar search can be conducted over additional text- and image-based principal components to determine whether including them improves model fit.
-To conduct inference based on the estimates from our approach, researchers must account for model selection to obtain valid confidence intervals.
-One potential solution is to perform sample-splitting: researchers can use one randomly selected subset of the data for model selection and the other for estimating the covariance matrix of the estimates.
-[^26] For a detailed discussion of sample-splitting, see Wasserman and Roeder (2009).
-Taylor and Tibshirani (2015) provide examples of how researchers can construct corrected p-values and confidence intervals for specific estimation problems.
-We mostly abstracted away from the issue of price endogeneity.
-Endogeneity concerns do not arise in our experiment because we randomize prices, whereas in the analysis of Amazon purchases, we rely on product fixed effects to capture product-level unobservables (e.g., unobserved quality) that may correlate with prices.
-In practice, researchers are often concerned that prices may be correlated with unobserved demand shocks that vary across markets and over time.
-To address this, researchers can combine our approach with existing methods for handling price endogeneity based on instrumental variables.
+Importantly, many of these results are difficult to predict in advance.
+For example, while we might expect visual features to be most relevant in clothing categories, where consumers care about visual design, only two of the five clothing categories (“Activewear” and “Sleep”) assign more weight to images than to text.
+In contrast, for “Tops & Blouses” and “Underwear,” product titles are most predictive of substitution, whereas for “Shirts,” reviews perform best.
+Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in the category of video games for Xbox, images contain substantially more information about substitution patterns than text.
+These results highlight that researchers cannot reliably predict in advance which data types will best capture substitution.
+Therefore, in practice, it is important to collect different data types and use model selection to identify the one that is most predictive of substitution.
+5 A Practitioner’s Guide for Using Text and Images
+In this section, we provide guidance for practitioners interested in applying our method.
+We highlight several practical lessons from Sections 3 and 4, clarify which counterfactuals our approach can and cannot accommodate, and suggest promising directions for future research.
+5.1 Data Choice and Model Selection
+The main takeaway from our results is that text and image data contain valuable information about substitution patterns, making these unstructured data useful for demand estimation.
+At the same time, our multi-category analysis in Section 4 shows that researchers may not be able to predict in advance which data type will best capture substitution in a given category.
+We therefore recommend collecting various types of unstructured data and performing model selection as in Sections 3 and 4.
+Algorithm 1 in Section 2 provides a practical heuristic for model selection: in our experiment, it successfully identified the specification that delivered the best counterfactual second-choice predictions.
+A natural question is whether researchers should add observed attributes to our approach when they are available.
+In our application, adding observed attributes to the selected review-based model did not improve fit (see Section 3.3).
+One explanation is that unstructured data may already encode the choice-relevant attributes, rendering structured attributes redundant.
+Nevertheless, in other applications, researchers may still want to test whether including observed attributes helps estimate substitution patterns.
+This can be done by checking whether adding some subset of observed attributes to the selected model with unstructured data further reduces $AIC$.
+When many attributes are available, researchers can reduce their dimensionality using PCA, as in Section 4.2.
+5.2 Counterfactual Analysis and its Boundaries
+We interpret embeddings as capturing time-invariant product aspects (e.g., attributes, functional benefits, and visual designs).
+This assumption is reasonable given that in our application, product images, titles, and descriptions rarely change over time (see Appendix F).
+Thus, our method is well-suited for a wide range of counterfactuals that require estimating consumer responses to price changes conditional on the other product attributes, including optimal pricing, simulating merger effects on prices, evaluating corrective taxes, or estimating markups (Berry and Haile, 2024).
+In such contexts, researchers can evaluate counterfactual market shares and prices just as in standard demand models, assuming that embeddings—and the resulting demand functions—do not change.
+Moreover, as usual, researchers can quantify the welfare effects of price changes by calculating the area under the demand curve (Small and Rosen, 1981).
+The fact that principal components have no direct interpretation does not preclude such analysis.
+Our approach can also be applied to study consumer responses to changes in product availability, for example, to quantify consumer surplus from new products or study how firms optimize assortments.
+In retrospective analyses, where the introduction of new products is observed in the data as in Petrin (2002), researchers can use our approach “as is” knowing that it performs well at predicting consumer responses to product removals.
+By contrast, evaluating a hypothetical new product that is not yet in the data poses an additional challenge: researchers need to obtain images and textual descriptions of this product from pre-launch information.
+In some applications, embeddings may change in counterfactual scenarios.
+For example, if customer reviews mention value relative to price (e.g., “Great quality for the price!” or “Good product, but overpriced.”), then the reviews and the resulting text embeddings will shift when prices change in counterfactuals.
+Researchers can diagnose this issue by checking how common such reviews are in their data and, if necessary, model how price changes affect embeddings.
+Another example is merger simulations: if firms redesign or reposition products after merging, this may alter embeddings and hence substitution patterns.
+If this is a first-order concern, researchers may want to specify a supply-side model in which firms choose both prices and non-price characteristics (Fan, 2013).
+That said, our method does allow for correlation between embeddings and variables that change in counterfactuals.
+For instance, embeddings may be correlated with prices if they capture unobserved product quality.
+This does not rule out counterfactuals where prices change (e.g., due to tariffs), since such changes do not alter the product’s intrinsic quality.
 5.3 Future Research Directions
-While we are not the first to use text and image data in demand estimation, our paper proposes and validates a concrete method for leveraging such data to flexibly estimate substitution patterns – a critical input for answering many empirical questions of interest.
-We do not claim that ours is the only or the best way to use unstructured data for this purpose.
-In fact, several questions remain open.
-First, other demand modeling approaches might be able to better take advantage of text and image data than the random coefficients logit model.
-In an earlier version of this paper, we tested a pairwise combinatorial logit model from Koppelman and Wen (2000), which generalizes the logit model by allowing each product pair to have its own correlation in utilities.
-We allowed these correlations to depend on the distance between products in the space of text and image embeddings.
-This model performed significantly worse than our current approach of incorporating principal components into the standard random coefficients logit model, which highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance.
-[^27] Researchers can also try relaxing the parametric assumptions of standard demand models.
-For example, with small number of products, one could incorporate principal components into a nonparametric demand model (Compiani, 2022).
-Second, our approach can be easily extended to newer, more advanced ML models that might better extract substitution patterns from texts and images.
-The approach we propose in this paper is modular in the sense that researchers can easily try alternative embeddings and test whether they improve counterfactual second-choice predictions.
-To facilitate such tests, we make our experimental dataset and estimation code publicly available.
+Our paper shows that text and image data contain valuable and easily extractable information for estimating substitution patterns. This finding opens several promising directions for future research.
+First, newer ML models might extract substitution patterns from texts and images more effectively.
+For example, Qwen3, OpenAI, and Gemini text embeddings perform robustly well across diverse natural language tasks and may thus generalize to demand estimation (Lee et al., 2025; Zhang et al., 2025).
+Because our method is modular, researchers can easily try alternative embeddings and test whether that improves performance.
+In addition, researchers can fine-tune existing models to construct embeddings optimized for counterfactual predictions.
+An open question is how much fine-tuning improves upon pre-trained models and whether these gains come without excessive computational costs.
+To facilitate future research on alternative embedding models, we make our experimental dataset and estimation code publicly available.
 [^27] Replication codes and experimental data are available in our public repository: github.com/ilyamorozov/DeepLogitReplication.
 Separately, the Python package for implementing our method is available on PyPI and at: github.com/deep-logit-demand/deeplogit.
-Third, researchers could improve upon our approach by constructing “optimal” embeddings specifically designed to improve the counterfactual predictions of interest.
-[^29] For example, Lee (2024) shows that fine-tuning LLMs helps researchers to better capture brand intercepts.
-Our current method relies on text and image models that were trained on datasets unrelated to our specific application, meaning they may not capture the most relevant dimensions of product differentiation.
-This suggests potential for improvement: researchers could fine-tune existing deep learning models or train custom ones to optimize for a specific objective, such as improving counterfactual predictions.
-It would be interesting to assess how much these tailored demand models improve upon pre-trained models and whether they can do so without significantly increasing the computational burden.
-Lastly, our approach could extend to other forms of unstructured data, such as music or videos, potentially opening doors to studying product differentiation and competition in markets where this would otherwise be challenging.
-One could also leverage tools from the explainable AI literature to construct interpretable embeddings, which would help tackle questions such as optimal product design.
-[^30] See Sisodia et al.
-(2024) for a detailed discussion of related ideas.
+Second, future research should provide a formal treatment of how to perform inference in a way that accounts for the uncertainty in the model selection step.
+One possibility would be to use sample-splitting as in Wasserman and Roeder (2009) and Taylor and Tibshirani (2015), randomly selecting a subset of the data for model selection and using the other for estimation and inference based on the selected model.
+Third, other demand models may leverage text and image data more effectively than the random coefficients logit model.
+In an earlier version of this paper, we tested the pairwise combinatorial logit model of Koppelman and Wen (2000), which allows each product pair to have its own utility correlation based on distance in the embedding space.
+The model, however, performed significantly worse than our mixed logit, not only when using unstructured data but also with observed attributes, where we let utility correlations depend on the distance between products in the attribute space.
+This highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance and suggests that combining unstructured data with more flexible demand models, such as that in Compiani (2022), may be a fruitful direction for future research.
+We note that this point applies more generally to any discrete choice model and is not specific to our method.
+Finally, because our focus is on recovering substitution patterns, we largely abstract away from price endogeneity.
+Endogeneity is not a concern in our experiment because we randomize prices.
+In the Amazon application, we rely on product fixed effects to capture unobserved attributes (e.g., quality) that may correlate with prices.
+More generally, researchers may worry that prices correlate with unobserved demand shocks that vary across markets and over time.
+To address this, one could combine our approach with existing methods for handling price endogeneity using instrumental variables (Goolsbee and Petrin, 2004; Berry et al., 2004).
+[^28] For example, one could estimate product-market fixed effects and random coefficient variances via maximum likelihood, and then estimate a two-stage least squares regression of the estimated fixed effects on prices and principal components.
+The principal components of competing products could serve as excluded instruments for prices.
 6 Conclusion
-In this paper, we demonstrated how researchers can incorporate unstructured text and image data in demand estimation to better recover substitution patterns.
-Our proposed estimation approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard discrete choice model.
-Using experimental data, we showed that our model outperforms standard attribute-based models in counterfactual predictions of second choices.
-We further validated our method with e-commerce data across dozens of categories and found that text and image data consistently help identify close substitutes within each category.
+In this paper, we demonstrate how researchers can incorporate unstructured text and image data in demand estimation to recover substitution patterns.
+Our approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard mixed logit model with random coefficients.
+Using experimental data, we show that our approach outperforms standard attribute-based models in counterfactual predictions of second choices.
+We further validate our method with e-commerce data across dozens of categories and find that text and image data consistently help identify close substitutes within each category.
 Appendix A Text Processing Steps
-Before applying our text models, listed in Section 2.1.2, we pre-process our text data as follows.
+Before applying our text models, listed in Section 2, we pre-process our text data as follows.
 Working with product titles is straightforward because each title is a short text that typically includes only 10-15 words.
@@
 We therefore apply these models directly to the unprocessed text data.
-Appendix B Choice Survey: Sanity Checks
+Appendix B Book Selection Procedure
+Which books we show to participants can significantly affect the performance of different demand models.
+Our general strategy was to select a set of books that does not bias our model comparisons in favor of any particular specification.
+To this end, we aimed to maximize variation in choice-relevant structured attributes as well as in text and image embeddings.
+We chose books from three genres—Mystery, Fantasy, and Self-Help—and identified 20 books of each genre from Amazon’s bestseller lists.
+We included books from three different genres in order to generate meaningful variation in this structured attribute, which we anticipated would be predictive of substitution patterns.
+To avoid arbitrary book selection within each genre, we implemented the following algorithm.
+We considered all possible combinations of ten books such that the three genres are roughly equally represented.
+Among these, we choose the combination of books that maximized the variance of text and image embeddings.
+[^29] Operationally, we computed the variance of image embeddings (using the VGG19 model) and the variance of text embeddings (using the USE model), then averaged the two.
+We performed a brute-force search over all possible sets of ten books and selected the set that maximized this average variance.
+The final set, shown in Figure 3, includes four mystery books, three fantasy books, and three self-help books.
+Appendix C Choice Survey: Sanity Checks
 Because choices in our experiment were hypothetical and not incentivized, we want to verify that participants made meaningful choices.
@@
 This observation suggests participants considered their genre preferences when making both choices.
-Appendix C Changes in Text and Images Over Time
-As discussed in Section 2.3, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product.
+Appendix D Computing Second-Choice Diversion Ratios
+To compute $RMSE$ in (2), we need to calculate predicted second-choice diversion ratios $\hat{s}_{j\rightarrow k}$ and their data counterparts $s_{j\rightarrow k}$.
+Recall that both $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$ reflect the probability that the consumer chooses book k in the second choice task conditional on having chosen book j in the first choice task.
+Using the analogy principle, we compute empirical diversions $s_{j\rightarrow k}$ using a simple frequency estimator:
+$s_{j\rightarrow k}=\frac{\sumop\displaylimits_{i=1}^{N}\mathbf{1}\{y_{i}^{(2)}=k,y_{i}^{(1)}=j\}}{\sumop\displaylimits_{i=1}^{N}\mathbf{1}\{y_{i}^{(1)}=j\}}\quad\text{for}\quad j\neq k,$ (4)
+where $y_{i}^{(1)}$ and $y_{i}^{(2)}$ are participant $i$’s first and second choices.
+By contrast, to compute diversions $\hat{s}_{j\rightarrow k}$ predicted by a given demand model, we use the identity
+$Pr(y_{i}^{(2)}=k|y_{i}^{(1)}=j;w_{i})=\frac{s_{k}^{(j)}(w_{i})-s_{k}(w_{i})}{s_{j}(w_{i})},$ (5)
+where $s_{j}(w_{i})$ and $s_{k}(w_{i})$ are the unconditional choice probabilities of books $j$ and $k$, $s_{k}^{(j)}(w_{i})$ is the probability of choosing book $k$ after book $j$ is removed from the choice set, and $w_{i}$ is a vector of prices $(\text{price}_{i1},\dots,\text{price}_{iJ})$ and rankings $(\text{rank}_{i1},\dots,\text{rank}_{iJ})$ that participant $i$ encounters in the experiment.
+We average diversions in (5) across all participants to compute $\hat{s}_{j\rightarrow k}$:
+$\hat{s}_{j\rightarrow k}=\frac{1}{N}\sumop\displaylimits_{i=1}^{N}\frac{s_{k}^{(j)}(w_{i})-s_{k}(w_{i})}{s_{j}(w_{i})}.$ (6)
+Appendix E Attributes Collected for Electronics Products
+For each of the four electronics categories described in Section 4.1, we collect standard attributes from Amazon product detail pages.
+Specifically, we include all specifications listed in the product descriptions and, when available, those in the “Technical Details”’ section.
+Below is the complete list of all attributes for each category:
+  1. 1.
+    Tablets: brand, model, memory, RAM, processor speed, number of cores, screen size, maximum resolution, charging time, battery life, front camera, back camera, front camera megapixels, back camera megapixels, warranty, whether the product comes with a case, and whether it includes an Amazon Kids subscription.
+  2. 2.
+    Monitors: brand, screen size, maximum resolution, refresh rate, blue light filter, frameless design, tilt adjustment, height adjustment, flicker-free display, built-in speaker, wall-mountable option, curved screen, and adaptive sync.
+  3. 3.
+    Memory Cards: brand, pack size, micro card, flash memory type, capacity, read speed, write speed, speed class, UHS speed class, device compatibility (smartphone, computer, camera, laptop, tablet), X-ray proof, temperature proof, waterproof, shockproof, magnetic proof, and whether an adapter is included.
+  4. 4.
+    Headphones: brand, color, connectivity, number of eartip sets, battery life on a single charge, battery life with charging case, deep bass feature, tangle-free design, waterproof, sweat-proof, IPX rating, whether the product comes with a case, a microphone, and a noise reduction option.
+Appendix F Changes in Text and Images Over Time
+As discussed in Section 2, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product.
 To validate this assumption, we examine whether text and images change over time.
@@
 Figure A3: Genres of participants’ first and second choices in the experiment.
+Figure A4: Share of variance of embeddings explained by principal components.
+[…]
 First Choices (Data) Second Choices (Counterf.)
 Model $\log L$ AIC RMSE Rel. to Plain Logit
-$\Delta$
 Panel A. Benchmark Model
-Plain Logit -20488.4 41006.7 0.089
+Plain Logit -20488.4 41006.7 0.091
 Panel B. Mixed Logit Models with Principal Components
-Images: InceptionV3 -20475.8 40985.6 0.080 -0.009 -9.8%
-Images: ResNet50 -20479.9 40991.8 0.080 -0.009 -9.8%
-Images: VGG19 -20481.2 40994.4 0.083 -0.006 -6.8%
-Images: Xception -20479.1 40990.3 0.086 -0.003 -3.1%
-Product Titles: Bag-of-Words Count -20479.3 40992.7 0.085 -0.004 -4.6%
-Product Titles: Bag-of-Words TF-IDF -20479.3 40992.7 0.085 -0.004 -4.6%
-Product Titles: Sentence Encoder (USE) -20479.1 40992.2 0.080 -0.009 -10.5%
-Product Titles: Sentence Transformer (ST) -20477.2 40986.5 0.082 -0.006 -7.1%
-Descriptions: Bag-of-Words Count -20480.1 40992.2 0.089 -$0$0.2%
-Descriptions: Bag-of-Words TF-IDF -20480.2 40992.3 0.089 -$0$0.2%
-Descriptions: Sentence Encoder (USE) -20476.2 40984.4 0.075 -0.014 -15.5%
-Descriptions: Sentence Transformer (ST) -20475.8 40983.6 0.073 -0.016 -18.0%
-Reviews: Bag-of-Words Count -20481.0 40995.9 0.084 -0.005 -5.2%
-Reviews: Bag-of-Words TF-IDF -20480.9 40995.7 0.084 -0.005 -5.1%
-Reviews: Sentence Encoder (USE) -20472.0 40975.9 0.069 -0.020 -22.3%
-Reviews: Sentence Transformer (ST) -20473.3 40978.6 0.072 -0.017 -18.8%
-Panel C. Attribute-Based Mixed Logit Models
-Price -20485.9 41003.9 0.088 -0.001 -0.8%
-Pages -20484.9 41001.7 0.083 -0.006 -6.8%
-Year -20484.7 41001.4 0.087 -0.002 -2.0%
-Genre -20483.3 41000.7 0.081 -0.008 -8.9%
-Price & Pages -20478.9 40991.9 0.077 -0.012 -13.0%
-Price & Year -20484.1 41002.2 0.088 -0.001 -1.4%
-Price & Genre -20483.2 41002.5 0.080 -0.008 -9.4%
-Pages & Year -20478.3 40990.7 0.077 -0.011 -12.9%
-Pages & Genre -20481.6 40999.2 0.076 -0.013 -14.9%
-Year & Genre -20483.3 41000.7 0.081 -0.008 -8.9%
-Price, Pages, & Year -20478.3 40990.7 0.077 -0.011 -12.9%
-Price, Pages, & Genre -20478.9 40991.9 0.077 -0.012 -13.0%
-Price, Year, & Genre -20482.2 41002.4 0.079 -0.010 -11.2%
-Pages, Year, & Genre -20478.3 40990.7 0.077 -0.011 -12.9%
-Price, Pages, Year, & Genre (All Attr.) -20476.9 40993.9 0.076 -0.013 -14.4%
+Images: InceptionV3 -20475.2 40990.5 0.085 -0.006 -7.0%
+Images: ResNet50 -20479.9 40997.8 0.083 -0.009 -9.4%
+Images: VGG19 -20481.2 41000.4 0.087 -0.005 -5.0%
+Images: Xception -20479.1 40996.3 0.089 -0.002 -2.1%
+Product Titles: Bag-of-Words Count -20478.7 40995.5 0.084 -0.007 -7.9%
+Product Titles: Bag-of-Words TF-IDF -20478.7 40995.5 0.084 -0.007 -7.9%
+Product Titles: Sentence Encoder (USE) -20481.2 40998.3 0.087 -0.004 -4.3%
+Product Titles: Sentence Transformer (ST) -20477.1 40992.3 0.085 -0.007 -7.3%
+Descriptions: Bag-of-Words Count -20475.5 40988.9 0.085 -0.006 -6.9%
+Descriptions: Bag-of-Words TF-IDF -20475.8 40989.6 0.085 -0.006 -6.5%
+Descriptions: Sentence Encoder (USE) -20474.2 40986.4 0.076 -0.015 -17.0%
+Descriptions: Sentence Transformer (ST) -20474.9 40987.8 0.075 -0.016 -18.1%
+Reviews: Bag-of-Words Count -20479.6 40997.1 0.082 -0.009 -9.8%
+Reviews: Bag-of-Words TF-IDF -20479.4 40996.9 0.082 -0.009 -10.1%
+Reviews: Sentence Encoder (USE) -20472.0 40981.9 0.070 -0.021 -23.0%
+Reviews: Sentence Transformer (ST) -20473.3 40984.6 0.073 -0.019 -20.4%
+Panel C. Mixed Logit with Observed Attributes
+Price -20485.9 41003.9 0.091 -0.000 -0.0%
+Pages -20484.9 41001.7 0.086 -0.005 -5.5%
+Year -20484.7 41001.4 0.091 0.000 0.1%
+Genre -20483.3 41000.7 0.084 -0.007 -7.7%
+Price & Pages -20478.9 40991.9 0.081 -0.010 -11.4%
+Price & Year -20484.1 41002.2 0.092 0.000 0.3%
+Price & Genre -20483.2 41002.5 0.086 -0.005 -5.2%
+Pages & Year -20478.3 40990.7 0.081 -0.011 -11.7%
+Pages & Genre -20481.6 40999.2 0.081 -0.010 -11.4%
+Year & Genre -20483.3 41002.7 0.084 -0.007 -7.7%
+Price, Pages, & Year -20478.3 40992.7 0.081 -0.011 -11.7%
+Price, Pages, & Genre -20478.9 40995.9 0.081 -0.010 -11.4%
+Price, Year, & Genre -20482.2 41002.4 0.084 -0.007 -7.4%
+Pages, Year, & Genre -20478.3 40994.7 0.081 -0.011 -11.7%
+Price, Pages, Year, & Genre (All Attr.) -20476.9 40993.9 0.078 -0.013 -14.1%
 Table A1: Model validation results.
-The table shows in-sample fit on first-choice data and counterfactual performance on second-choice data for all specifications considered in Figures 3 and 4 (see Section 3.2 for detailed description of these models).
+The table shows in-sample fit on first-choice data and counterfactual performance on second-choice data for all specifications considered in Figure 2 (see Section 3.2 for detailed description of these models).
+Book Plain Logit Mixed Logit with Attributes Mixed Logit with Texts
+Don’t Believe 3.7% 6.7% 8.9%
+Art of Letting Go 3.0% 4.3% 5.6%
+The Ritual 3.1% 1.6% 2.5%
+The Inmate 5.0% 4.4% 4.3%
+Ashes & Star 2.7% 1.3% 2.0%
+Court of Ravens 3.3% 3.5% 2.5%
+Serpent & Wings 3.4% 2.3% 2.3%
+Please Tell Me 6.9% 6.1% 5.6%
+The Housemaid 5.4% 5.2% 4.0%
+Table A2: Merger Simulation Results (Dopamine Detox).
+The table shows predicted relative price increases resulting from a simulated merger between Dopamine Detox and each other book respectively.
+For each simulated merger between Dopamine Detox and another book (first column), the table reports the expected average price increase for the two merging books across three estimated demand models.
+Category Data Type Model Type AIC Diversion to Closest Substitute
+1. Clothing Active Images VGG16 -25.1 11.5%
+2. Clothing Shirts Reviews USE -11.9 12.5%
+3. Clothing Underwear Titles TFIDF -16.3 45.5%
+4. Clothing Sleep Images Inceptionv3 -13.1 32.5%
+5. Clothing Tops & Blouses Titles USE -12.3 30.4%
+6. Electronics Cables Reviews TFIDF -16.7 20.1%
+7. Electronics Accessories Descriptions COUNT -19.6 19.0%
+8. Electronics Keyboards Descriptions ST -19.5 20.0%
+9. Electronics Memory Cards Images VGG16 -96.8 33.4%
+10. Electronics Tablets Descriptions ST -111.5 18.2%
+11. Electronics Monitors Images Resnet50 -22.4 19.9%
+12. Electronics Headphones Images VGG19 -10.7 6.1%
+13. Electronics Media Players Images VGG19 -90.2 23.4%
+14. Groceries Water Titles USE -19.1 19.0%
+15. Groceries Coffee Descriptions TFIDF -9.5 31.8%
+16. Groceries Tea Images Xception -17.1 37.4%
+17. Groceries Chips Descriptions ST -5.6 19.2%
+18. Household Aromatherapy Descriptions COUNT -11.2 17.6%
+19. Household Batteries Images VGG16 -10.7 -0.6%
+20. Household Trash Bags Descriptions TFIDF -16.2 40.1%
+21. Household Paper Towels Titles TFIDF -16.9 33.1%
+22. Home Sheets & Pillowcases Images Resnet50 -21.3 6.8%
+23. Bedroom Beds Descriptions USE -11.2 26.0%
+24. Bedroom Mattresses Reviews ST -12.9 23.5%
+25. Kitchen Food Storage Images VGG19 -20.5 28.9%
+26. Office Folders Descriptions ST -17.4 31.9%
+27. Office Paper Reviews ST -13.5 21.0%
+28. Office Markers Images Inceptionv3 -13.6 19.8%
+29. Office Pens Images VGG19 -13.5 31.6%
+30. Office Printer Supplies Titles USE -15.1 34.1%
+31. Pet Cat Food Descriptions COUNT -11.1 36.2%
+32. Pet Cat Litter Titles ST -12.1 34.1%
+33. Pet Cat Snacks Reviews COUNT -16.9 29.1%
+34. Pet Dog Food Images Inceptionv3 -11.7 24.5%
+35. Pet Dog Treats Titles USE -18.2 38.0%
+36. Game Consoles Nintendo Images Resnet50 -6.8 11.9%
+37. Video Games Nintendo Descriptions ST -23.5 13.3%
+38. Video Games PC Titles COUNT -9.5 20.7%
+39. Video Games PS4 Descriptions ST -34.7 32.2%
+40. Video Games Xbox Images Xception -74.7 28.9%
+Averaged -23.3 24.6%
+Table A3: Estimation results across 40 categories in Comscore data.
+For each category, the table shows the selected model and data type yielding the lowest AICAIC and the $AIC$ improvement relative to plain logit.
+The last column shows the increase in the predicted diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$, relative to plain logit (in percentage points).
+[…]
+Category: Tablets
+Mixed Logit with Attributes 4293.2 -54.9
+Mixed Logit with Unstructured Data 4236.6 -111.5
+Category: Monitors
+Mixed Logit with Attributes 1376.4 -12.3
+Mixed Logit with Unstructured Data 1366.4 -22.4
+Category: Memory Cards
+Mixed Logit with Attributes 2709.3 -72.6
+Mixed Logit with Unstructured Data 2685.1 -96.8
+Category: Headphones
+Mixed Logit with Attributes 9882.0 -4.5
+Mixed Logit with Unstructured Data 9875.8 -10.7
+Table A4: AICAIC Comparison of Our Approach and Mixed Logit with Attributes
+Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
+8 10 Kids 10.2 8 Kids Touch 9.7
+Fire 7 0.000 0.330 0.381 0.301 0.312 0.283 0.279 0.286
+Fire HD 8 0.225 0.000 0.229 0.175 0.178 0.165 0.162 0.166
+Fire HD 10 0.386 0.340 0.000 0.314 0.320 0.297 0.291 0.295
+Fire 7 Kids 0.123 0.105 0.128 0.000 0.100 0.091 0.089 0.091
+iPad 10.2 0.153 0.129 0.148 0.122 0.000 0.113 0.112 0.114
+Fire HD 8 Kids 0.043 0.038 0.049 0.035 0.035 0.000 0.032 0.033
+Dragon Touch 0.022 0.018 0.021 0.017 0.017 0.016 0.000 0.016
+iPad 9.7 0.048 0.041 0.044 0.037 0.037 0.035 0.034 0.000
+Table A5: Estimated Diversion Ratios for Tablets (Plain Logit).
+Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
+Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
+8 10 Kids 10.2 8 Kids Touch 9.7
+Fire 7 0.000 0.590 0.162 0.362 0.143 0.033 0.024 0.080
+Fire HD 8 0.753 0.000 0.429 0.336 0.200 0.193 0.108 0.116
+Fire HD 10 0.058 0.144 0.000 0.079 0.142 0.297 0.022 0.077
+Fire 7 Kids 0.127 0.131 0.080 0.000 0.012 0.440 0.293 0.007
+iPad 10.2 0.049 0.079 0.141 0.010 0.000 0.015 0.310 0.715
+Fire HD 8 Kids 0.003 0.034 0.151 0.200 0.004 0.000 0.188 0.001
+Dragon Touch 0.000 0.001 0.000 0.011 0.013 0.019 0.000 0.003
+iPad 9.7 0.011 0.020 0.036 0.002 0.485 0.003 0.055 0.000
+Table A6: Estimated Diversion Ratios for Tablets (Selected Model: Descriptions ST).
+Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.

### v4

@@
 Giovanni Compiani Ilya Morozov Stephan Seiler
-Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Jon McClure, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments.
+Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Jon McClure, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, QME Conference 2025, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments.
 We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance.
@@
 & CEPR
-We propose a demand estimation method that leverages unstructured data to infer substitution patterns.
+We propose a demand estimation approach that leverages unstructured data to infer substitution patterns.
 Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a mixed logit demand model.
-This approach enables demand estimation even when researchers lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design.
-Using a choice experiment, we show this approach outperforms standard attribute-based models at counterfactual predictions of second choices.
-We also apply it to 40 product categories on Amazon.com and consistently find that unstructured data identifies close substitutes within each category.
+This approach enables demand estimation even when researchers lack data on product attributes or when consumers value hard-to-quantify attributes such as visual design.
+Using a choice experiment, we show this approach substantially outperforms standard attribute-based models at counterfactual predictions of second choices.
+We also apply it to 40 product categories offered on Amazon.com and consistently find that unstructured data are informative about substitution patterns.
 Keywords: Demand Estimation, Unstructured Data, Deep Learning
 1 Introduction
-Many problems in economics and marketing—such as merger analysis, tariff evaluation, and optimal pricing—require researchers to estimate demand for differentiated products.
-A standard approach to these problems has been to estimate demand models that capture substitution through the similarity of product attributes.
+Many problems in economics and marketing—such as merger analysis, tariff evaluation, and optimal pricing—require estimating demand for differentiated products.
+A standard approach has been to estimate demand models that capture substitution through the similarity of product attributes.
 While common, this approach faces two practical challenges.
-First, researchers rarely observe all attributes that differentiate products.
-Instead, they rely on third-party data, where attributes are chosen based on unknown criteria, or gather their own data, subjectively selecting which attributes to collect.
-The collected attributes may not align with those most relevant to consumer choices.
-Second, consumers often consider visual design and functional benefits of products—dimensions that are difficult to capture through observed attributes.
+First, researchers rarely observe all choice-relevant attributes.
+Instead, they rely on third-party data in which attributes are chosen based on unknown criteria, or gather their own data, subjectively choosing which attributes to collect.
+Second, consumers often consider visual design and functional benefits of products, dimensions that are difficult to capture through observed attributes.
 [^1] This criticism of attribute-based models has a long history in the economics literature.
 For example, Hausman (1994, p.229) skeptically remarks that applying such models to French champagne choices would require researchers to somehow quantify the bubble content.
-In this paper, we show how researchers can incorporate text and image data in demand estimation to recover substitution patterns.
-We propose an approach that uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging.
-Using pre-trained deep learning models, we transform images and texts into vector representations—embeddings.
-We then apply Principal Component Analysis (PCA) to reduce dimensionality and capture the main dimensions of product differentiation.
-We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation.
-This approach enables researchers to incorporate hard-to-quantify product attributes, such as visual design from images and functional benefits from text, while also circumventing the need to subjectively choose which observed attributes to collect.
-Given the prevalence of unstructured data, we see our approach as a valuable addition to the toolbox of empirical researchers.
-To make it easier for others to apply our approach, we provide a publicly available Python package, DeepLogit, alongside this paper.
+In this paper, we show how researchers can incorporate unstructured data—such as product images, descriptions, or reviews—in demand estimation to recover substitution patterns.
+Using pre-trained deep learning models, we transform images and texts into vector representations (embeddings) and apply Principal Component Analysis (PCA) to reduce dimensionality and capture the main dimensions of product differentiation.
+We incorporate these principal components into a logit model with random coefficients, treating them as researchers usually treat observed product attributes.
+Lastly, we implement a model selection algorithm to choose the type of unstructured data and deep learning model that extract the strongest signal of substitution.
+This approach lets researchers incorporate hard-to-quantify product attributes, such as visual design in images and functional benefits from text, while avoiding subjective choices about which attributes to collect.
+It relies on product images and textual descriptions, which are widely available even in markets where collecting product attributes is challenging.
+Given the prevalence of unstructured data, the approach offers a valuable addition to the toolbox of empirical researchers.
+To ease adoption, we provide a Python package, DeepLogit, alongside this paper.
 [^2] The package is available on PyPI and in the public GitHub repository: github.com/deep-logit-demand/deeplogit.
@@
 (2022).
-To show that the proposed method performs well at recovering substitution patterns, we validate it extensively on several datasets.
-Validation in demand estimation is challenging because substitution patterns are usually unobserved, leaving no clear “ground truth” to evaluate model predictions.
-To overcome this challenge, we design a choice experiment where participants choose a book from a list of options.
-We randomize both prices and rankings to identify substitution patterns.
-Crucially, we elicit both first and second choices, estimate different demand models using only first choices, and evaluate their performance based on how well they predict second choices counterfactually.
+We validate this approach empirically.
+A key challenge is that substitution patterns are unobserved, so there is no ground truth to test model predictions.
+An ideal dataset would let us estimate demand precisely from consumers’ choices and then evaluate how well the model predicts an empirical measure of substitution not used in estimation.
+To achieve this, we design an experiment in which participants choose a book from a list of options, and we elicit both first and second choices.
+We randomize prices and rankings of all options.
+This variation lets us estimate substitution patterns using only first choices and without instruments, while reserving second-choice data for validation.
+We then test how well the model predicts second choices counterfactually.
 [^3] We refer to second choice predictions as “counterfactual predictions” in the sense that second choices are not used during estimation, which is based solely on first choices.
-Conditional second choices are informative about substitution because they reveal which product a consumer would substitute to if their preferred option was unavailable.
-Recording second choices thus enables us to assess how well each demand model predicts substitution patterns.
-[^4] Second choices are related to the notion of diversion ratios, which reflect the degree of substitutability and thus the intensity of price competition between two products (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2023).
-Recovering them is crucial for antitrust authorities trying to predict price increases from horizontal mergers and for managers optimizing assortments and prices.
-Using these experimental data, we show that both text and images contain useful information about substitution patterns.
-Intuitively, images capture substitution because book covers visually convey information about genre.
-Text performs even better, likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre.
-We then show that the best performing model—based on customer reviews—improves counterfactual predictions of second-choices relative to the standard mixed logit with observed attributes.
-In fact, the improvement relative to that model is on par with the improvement that the mixed logit achieves over the plain logit model without random coefficients.
-Thus, in this application, we can recover substitution from unstructured data alone, and in fact can achieve better performance than is possible even with attribute data.
-To assess the economic importance of these improvements, we simulate hypothetical mergers of book publishers.
-We find that differences in estimated substitution patterns are large enough such that an antitrust agency relying on the standard mixed logit with attributes might incorrectly approve mergers that should be challenged and block mergers that should be approved.
-Thus, our results show that incorporating text and image data yields improvements in estimated substitution patterns that are both statistically and economically significant.
-We also validate our approach on observational data.
-We apply it to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing.
-We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages.
-Here, our contribution is twofold.
-First, we show that text and images contain useful signals about substitution in all studied categories, highlighting that our approach is broadly applicable across markets.
-Second, we find—perhaps unexpectedly—that text performs best in some categories where images might be expected to be more useful, such as clothing, and vice versa.
-A practical takeaway is that researchers may benefit from collecting both text and images, even when only one seems relevant a priori.
+Because second choices reveal which product a consumer substitutes to if their preferred option is unavailable, predicting them directly validates the model’s ability to recover substitution patterns.
+Using these experimental data, we show that both text and images contain useful information about substitution.
+Intuitively, book covers visually convey information about genre, while descriptions and reviews capture plot details that identify similar books even within the same genre.
+Our model selection algorithm chooses a review-based mixed logit.
+This selected model improves counterfactual predictions of second-choices relative to a standard attribute-based mixed logit.
+In fact, the improvement relative to that model is on par with the improvement that the attribute-based mixed logit achieves over plain logit without random coefficients.
+Thus, using unstructured data alone, our approach recovers substitution patterns substantially better than standard attribute-based models.
+We further validate our approach using observational data from 40 product categories offered on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing.
+We combine online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product pages.
+We show that text and images contain useful signals about substitution in all studied categories, highlighting that our approach is broadly applicable across markets.
+We find that text performs best in some categories where images might seem ex ante more relevant (e.g., clothing), and vice versa.
+Thus, it is helpful to collect both text and image data even when only one seems relevant a priori.
 Our paper contributes to the extensive literature on demand estimation.
-Specifically, we build on papers that model substitution through the heterogeneity of preferences over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Nevo, 2001; Berry et al., 2004).
-We show that incorporating the principal components of text and image embeddings into these models captures substitution patterns well.
-A key contribution of our paper is that we validate this approach extensively using both experimental and observational data.
-This paper is also related to a vast literature in computer science that transforms unstructured data into lower-dimensional representations (Krizhevsky et al., 2012; Mikolov et al., 2013; Goodfellow et al., 2016).
-There is a growing literature on applications of pre-trained deep learning models in economics and marketing (see Battaglia et al., 2024 for a review).
-While most prior work focuses on predicting observed outcomes, our goal is fundamentally different: we aim to predict counterfactual quantities, such as substitution patterns.
-Predicting these quantities well is crucial for applications like merger simulations and optimal pricing.
-Yet, they are usually unobserved in choice data.
-As a result, we cannot apply standard cross-validation techniques and instead need to design a custom experiment that allows us to measure these otherwise unobserved quantities.
-To our knowledge, we are the first to use such an experiment for validation of demand models.
+We build on papers that model substitution through the heterogeneity of preferences over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Nevo, 2001; Berry et al., 2004).
+We show how researchers can incorporate unstructured data into these standard models to improve estimation of substitution patterns.
+Our key contribution is to validate this approach extensively using both experimental and observational data.
+We also build on a vast computer science literature that transforms unstructured data into lower-dimensional representations (Krizhevsky et al., 2012; Mikolov et al., 2013; Goodfellow et al., 2016) and a growing economics and marketing literature that applies pre-trained deep learning models (Battaglia et al., 2024; Dew, 2024).
+Unlike this prior work, which focuses on predicting observed outcomes, we aim to predict substitution patterns counterfactually—a crucial task for applications like merger simulations and optimal pricing.
+Because substitution patterns are usually unobserved, we cannot apply standard cross-validation and instead need to design an experiment to measure them.
+To our knowledge, we are the first to use such an experiment to validate demand models.
 [^4] Berry et al.
@@
 (2022) estimate demand from aggregate data on first-choice probabilities and a subset of second-choice probabilities.
-Conceptually, our validation approach reverses the logic of these papers: we only use first-choice data for estimation and treat second choices as “holdout” data for model validation.
-Additionally, Raval et al.
+Conceptually, our validation approach reverses the logic of these papers: we only use first-choice data for estimation and reserve second choices for model validation.
+In a similar spirit to our approach, Raval et al.
 (2022) use hospital closures induced by natural disasters to compare the ability of various models to predict diversions.
-Several other papers have proposed leveraging text and image data in demand estimation.
-[^6] Netzer et al.
-(2012) use data on the co-occurrence of products mentioned in online discussion forums to generate a visual representation of competing products, but they do not incorporate these measures into demand estimation.
-Dew (2024) uses image embeddings from pre-trained models to predict consumers’ ratings for products based on past ratings elicited in a survey.
+Several other papers leverage text and image data in demand estimation.
+Quan and Williams (2019) and Lee (2024) incorporate product images and textual descriptions as mean utility shifters.
+Our approach models utility covariances and can thus capture substitution patterns more flexibly.
+Han and Lee (2025) use image data to estimate substitution among fonts and study the effects of copyright protection policies.
+We contribute by validating our approach on ground truth substitution patterns, comparing counterfactual prediction performance across several embedding models and types of unstructured data, and providing practical guidance for demand estimation.
 Sisodia et al.
-(2024) extract interpretable product attributes from images and use them to estimate preferences.
-Our approach is simpler and scales better across product categories, and thus, it may be preferable in applications where interpretability of extracted attributes is not essential.
-Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities.
-Similarly, Lee (2024) employs large-language models to predict utility intercepts of new products based on their textual descriptions.
-These methods predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution.
-[^7] For example, Lee (2024) inverts a plain logit model without random coefficients to extract product fixed effects for the prediction part of his analysis.
-Thus, his model restricts substitution patterns due to the IIA property.
-In contrast, our approach models utility covariances and can thus capture flexible substitution patterns—a key input for many empirical applications of interest.
-Further, while utility intercepts can be recovered from market shares, our goal is to capture substitution patterns that are unobserved.
-We measure them in an experiment by eliciting second choices, and we use these data for validation.
-[^8] A recent paper by Han and Lee (2025) uses unstructured data to estimate substitution patterns among fonts in order to study the effects of copyright protection policies.
-A possible alternative to our use of unstructured data is to estimate substitution patterns using survey data.
+(2024) extract interpretable product attributes from images and use them to estimate preferences in a choice-based conjoint experiment.
+Because our approach is easier to implement and scales better across categories, it is preferable when researchers do not need the extracted attributes to be interpretable.
+An alternative way to learn substitution patterns is to use survey data.
 For example, Dotson et al.
@@
 Similarly, Magnolfi et al.
-(2022) elicit relative rankings from survey participants (e.g., “Product A is closer to B than to C”), generate embeddings from these rankings, and use them in a random coefficients logit model.
+(2022) elicit relative rankings from survey participants (e.g., “Product A is closer to B than to C”), generate embeddings from these rankings, and use them in a mixed logit model.
 Our approach complements these survey-based methods but has the advantage of using only widely available text and image data, thereby avoiding the need for costly category-specific surveys.
-Another complementary approach is proposed by McClure (2025), who shows how publicly available recommendations data from online platforms can also be used to estimate substitution patterns.
-Our approach is well-suited for several empirical applications.
-Researchers can use it to address a wide range of economics and marketing questions that require accurate and flexible estimates of product substitution.
-This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through altering assortments and prices.
-Additionally, researchers can apply this approach to study how multi-product firms optimize prices and promotions (Hoch et al., 1995; Vilcassim and Chintagunta, 1995).
-Finally, because our approach circumvents the need to collect distinct attributes for each category, it may be valuable for researchers trying to estimate demand across many categories (Döpper et al., 2024).
+Our approach is well-suited for empirical applications that require accurate and flexible estimates of product substitution.
+This includes evaluating how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through changes in assortments and prices.
+It can also be used to study how multi-product firms should optimize prices and promotions (Hoch et al., 1995; Vilcassim and Chintagunta, 1995).
+The approach avoids the need to collect category-specific attributes, making it easier to estimate demand across many categories (Döpper et al., 2024).
 [^5] Döpper et al.
 (2024) remark that estimating demand across multiple categories “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach can address this problem if researchers have access to unstructured data.
+Finally, this approach might be particularly useful for estimating demand for creative goods—such as books, art, and movies—where unstructured data is likely to capture product differentiation more fully than observed attributes alone.
 2 Proposed Approach
@@
 Step 1. Extracting Embeddings from Texts and Images
-To extract information from product images, we extract their low-dimensional representations—embeddings—using pre-trained deep learning models.
+To extract information from product images, we compute low-dimensional representations—embeddings—using pre-trained deep learning models.
 We use four convolutional neural networks: VGG19 (Simonyan and Zisserman, 2015), ResNet50 (He et al., 2016), InceptionV3 (Szegedy et al., 2016), and Xception (Chollet, 2017).
@@
 However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings.
-The key advantage of using pre-trained models is that it reduces the computational burden of estimation while leveraging models trained on large-scale datasets, known for their strong performance in image classification tasks (Russakovsky et al., 2015).
+The key advantage of using pre-trained models is that it reduces the computational burden of estimation while leveraging models that were trained on large-scale datasets and are known for their strong performance in image classification tasks (Russakovsky et al., 2015).
 Because these models perform well at distinguishing visually similar objects, we expect embeddings to capture the key visual features that differentiate products.
 Rather than committing to a single model, we perform model selection to determine which one best captures substitution patterns, as described below.
-We also extract information from texts, including product titles, descriptions, and customer reviews.
+We also extract information from texts, including product titles, descriptions, and reviews.
 We use a bag-of-words model, which represents text as fixed-length vectors based on word counts, as well as a variation with a TF-IDF vectorizer.
@@
 This makes it more likely that text embeddings capture distinctive words that differentiate products from one another.
-While these count models are relatively simple, we view them as useful benchmarks because they can still detect attributes mentioned in titles, descriptions, and reviews.
+While these count models are relatively simple, we view them as useful benchmarks because they can still detect attributes mentioned in textual product descriptions.
 In addition, we employ two pre-trained deep learning models: Universal Sentence Encoder (USE) (Cer et al., 2018), and BERT Sentence Transformer (ST) (Reimers and Gurevych, 2019).
@@
 Step 2. Generating Principal Components
-Although embeddings compress texts and images into lower-dimensional representations, they remain high-dimensional compared to the number of attributes typically used in demand models.
+Although embeddings compress texts and images into lower-dimensional representations, they remain high-dimensional compared to the number of attributes typically included in demand models.
 For example, incorporating 512-dimensional VGG19 embeddings into a random coefficients logit model would require prohibitively costly numerical integration over all dimensions.
-We therefore apply PCA to further reduce dimensionality (Backus et al., 2021).
-The dimensionality reduction via PCA is attractive for multiple reasons.
-Raw embeddings are trained on general-purpose datasets to classify images into broad categories, such as “tablet” or “laptop.” However, to estimate demand, we need to analyze consumer choices within a given category.
-PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution within a category.
-Additionally, principal components are appealing because they are orthogonal to each other, which avoids multicollinearity issues and simplifies estimation of random coefficients in demand models.
+We apply PCA to further reduce dimensionality (Backus et al., 2021).
+PCA is appealing for several reasons.
+Because embeddings are trained on general-purpose datasets to classify images into broad categories, they contain variation that distinguishes each category from one another (e.g., laptops from tablets).
+By applying PCA within each category, we downweight this common variation and focus on differences across products within that category.
+Additionally, the resulting principal components are orthogonal to each other, which avoids multicollinearity issues and thus simplifies estimation of logit demand models with random coefficients.
 Step 3. Including Principal Components into a Demand Model
@@
 We thus propose a model selection algorithm that relies solely on first choices.
-We perform model selection using in-sample $AIC$.
-An advantage of using $AIC$ is that it can be interpreted as a bias-corrected estimator of the expected relative Kullback-Leibler (KL) information based on the maximized log-likelihood function (Akaike, 1998).
-Further, in our experimental results, in-sample $AIC$ strongly correlates with counterfactual performance, justifying its use for model selection (see Section 3.3 where we also implement $BIC$ and cross-validation).
-Our model selection algorithm considers a pool of candidate variables consisting of price and the first $P$ principal components.
-In our application, we choose $P=6$, which collectively account for 70-80% of variance in embeddings (see Appendix Figure A4).
-We first estimate all possible models with a random coefficient on a single variable and retain the best-fitting model.
-We then estimate all possible models with two random coefficients and assess whether fit improves relative to the best single-variable model.
-We continue to increase the number of variables $K$ that have random coefficients until there is no further $AIC$ improvement.
-We repeat this across all specifications—defined as a combination of a text or image model and data type (e.g., USE Reviews)—and choose the one with the lowest overall $AIC$.
-The full procedure is given in Algorithm 1.
+We perform model selection using in-sample $AIC$, which balances model fit and complexity and can be interpreted as approximating expected out-of-sample predictive performance (Akaike, 1998).
+In our experimental results, in-sample $AIC$ strongly correlates with counterfactual performance, justifying its use for model selection (see Section 3.3 for a discussion and comparison to $BIC$ and cross-validation).
+Algorithm 1 describes our model selection procedure.
+We define the candidate set as the variables for which random coefficients may be included.
+In our applications, this set includes price and the first $P$ principal components.
+We first estimate the plain logit model without random coefficients ($K=0$) and record its $AIC$.
+For each $K\geq 1$, we estimate all models with random coefficients on $K$ candidate variables and retain the model with the lowest $AIC$.
+If this $AIC$ is lower than that of the previously selected model, we update the selected specification and increase $K$; otherwise, we stop and select the previous model.
+We increase $K$ until $AIC$ no longer improves.
+We repeat this procedure for each specification—defined by data type and embedding model (e.g., USE Reviews)—and choose the specification with the lowest overall $AIC$.
 This algorithm avoids costly combinatorial search over all possible specifications.
-At the same time, it ensures that we do not rely too heavily on the ordering of principal components, which reflects the explained variance of embeddings but need not be predictive of preferences.
-For example, the approach might select a model that has random coefficients on the second and fourth principal components, but not the first or third.
+At the same time, it ensures that we do not rely too heavily on the default ordering of principal components, which reflects the explained variance of embeddings but need not match how predictive they are of preferences.
+For example, substitution patterns may be driven by the second and fourth principal components.
+In this case, our approach may select a model with random coefficients on those components, but not on the first or third.
       Algorithm 1: Model Selection
@@
      on subsets of $K$ variables from the candidate set $\mathcal{R}=\{\text{price, P$C_{1}^{(m)}$, $\dots$, P$C_{P}^{(m)}$}\}$
-9:     Find subset $R_{K}^{*}\in\mathcal{R}$ that minimizes $AIC$ at $AIC_{K}^{*}$
+9:     Find subset $R_{K}^{*}\subseteq\mathcal{R}$ that minimizes $AIC$ at $AIC_{K}^{*}$
 10:     if $AIC_{K}^{*}<\text{BestAIC}^{(m)}$ do
-11:       $\text{BestSpec}^{(m)}\leftarrow R_{K}^{*}$ update the best specification
-12:       $\text{BestAIC}^{(m)}\leftarrow AIC_{K}^{*}$ update the lowest AICAIC
-13:   end while
-14: end for
-15: Choose the best-fitting specification $m^{*}=\arg\min_{m}\text{BestAIC}^{(m)}$
+11:       $\text{BestSpec}^{(m)}\leftarrow R_{K}^{*}$ : update the best specification
+12:       $\text{BestAIC}^{(m)}\leftarrow AIC_{K}^{*}$ : update the lowest AICAIC
+13:     end if
+14:   end while
+15: end for
+16: Choose the best-fitting specification $m^{*}=\arg\min_{m}\text{BestAIC}^{(m)}$
 Why Does This Approach Work?
-There are several reasons why texts and images may predict substitution.
-Consumers may choose based on how a product looks and how it is described by sellers or other consumers.
-Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect attributes that drive substitution.
+The key feature of our approach is that it uses text and images as proxies for product attributes that shape substitution.
 Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly.
 Similarly, photos may showcase design features like color and casing style.
-Our approach extracts this information and uses it to capture substitution.
-Given this interpretation, embeddings correlate with substitution patterns because they capture choice-relevant attributes, but the relationship need not be causal: changing product images or descriptions will not necessarily alter substitution patterns.
-Consequently, we cannot study questions like optimal product design or positioning unless we further unpack the link between embeddings and substitution, which is outside the scope of this paper.
-By contrast, our method is well-suited for counterfactuals where embeddings can be held fixed, such as optimal pricing or merger simulations.
-We provide a more detailed discussion of the counterfactuals enabled by our approach in Section 5.2.
+Our approach extracts this information in an automated way and uses it to estimate substitution patterns.
+We interpret the relationship between embeddings and substitution as correlational.
+In particular, our approach does not assume that consumers directly consider product images or textual descriptions when making choices.
+Under this interpretation, our approach is well-suited for counterfactuals where embeddings can be held fixed, such as optimal pricing or merger simulations.
+By contrast, it is less appropriate for settings in which counterfactual changes that may alter product design or positioning and, in turn, the embeddings themselves.
+In such cases, researchers are required to estimate the causal effect of embeddings on demand.
+Section 5.2 provides additional discussion.
 Advantages Over Standard Methods
@@
 This is valuable for researchers seeking to estimate demand and study competition and pricing across many product categories.
+When some data on observed attributes are available, researchers can incorporate them into our approach.
+For example, they can expand the set of candidate variables in Algorithm 1 to include the observed attributes themselves or, when the number of attributes is large, their principal components.
+In our experiment, adding observed attributes does not improve counterfactual predictions, indicating that the information in the attributes is subsumed by the unstructured data.
+See Sections 3.3 and 5.1 for details.
 3 Validation with Experimental Data
-We apply our method to data from a choice experiment to show it recovers substitution patterns more effectively than standard attribute-based methods.
-A key challenge is that we usually do not directly observe substitution patterns in the data, and thus do not have a “ground truth” to validate demand models.
-To address this, we design and implement a choice experiment that measures second-choice diversion ratios, which directly reflect substitution.
+Next, we validate our approach empirically.
+A key challenge is that substitution patterns are unobserved, so there is typically no ground truth against which to test model predictions.
+To address this challenge, we design a choice experiment that provides such a ground-truth measure of substitution.
 3.1 Experiment Design
@@
 We limit the number of options to ten to ensure participants pay attention to most options, which allows us to abstract away from limited consideration.
-After a participant selects a book from the list of ten options, we remove this book from the list.
+After a participant selects a book, we remove it from the list.
 The participant then proceeds to a second choice task, where they choose among the remaining nine books.
@@
 To give all models a fair chance at capturing substitution, we display attributes (author, year, genre, pages), cover images, and texts (plot description and five reviews), collected from Amazon product pages.
+[^10] We collect the first five reviews displayed on each product detail page at the time of data collection.
 We take all books from Amazon’s bestseller lists, sampling them in a way that does not bias our model comparisons in favor of any particular specification (see Appendix B for details).
-Lastly, we randomize books’ rankings and prices across participants, keeping them fixed across choice tasks for the same participant.
+Lastly, we randomize book rankings and prices across participants, keeping them fixed across choice tasks for the same participant.
 [^11] Prices were drawn from a discrete uniform distribution ranging from \$3 to \$7.
-This double randomization generates clean variation for estimating substitution patterns (Berry and Haile, 2024).
-[^15] Although all participants technically see the same set of ten books in the first choice task, they rarely choose books that are displayed near the bottom of the list.
-Therefore, by randomizing the order in which books are displayed to participants, we vary their effective choice sets.
+This variation allows us to cleanly estimate substitution patterns from only first choices and reserve second-choice data for validation (Berry and Haile, 2024).
 The book category is well-suited for our analysis because participants are likely to consider both structured attributes (e.g., genre) and unstructured information (e.g., plot descriptions).
-It is also unclear a priori whether substitution patterns are better predicted by images or texts—while texts describe rich plot details, cover images contain clear cues of genres, and whether the book is fiction or non-fiction (see Section 3.3 for a detailed discussion).
+It is also unclear a priori whether substitution patterns are better predicted by images or texts—while texts describe rich plot details, cover images contain clear cues of genres, and whether the book is fiction or non-fiction (see Section 3.3 for further discussion).
 Figure 1: Example of a choice task in our experiment. The screenshot displays the top portion of the page as it appeared to participants.
@@
 3.2 Model Specifications and Estimation Results
-We compare our approach against two benchmark models: the plain logit model and a standard logit model with random coefficients model on attributes.
+We compare our approach against two benchmark models: the plain logit model and a mixed logit model with attributes.
 All three models fall into a framework where the indirect utility of participant $i$ from book $j$ is
@@
 [^13] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns using observed characteristics or principal components.
-This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does.
+This variation gives us an additional exogenous shifter that helps us identify substitution patterns, analogous to price variation.
 We estimate three demand models:
@@
     Plain Logit.
-We include neither attributes $x_{j}$ nor principal components $PC_{j}$, estimating a constant price coefficient $\alpha_{i}=\alpha$, rank coefficient $\gamma$ and fixed effects $\delta_{j}$.
+We include neither attributes $x_{j}$ nor principal components $PC_{j}$, estimating only a constant price coefficient $\alpha_{i}=\alpha$, rank coefficient $\gamma$, and fixed effects $\delta_{j}$.
 Table 1 compares in-sample $AIC$.
-The best mixed logit with attributes includes random coefficients on the number of pages and year, reducing $AIC$ by $16.0$ relative to plain logit.
+The best mixed logit specification with attributes includes random coefficients on the number of pages and year, reducing $AIC$ by $16.0$ relative to plain logit.
 The best model with unstructured data, Reviews USE, puts random coefficients on the first two principal components.
@@
 3.3 Validation Using Second-Choice Data
-To evaluate demand models, we let them confront a difficult counterfactual prediction problem—predicting second choices, i.e.
-which books participants switch to when their most preferred option becomes unavailable.
-A model can only predict second choices well if it has accurately captured the true substitution patterns.
-Predicting second choices is also directly relevant in antitrust where second-choice diversion ratios are often used to measure substitutability and the intensity of price competition (Conlon and Mortimer, 2021).
-Operationally, we first estimate each model via Maximum Likelihood Estimation (MLE) using first choices only.
+We assess how well different demand models predict second choices counterfactually.
+Because second choices reveal which books participants choose when their preferred option is unavailable, predicting them accurately requires a model that captures substitution patterns well.
+Further, predicting second choices is directly relevant in antitrust where second-choice diversion ratios are often used as measures of substitutability and price competition (Conlon and Mortimer, 2021).
+We first estimate each model via Maximum Likelihood Estimation (MLE) using only first choices.
 We then use the estimated model to predict counterfactual second-choice diversion ratios, $\hat{s}_{j\rightarrow k}$, defined as the probability that the participant chooses product $k$ in the second choice task conditional on having chosen book $j$ in the first choice task.
@@
 $RMSE=\sqrt{\frac{1}{J(J-1)}\sumop\displaylimits_{j}\sumop\displaylimits_{k\neq j}\left(s_{j\rightarrow k}-\hat{s}_{j\rightarrow k}\right)^{2}}$ (2)
-Since we do not use second-choice data in estimation, lower $RMSE$ indicates that the model performs better at counterfactual predictions of second choices.
+Since we do not use second-choice data in estimation, lower $RMSE$ indicates better performance at counterfactual predictions.
 Appendix D provides further details on how we compute diversions $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$.
 Figure 2: Comparison of models in terms of counterfactual RMSERMSE on second choices.
-The two benchmarks in the first panel are the plain logit without random coefficients and the best-fitting mixed logit with random coefficients on observed attributes.
+The two benchmarks in the first panel are the plain logit without random coefficients and the lowest-$AIC$ mixed logit with random coefficients on observed attributes.
 The remaining specifications correspond to mixed logit models with random coefficients on principal components extracted from image or text embeddings.
@@
 Images
-All four image models outperform the plain logit model, with the lowest-$AIC$ model, InceptionV3, reducing $RMSE$ by 7.0% relative to plain logit.
+All four image models outperform the plain logit model, with the selected (i.e., lowest-$AIC$) model, InceptionV3, reducing $RMSE$ by 7.0% relative to plain logit.
 To understand why images predict substitution, consider the book covers used in our study (see Figure 3).
@@
 All text models outperform the plain logit, and their relative performance highlights several notable patterns.
-First, performance tends to improve when we move from simple bag-of-words models to the more advanced USE and ST models.
-This is particularly true for descriptions and reviews, which contain detailed information about book plots but do not always describe them in the same words or phrases.
+First, performance improves as we move from simple bag-of-words models to the more advanced USE and ST models.
+This is particularly true for descriptions and reviews, which contain detailed information about book plots but do not always use the same words or phrases.
 Therefore, extracting substitution patterns from text requires natural language models that can accurately measure semantic similarity.
 Second, performance improves as we include richer text data.
-For instance, the USE model reduces $RMSE$ by 4.3% with book titles, 17% with descriptions, and 23% with reviews relative to the plain logit.
+For instance, relative to the plain logit, the USE model reduces $RMSE$ by 4.3% using titles, 17% using descriptions, and 23% using reviews.
 This makes intuitive sense.
-Although book titles are brief, they often contain subtle genre cues.
-For example, mystery titles frequently hint at secrets (Please Tell Me) or characters in confined situations (Housemaid, Inmate), whereas fantasy titles often include words that evoke a gothic atmosphere (Serpent, Wings, Ravens, Ruin).
-Descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals.
-Reviews go even further: consumers summarize plots while also expressing opinions.
-When reviewing mystery novels, for instance, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings.
-Thus, it is unsurprising that reviews provide more information about substitution than descriptions, which, in turn, are more informative than titles.
+Take mystery books, for example.
+Titles may contain subtle genre cues, hinting at characters in confined situations (Housemaid, Inmate).
+Descriptions reveal further plot details, such as hidden identities, past relationships, and betrayals.
+Lastly, customer reviews contain the richest information about a book’s style and pacing: for instance, readers may praise cliffhangers and intriguing twists, or critique pacing issues like slow starts or rushed endings.
 Extension: Combining Data Types
-If text, images, and attributes provides distinct signals of substitution, combining them might improve fit and counterfactual performance.
+If text, images, and attributes provide distinct signals of substitution, combining them might improve fit and counterfactual performance.
 To explore this, we start from the selected Review USE model and attempt to extend it in two ways: (1) by adding a combination of observed product attributes with random coefficients, or (2) by adding a combination of principal components from InceptionV3, the lowest-$AIC$ image model.
-Neither extension improves $AIC$ or $RMSE$.
-This result suggests information across data types is highly correlated, with text providing the strongest signal of substitution.
+In each case, we evaluate all subsets of added variables and estimate random coefficients for these added variables, in addition to those already included in the selected Review USE model.
+[^16] We chose this approach because including all principal components from unstructured data and all observed attributes in the candidate set for Algorithm 1 would be too computationally burdensome.
+We find that in all evaluated specifications, the variances of the added random coefficients are zero.
+As a result, $AIC$ increases and second-choice $RMSE$ does not improve.
+This result suggests information is highly correlated across data types, with text providing the strongest signal of substitution.
 What do Principal Components Capture?
-To better understand the variation captured by embeddings, in Figure 4 we show book locations in the space of the two principal components selected into our best-fitting Review USE model.
+To better understand the variation captured by embeddings, in Figure 4 we show book locations in the space of the two principal components selected into our best-performing Review USE model.
 These principal components align with intuitive substitution patterns: the first one (horizontal axis) separates non-fiction on the left from fiction on the right, while the second (vertical axis) further distinguishes science fiction from mystery.
@@
 These deviations from observed second choices suggest that there is not enough variation in the first-choice data to reliably estimate substitution patterns for some alternatives.
+Finally, to illustrate how estimated substitution patterns affect counterfactuals, we perform a stylized merger simulation in Appendix E.
+We find that the selected review-based model predicts substantially different post-merger prices than benchmark models.
 Relation Between In-Sample and Counterfactual Performance
@@
 First, across all specifications with principal components considered by our model selection algorithm, the correlation between first-choice $AIC$ and counterfactual second-choice $RMSE$ is 0.78.
-Further, $AIC$ selects the specification that has the lowest second-choice $RMSE$ across all considered specifications.
-Using $BIC$ instead of $AIC$ selects the same specification, as does a five-fold cross-validation procedure on the first choices.
+Second, $AIC$ selects the specification that has the lowest second-choice $RMSE$ across all considered specifications.
+Lastly, using $BIC$ instead of $AIC$ selects the same specification, as does a five-fold cross-validation procedure on the first choices.
 This reassures us that the key results are robust to the choice of model selection method.
-3.4 Implications for Pricing: Merger Simulations
-To illustrate how estimated substitution patterns can influence counterfactuals of interest, we conduct simulations of horizontal mergers—a natural application of our approach.
-Antitrust agencies routinely use demand models to assess whether a hypothetical merger would lead to a significant price increase (Federal Trade Commission, 2022).
-If the merging firms’ products are close substitutes, the merger creates strong upward pricing pressure, as the firm can “recapture” some consumers after raising prices.
-Therefore, predicting how a merged firm would set prices requires accurate estimates of substitution patterns.
-Figure 5: Results of hypothetical merger simulations.
-For each simulated merger of Dopamine Detox with one other book (horizontal axis), the figure shows the average price increase among the two merging books.
-The horizontal dashed line represents a hypothetical policy where the decision-maker challenges all mergers expected to raise prices by at least 5%.
-Appendix Table A2 reports the exact price increase estimates used to construct this graph.
-For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher setting prices by solving a joint first-order condition.
-In both cases, publishers take the prices of the other eight books as given, fixed at \$5, and have zero marginal costs.
-[^19] We fix the prices of the other books at \$5, the average value in the experimental dataset.
-We do not optimize prices of other books because the model does not include an outside option.
-Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
-We recognize that limiting the choice set to ten books makes our analysis somewhat artificial, as major publishers typically manage vast assortments with hundreds of thousands of titles.
-Nevertheless, our simulations offer a natural way to evaluate how estimated substitution patterns translate into pricing decisions.
-This application is also policy-relevant given the high-profile mergers among major book publishers in the past few decades, which have drawn attention and regulatory scrutiny.
-[^20] In 2022, a federal judge blocked the proposed merger between Penguin Random House and Simon & Schuster, citing concerns that it would stifle competition (U.S.
-Department of Justice, 2022).
-This merger would have reduced the “Big Five” publishers to the “Big Four” and significantly increased the concentration of the publishing market.
-Prior to that, Penguin Books and Random House merged in 2013 to form Penguin Random House—the largest trade book publisher globally.
-Figure 5 illustrates the predicted price increase resulting from the joint ownership of Dopamine Detox with each of the other books.
-We select Dopamine Detox as an example where our approach outperforms alternative models in capturing substitution patterns (see Table 2).
-For each simulated merger, Figure 5 reports the average price increase across the two merging books.
-As discussed above, the plain logit model fails to capture strong within-genre substitution.
-Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones.
-This misclassification affects the predicted price changes in merger simulations.
-Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think.
-These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 8.9%—approximately 2 to 2.5 times higher (see Appendix Table A2).
-The attribute-based mixed logit model produces predictions more aligned with the review-based model.
-In most cases, both models deviate from the plain logit in the same direction, yielding similar or even identical predicted price increases.
-However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to the review-based model.
-These divergent predictions could lead decision-makers to different conclusions.
-As an example, consider an antitrust agency that applies a heuristic rule, challenging all mergers expected to increase prices by more than 5% (Bhattacharya et al., 2023).
-Compared to decisions made using the review-based model, which best captures substitution patterns, a decision-maker relying on the plain logit model would approve two mergers that should be challenged (Don’t Believe Everything You Think, The Art of Letting Go) and challenge one that should be approved (The Housemaid).
-Similarly, a decision-maker using the attribute-based mixed logit model would approve a merger that should be challenged (The Art of Letting Go) and challenge one that should be approved (The Housemaid).
 4 Application to Online Retail Data
-Next, we apply the same approach to choice data from several online markets.
+Next, we apply our approach to choice data from several online markets.
 Our first goal is to show that this approach can be applied broadly across categories without being tailored.
-Our second goal is to identify which text and image data best predict substitution patterns in various categories, thus offering practical guidance on what types of unstructured data researchers should collect for demand estimation.
+Our second goal is to determine which types of unstructured data best predict substitution patterns in various categories and to offer practical guidance on what data researchers should collect for demand estimation.
 4.1 Data
@@
 In estimation, we treat text and image embeddings as fixed over time.
-While sellers could theoretically modify images and textual descriptions to reposition their products in response to unobserved demand shocks, we find that product images, titles, and descriptions very rarely change over time (see Appendix F).
+While sellers could theoretically modify images and textual descriptions to reposition their products, we find that product images, titles, and descriptions very rarely change over time (see Appendix G).
 We apply our method to 40 product categories spanning clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation).
 We select the 15 most-purchased products in each category and retain only those categories where these products collectively account for at least 2,000 purchases.
-This criterion ensures that we observe enough purchases to estimate both product-fixed effects and substitution patterns.
-[^22] We omit products purchased fewer than 10 times and those with missing price data, which means some categories in our sample include fewer than 15 products.
-The average product price in selected categories is \$43.
-For four electronics categories, we collect standard attributes from product detail pages to compare our approach with attribute-based mixed logit.
-Specifically, we collect 18 attributes for “Tablets,” 13 attributes for “Monitors,” 20 attributes for “Memory Cards,” and 14 attributes for “Headphones” (see Appendix E for a full list).
+These selection criteria ensure that we can estimate product fixed effects precisely, which helps capture time-invariant differences across products, and that we have sufficient variation to estimate substitution patterns.
+For four electronics categories, we collect standard attributes from product detail pages to compare our approach with attribute-based mixed logit: 18 attributes for “Tablets,” 13 for “Monitors,” 20 for “Memory Cards,” and 14 for “Headphones” (see Appendix F for a full list).
 4.2 Estimation Results
-We apply our approach to each product category and compare it to the plain logit model.
-In the four electronics categories for which we collected attribute data, we also compare it to the attribute-based mixed logit model.
+We first compare our approach to benchmark models in the four electronics categories for which we collected attribute data.
 Since we have a relatively large number of attributes in these categories, we reduce their dimensionality via PCA and apply Algorithm 1 for model selection.
-[^23] We normalize each attribute to have mean zero and variance one before applying PCA.
+[^18] We normalize each attribute to have mean zero and variance one before applying PCA to prevent attributes with larger variance or units from dominating the principal components.
+As in Section 3, we let the candidate set include price and the first $P=6$ principal components.
+Table 3 shows the results.
+In all four categories, unstructured data meaningfully improves model fit relative to both plain logit and attribute-based mixed logit, suggesting that our approach captures information about substitution patterns beyond that contained in standard attributes.
+This finding is particularly noteworthy given that one might expect technical specifications of electronics products—captured by our extensive list of attributes—to be highly relevant for consumer choices.
+[…]
+Category: Tablets
+Mixed Logit with Attributes 4293.2 -54.9
+Mixed Logit with Unstructured Data 4236.6 -111.5
+Category: Monitors
+Mixed Logit with Attributes 1376.4 -12.3
+Mixed Logit with Unstructured Data 1366.4 -22.4
+Category: Memory Cards
+Mixed Logit with Attributes 2709.3 -72.6
+Mixed Logit with Unstructured Data 2685.1 -96.8
+Category: Headphones
+Mixed Logit with Attributes 9882.0 -4.5
+Mixed Logit with Unstructured Data 9875.8 -10.7
+Table 3: AICAIC comparison of our approach and attribute-based mixed logit. The table reports the AICAIC of the specification selected by Algorithm 1 within each class of demand models.
+These performance differences translate into different substitution patterns.
+For example, consider diversion ratios for the category of tablets (see Appendix Tables A2-A4).
+Relative to the plain logit, the selected Description ST model produces more intuitive substitution patterns: kids’ tablets are close substitutes (Fire Kids and Fire HD8 Kids), iPads are close substitutes (iPad 9.7 and iPad 10.2), and so on.
+The predicted diversions in this model also differ markedly from those in the mixed logit with attributes.
+However, in the absence of second choice data, we cannot directly test how well different models predict diversions counterfactually, as we do in Section 3.
+We therefore rely on the fit comparisons reported in Table 3, which favor the Description ST model.
 […]
@@
 This figure plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$ in each category.
-Appendix Table A4 shows results from the electronics categories with attribute data.
-In all four categories, unstructured data meaningfully improves model fit relative to observed attributes, suggesting that our approach captures information about substitution patterns beyond that contained in standard attributes.
-This finding is particularly noteworthy given that one might expect technical specifications of electronics products—captured by our extensive list of attributes—to be highly relevant for consumer choices.
-Next, we examine the estimated substitution patterns.
-While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how estimated diversion ratios deviate from those in the plain logit model.
-A well-known limitation of plain logit is that diversion ratios depend only on market shares and not on attribute similarities, so this model often fails to identify close substitutes, producing overly flat diversions (Conlon et al., 2023).
-If our approach recovers substitution better, it should produce more variable diversion ratios.
-Figure 7 plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$.
-Consistent with the intuition above, the plain logit yields relatively small diversion ratios, about 22% on average, while our approach increases this average substantially to 47%.
-In some categories, our approach estimates diversion ratios as high as 60-80%, which the plain logit does not produce.
-These results confirm that our method generates significantly more variable diversion ratios, suggesting it better identifies which products are close substitutes.
-Apart from showing greater variability, the diversion ratios predicted by our method are also more intuitive.
-For example, Appendix Tables A5 and A6 report predicted diversions for tablets.
-Because the plain logit relies on product fixed effects, it predicts that if any tablet is removed, consumers substitute to the most popular alternatives (e.g., Fire 7 or Fire HD10), producing unrealistically flat diversions.
-By contrast, the Description ST model chosen by our model selection algorithm yields more intuitive diversions: the two kids’ tablets are estimated to be close substitutes (Fire Kids and Fire HD8 Kids), the two iPads are close substitutes (iPad 9.7 and iPad 10.2), and so on.
+A well-known limitation of the plain logit is that diversion ratios depend only on market shares rather than attribute similarity, leading to overly flat diversions (Conlon et al., 2023).
+Figure 6 plots the estimated diversion ratio to the closest substitute, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$.
+The plain logit yields relatively small diversions, about 22% on average, while our approach increases this average to 47% and, in some categories, to as high as 60-80%.
+These results indicate that our approach produces substantially more variable diversion ratios, suggesting it better identifies close substitutes.
 4.3 Relevance of Different Data Types
 Lastly, we analyze which types of unstructured data yield the largest fit improvements.
-It would be misleading to report only the best-performing model in each category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit.
+It would be misleading to report only the best-performing model in each product category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit.
 We therefore follow the statistical literature on model selection and report Akaike weights (Burnham and Anderson, 2004).
-Formally, for each data type $d$ (images, titles, descriptions, or reviews) we compute the Akaike weight $w_{d}$ as
-$w_{d}=\frac{\sumop\displaylimits_{\{r:d(r)=d\}}\exp(-{}_{r}/2)}{\sumop\displaylimits_{i}\exp(-{}_{i}/2)}$ (3)
-where $i$ indexes estimated specifications, and $_{i}=AIC_{i}-AIC_{min}$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category.
-The numerator in (3) sums over estimated specifications using data type $d$, while the denominator sums over all specifications estimated during model selection.
+Formally, let $d$ denote data type (images, titles, descriptions, or reviews), and let $AIC_{d}$ denote the lowest $AIC$ among all specifications that use data type dd.
+[^21] The summation in the denominator is over all four data types.
+For each data type $d$, we compute the Akaike weight $w_{d}$ as
+$w_{d}=\frac{\exp(-{}_{d}/2)}{\sumop\displaylimits_{k}\exp(-{}_{k}/2)}$ (3)
+where $_{d}=AIC_{d}-AIC_{min}$ and $AIC_{min}=\min_{d}AIC_{d}$ is the lowest $AIC$ across all four data types.
 We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data.
@@
 Figure 7: Akaike weights of different types of unstructured data across 40 product categories.
-These weights reflect the relative importance of data types for predicting substitution patterns in the data.
+These weights reflect the relative performance of data types in capturing substitution patterns from choice data.
 We compute Akaike weights using the formula in (3).
-Figure 8 shows the estimated Akaike weights by category and data type.
-We find considerable variation across categories in terms of which types of data are most important for predicting substitution patterns.
-Importantly, many of these results are difficult to predict in advance.
-For example, while we might expect visual features to be most relevant in clothing categories, where consumers care about visual design, only two of the five clothing categories (“Activewear” and “Sleep”) assign more weight to images than to text.
-In contrast, for “Tops & Blouses” and “Underwear,” product titles are most predictive of substitution, whereas for “Shirts,” reviews perform best.
-Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in the category of video games for Xbox, images contain substantially more information about substitution patterns than text.
-These results highlight that researchers cannot reliably predict in advance which data types will best capture substitution.
-Therefore, in practice, it is important to collect different data types and use model selection to identify the one that is most predictive of substitution.
+Figure 7 shows the estimated Akaike weights by category and data type, revealing substantial variation across categories.
+Many results are difficult to anticipate a priori.
+For example, although one might expect visual features to be most relevant for clothing, only two of the five clothing categories (“Activewear” and “Sleep”) place more weight on images than text.
+In “Tops & Blouses” and “Underwear,” product titles are most predictive of substitution, while in “Shirts,” reviews perform best.
+Similarly, we may expect descriptions and reviews of video games to be most informative, analogous to books.
+In practice, we find that images contain substantially more information about substitution in the “Video Games Xbox” category.
+These results highlight that researchers cannot reliably predict which data types best capture substitution.
+In practice, it is therefore important to collect multiple data types and use model selection to choose the one that best captures substitution.
 5 A Practitioner’s Guide for Using Text and Images
@@
 We therefore recommend collecting various types of unstructured data and performing model selection as in Sections 3 and 4.
-Algorithm 1 in Section 2 provides a practical heuristic for model selection: in our experiment, it successfully identified the specification that delivered the best counterfactual second-choice predictions.
+Algorithm 1 in Section 2 provides a practical heuristic for model selection: in our experiment, it successfully identifies the specification that produces the best counterfactual second-choice predictions.
 A natural question is whether researchers should add observed attributes to our approach when they are available.
-In our application, adding observed attributes to the selected review-based model did not improve fit (see Section 3.3).
+In our application, adding observed attributes to the selected review-based model does not improve fit (see Section 3.3).
 One explanation is that unstructured data may already encode the choice-relevant attributes, rendering structured attributes redundant.
 Nevertheless, in other applications, researchers may still want to test whether including observed attributes helps estimate substitution patterns.
-This can be done by checking whether adding some subset of observed attributes to the selected model with unstructured data further reduces $AIC$.
-When many attributes are available, researchers can reduce their dimensionality using PCA, as in Section 4.2.
+For example, researchers can expand the set of candidate variables in Algorithm 1 to include the observed attributes themselves or, when many attributes are available, their principal components.
 5.2 Counterfactual Analysis and its Boundaries
-We interpret embeddings as capturing time-invariant product aspects (e.g., attributes, functional benefits, and visual designs).
-This assumption is reasonable given that in our application, product images, titles, and descriptions rarely change over time (see Appendix F).
-Thus, our method is well-suited for a wide range of counterfactuals that require estimating consumer responses to price changes conditional on the other product attributes, including optimal pricing, simulating merger effects on prices, evaluating corrective taxes, or estimating markups (Berry and Haile, 2024).
-In such contexts, researchers can evaluate counterfactual market shares and prices just as in standard demand models, assuming that embeddings—and the resulting demand functions—do not change.
-Moreover, as usual, researchers can quantify the welfare effects of price changes by calculating the area under the demand curve (Small and Rosen, 1981).
-The fact that principal components have no direct interpretation does not preclude such analysis.
-Our approach can also be applied to study consumer responses to changes in product availability, for example, to quantify consumer surplus from new products or study how firms optimize assortments.
-In retrospective analyses, where the introduction of new products is observed in the data as in Petrin (2002), researchers can use our approach “as is” knowing that it performs well at predicting consumer responses to product removals.
-By contrast, evaluating a hypothetical new product that is not yet in the data poses an additional challenge: researchers need to obtain images and textual descriptions of this product from pre-launch information.
-In some applications, embeddings may change in counterfactual scenarios.
-For example, if customer reviews mention value relative to price (e.g., “Great quality for the price!” or “Good product, but overpriced.”), then the reviews and the resulting text embeddings will shift when prices change in counterfactuals.
-Researchers can diagnose this issue by checking how common such reviews are in their data and, if necessary, model how price changes affect embeddings.
-Another example is merger simulations: if firms redesign or reposition products after merging, this may alter embeddings and hence substitution patterns.
-If this is a first-order concern, researchers may want to specify a supply-side model in which firms choose both prices and non-price characteristics (Fan, 2013).
-That said, our method does allow for correlation between embeddings and variables that change in counterfactuals.
-For instance, embeddings may be correlated with prices if they capture unobserved product quality.
-This does not rule out counterfactuals where prices change (e.g., due to tariffs), since such changes do not alter the product’s intrinsic quality.
+Our approach is well suited for applications that assess consumer responses to price changes while holding embeddings fixed, including optimal pricing, merger simulations, corrective taxes, and markup estimation (Berry and Haile, 2024).
+In these applications, embeddings may be correlated with prices in the observed data, but they, and hence the demand system, should be held fixed in counterfactual simulations.
+Under this assumption, researchers can compute counterfactual market shares, prices, and welfare using standard tools from demand analysis.
+For instance, consumer welfare can be computed as the area under the demand curve (Small and Rosen, 1981).
+The fact that embeddings have no direct interpretation does not preclude these analyses.
+Our approach can also predict responses to changes in product availability, for example, to study firms’ assortment choices or consumer surplus from new products (Petrin, 2002).
+If new products launches are observed in the data, researchers can apply our method directly, since our results show it performs well at predicting consumer responses to product removals.
+On the other hand, when evaluating a hypothetical new product, researchers must construct its embeddings from pre-launch information, such as images and descriptions from marketing materials or early retailer listings.
+By contrast, our approach is less suitable in applications where embeddings may change in counterfactuals.
+For example, if customer reviews mention value relative to price (e.g., “Great quality for the price!” or “Good product, but overpriced.”), embeddings constructed from them may shift when prices change.
+Researchers can diagnose this issue by checking how frequently price-related language appears in reviews and, if necessary, modeling how price changes affect embeddings.
+Another example is simulations of mergers that may induce firms to redesign products, which would alter embeddings and hence substitution patterns.
+Studying such mergers is not possible with our approach unless researchers impose additional assumptions on firm conduct and on how embeddings map into attributes that firms can control.
 5.3 Future Research Directions
@@
 Separately, the Python package for implementing our method is available on PyPI and at: github.com/deep-logit-demand/deeplogit.
-Second, future research should provide a formal treatment of how to perform inference in a way that accounts for the uncertainty in the model selection step.
-One possibility would be to use sample-splitting as in Wasserman and Roeder (2009) and Taylor and Tibshirani (2015), randomly selecting a subset of the data for model selection and using the other for estimation and inference based on the selected model.
-Third, other demand models may leverage text and image data more effectively than the random coefficients logit model.
-In an earlier version of this paper, we tested the pairwise combinatorial logit model of Koppelman and Wen (2000), which allows each product pair to have its own utility correlation based on distance in the embedding space.
-The model, however, performed significantly worse than our mixed logit, not only when using unstructured data but also with observed attributes, where we let utility correlations depend on the distance between products in the attribute space.
-This highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance and suggests that combining unstructured data with more flexible demand models, such as that in Compiani (2022), may be a fruitful direction for future research.
-We note that this point applies more generally to any discrete choice model and is not specific to our method.
+Second, our method incorporates embeddings into the demand model as if they were the true attributes driving substitution, when in practice embeddings may be only imperfect proxies of these attributes.
+Our results show that, despite this limitation, embeddings are sufficiently informative to improve counterfactual predictions.
+Nevertheless, future research should explore whether explicitly accounting for error in these proxies can further improve performance.
+For an early contribution in this area, see Christensen and Compiani (2026).
+Third, other demand models may leverage unstructured data more effectively.
+In an earlier version of this paper, we tested the pairwise combinatorial logit model of Koppelman and Wen (2000), which allows pairwise utility correlations to depend on distances in the embedding space.
+That model performed worse than mixed logit, not only when using unstructured data but also when using distances in observed attributes, suggesting that functional form assumptions can significantly impact counterfactual performance.
+A fruitful direction for future research might be to combine unstructured data with more flexible demand models, such as that in Compiani (2022).
 Finally, because our focus is on recovering substitution patterns, we largely abstract away from price endogeneity.
 Endogeneity is not a concern in our experiment because we randomize prices.
-In the Amazon application, we rely on product fixed effects to capture unobserved attributes (e.g., quality) that may correlate with prices.
-More generally, researchers may worry that prices correlate with unobserved demand shocks that vary across markets and over time.
+In the e-commerce application, we rely on product fixed effects to capture unobserved attributes (e.g., quality) that may correlate with prices.
+More generally, prices may correlate with unobserved demand shocks that vary across markets and over time.
 To address this, one could combine our approach with existing methods for handling price endogeneity using instrumental variables (Goolsbee and Petrin, 2004; Berry et al., 2004).
@@
 6 Conclusion
-In this paper, we demonstrate how researchers can incorporate unstructured text and image data in demand estimation to recover substitution patterns.
+In this paper, we show how researchers can incorporate unstructured text and image data in demand estimation to recover substitution patterns.
 Our approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard mixed logit model with random coefficients.
@@
 Appendix A Text Processing Steps
-Before applying our text models, listed in Section 2, we pre-process our text data as follows.
-Working with product titles is straightforward because each title is a short text that typically includes only 10-15 words.
-When assessing similarity based on product descriptions, we merge the text from all bullet points, and apply our models to the merged text.
-For customer reviews, we transform the text of each review into a separate vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
-For both bag-of-words approaches, we further pre-process text data by removing stopwords and lemmatizing words.
-We remove stopwords using the standard dictionary of common English words in the NLTK package.
+Before computing embeddings, we pre-process text as follows.
+We keep product titles unchanged because they typically consist of only 10-15 words.
+For product descriptions, we merge the text from all bullet points and apply the embedding models to the merged text.
+For customer reviews, we transform the text of each review into a vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
+For both bag-of-words models, we further pre-process text data by removing stopwords and lemmatizing words.
+Stopwords are removed using the standard dictionary of common English words in the NLTK package.
 We lemmatize words using the WordNet Lemmatizer from the same package, NLTK.
-Then, we convert each pre-processed text into a vector of word occurrences (weighted word occurrences for TF-IDF).
-The other two models, USE and ST, have a built-in text pre-processing step.
-We therefore apply these models directly to the unprocessed text data.
+We then convert each processed text into a vector of word occurrences (or weighted word occurrences for TF-IDF).
+By contrast, the USE and ST models have built-in text pre-processing, so we apply them directly to the raw text.
 Appendix B Book Selection Procedure
-Which books we show to participants can significantly affect the performance of different demand models.
-Our general strategy was to select a set of books that does not bias our model comparisons in favor of any particular specification.
-To this end, we aimed to maximize variation in choice-relevant structured attributes as well as in text and image embeddings.
+The set of books shown to participants can significantly affect the performance of demand models.
+Our goal was to select books in a way that does not favor any particular specification.
+To this end, we maximized variation in both observed attributes as well as in text- and image-based embeddings.
 We chose books from three genres—Mystery, Fantasy, and Self-Help—and identified 20 books of each genre from Amazon’s bestseller lists.
@@
 Among these, we choose the combination of books that maximized the variance of text and image embeddings.
-[^29] Operationally, we computed the variance of image embeddings (using the VGG19 model) and the variance of text embeddings (using the USE model), then averaged the two.
+[^25] We computed the variance of image embeddings (using the VGG19 model) and the variance of text embeddings (using the USE model), then averaged the two.
 We performed a brute-force search over all possible sets of ten books and selected the set that maximized this average variance.
@@
 Appendix C Choice Survey: Sanity Checks
-Because choices in our experiment were hypothetical and not incentivized, we want to verify that participants made meaningful choices.
+Because choices in our experiment were not incentivized, it is crucial to verify that participants made meaningful choices.
 Several summary statistics suggest that participants took the choice tasks seriously.
-First, only 50 of the initially recruited participants (less than 1%) completed the entire study in less than one minute and thus had to be dropped from our sample.
+First, only 50 of the initially recruited participants (less than 1%) completed the entire study in under one minute and thus were excluded from the sample.
 The remaining participants spent, on average, 7 minutes on the survey overall and 1.3 minutes on the choice tasks, indicating they took their time to make careful selections and did not mindlessly click through the survey (Figure A2).
-Second, in the choice tasks, participants were disproportionately more likely to select books of the genre they reported to be their favorite in a questionnaire before the choice tasks (Figure A2), suggesting they considered the book attributes when making their choices.
-Finally, participants’ choices were consistent across the two choice tasks (Figure A3).
+Second, participants were disproportionately more likely to select books of the genre they reported to be their favorite in a questionnaire before the choice tasks (Figure A3).
+Further, their choices were consistent across the two choice tasks (Figure A4).
 For example, over 60% of participants who selected a mystery book in the first task chose another mystery book in the second choice task.
-This observation suggests participants considered their genre preferences when making both choices.
+These observations suggest that participants considered their genre preferences when making choices.
 Appendix D Computing Second-Choice Diversion Ratios
@@
 where $y_{i}^{(1)}$ and $y_{i}^{(2)}$ are participant $i$’s first and second choices.
-By contrast, to compute diversions $\hat{s}_{j\rightarrow k}$ predicted by a given demand model, we use the identity
+By contrast, to compute diversions $\hat{s}_{j\rightarrow k}$ predicted by a given demand model, we use the following identity from Conlon and Mortimer (2021):
 $Pr(y_{i}^{(2)}=k|y_{i}^{(1)}=j;w_{i})=\frac{s_{k}^{(j)}(w_{i})-s_{k}(w_{i})}{s_{j}(w_{i})},$ (5)
@@
 $\hat{s}_{j\rightarrow k}=\frac{1}{N}\sumop\displaylimits_{i=1}^{N}\frac{s_{k}^{(j)}(w_{i})-s_{k}(w_{i})}{s_{j}(w_{i})}.$ (6)
-Appendix E Attributes Collected for Electronics Products
+Appendix E Implications for Pricing: Merger Simulations
+To illustrate how estimated substitution patterns can influence counterfactuals of interest, we conduct a stylized simulation of horizontal mergers—a natural application of our approach.
+Antitrust agencies routinely use demand models to assess whether a hypothetical merger would lead to a significant price increase (Federal Trade Commission, 2022).
+If the merging firms’ products are close substitutes, the merger creates strong upward pricing pressure, as the firm can “recapture” some consumers after raising prices.
+Therefore, predicting how a merged firm would set prices requires accurate estimates of substitution patterns.
+Figure A1: Results of hypothetical merger simulations.
+For each simulated merger of Dopamine Detox with one other book (horizontal axis), the figure shows the predicted average price increase among the two merging books.
+The horizontal dashed line represents a hypothetical policy where the decision-maker challenges all mergers expected to raise prices by at least 5%.
+We perform simulations using our demand estimates from Section 3.
+For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher setting prices by solving a joint first-order condition.
+In both cases, publishers take the prices of the other eight books as given, fixed at \$5, and have zero marginal costs.
+[^26] We fix the prices of the other books at \$5, the average value in the experimental dataset.
+We do not optimize prices of other books because the model does not include an outside option.
+Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
+We recognize that restricting the choice set to ten books and excluding an outside option makes the setting somewhat artificial.
+Still, we view it as a useful way to evaluate how estimated substitution patterns translate into counterfactuals of interest.
+Figure A1 illustrates the predicted price increase resulting from the joint ownership of Dopamine Detox with each of the other books.
+We select Dopamine Detox as an example where our approach outperforms alternative models in capturing substitution patterns (see Table 2).
+For each simulated merger, Figure A1 reports the average price increase across the two merging books.
+As discussed in Section 3, the plain logit model fails to capture strong within-genre substitution.
+Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones.
+This misclassification affects the predicted price changes in merger simulations.
+Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think.
+These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 8.9%—approximately 2 to 2.5 times higher.
+The attribute-based mixed logit model produces predictions more aligned with the review-based model.
+In most cases, both models deviate from the plain logit in the same direction, yielding similar or even identical predicted price increases.
+However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to the review-based model.
+These divergent predictions could lead decision-makers to different conclusions.
+As an example, consider an antitrust agency that applies a heuristic rule, challenging all mergers expected to increase prices by more than 5% (Bhattacharya et al., 2023).
+Compared to decisions made using the review-based model, which best captures substitution patterns, a decision-maker relying on the plain logit model would approve two mergers that should be challenged (Don’t Believe Everything You Think, The Art of Letting Go) and challenge one that should be approved (The Housemaid).
+Similarly, a decision-maker using the attribute-based mixed logit model would approve a merger that should be challenged (The Art of Letting Go) and challenge one that should be approved (The Housemaid).
+Appendix F Attributes Collected for Electronics Products
 For each of the four electronics categories described in Section 4.1, we collect standard attributes from Amazon product detail pages.
-Specifically, we include all specifications listed in the product descriptions and, when available, those in the “Technical Details”’ section.
+Specifically, we include all specifications listed in the product descriptions and, when available, those in the “Technical Details” section of the product detail page.
 Below is the complete list of all attributes for each category:
@@
     Headphones: brand, color, connectivity, number of eartip sets, battery life on a single charge, battery life with charging case, deep bass feature, tangle-free design, waterproof, sweat-proof, IPX rating, whether the product comes with a case, a microphone, and a noise reduction option.
-Appendix F Changes in Text and Images Over Time
-As discussed in Section 2, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product.
+Appendix G Changes in Text and Images Over Time
+Throughout the paper, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product.
 To validate this assumption, we examine whether text and images change over time.
@@
 While we do not have review data to verify this claim, future research should examine whether review embeddings are stable in online markets.
-The selected categories are: “Shirts,” “Coffee,” “Aromatherapy,” “Mattresses,” “Markers,” “Pet Litters,” “Nintendo Games,” “Tablets,” “Memory Cards,” “Monitors,” and “Earbuds.” In each category, we collect data for the products used in our estimation in Section 4 that were not discontinued, totaling 136 products.
+The selected categories are: “Shirts,” “Coffee,” “Aromatherapy,” “Mattresses,” “Markers,” “Pet Litters,” “Nintendo Games,” “Tablets,” “Memory Cards,” “Monitors,” and “Earbuds.” In each category, we collect data for the products used in estimation in Section 4 that were not discontinued, totaling 136 products.
 We find that product images do not change over time.
@@
 Similarly, descriptions change for just 21 products (15%).
-Most changes do not affect which product features are revealed, but they alter which ones are immediately visible versus being revealed only after clicking on “See more product details.” Thus, we conclude that changes in text and images are not a significant concern for the products in our empirical application.
+Most changes do not affect which product features are revealed, but they alter which ones are immediately visible versus being revealed only after clicking on “See more product details.” Thus, we conclude that changes in text and images are not a significant concern in our empirical application.
 Figure A2: Time spent by participants on choice tasks in the experiment.
@@
 The table shows in-sample fit on first-choice data and counterfactual performance on second-choice data for all specifications considered in Figure 2 (see Section 3.2 for detailed description of these models).
-Book Plain Logit Mixed Logit with Attributes Mixed Logit with Texts
-Don’t Believe 3.7% 6.7% 8.9%
-Art of Letting Go 3.0% 4.3% 5.6%
-The Ritual 3.1% 1.6% 2.5%
-The Inmate 5.0% 4.4% 4.3%
-Ashes & Star 2.7% 1.3% 2.0%
-Court of Ravens 3.3% 3.5% 2.5%
-Serpent & Wings 3.4% 2.3% 2.3%
-Please Tell Me 6.9% 6.1% 5.6%
-The Housemaid 5.4% 5.2% 4.0%
-Table A2: Merger Simulation Results (Dopamine Detox).
-The table shows predicted relative price increases resulting from a simulated merger between Dopamine Detox and each other book respectively.
-For each simulated merger between Dopamine Detox and another book (first column), the table reports the expected average price increase for the two merging books across three estimated demand models.
+Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
+8 10 Kids 10.2 8 Kids Touch 9.7
+Fire 7 0.000 0.330 0.381 0.301 0.312 0.283 0.279 0.286
+Fire HD 8 0.225 0.000 0.229 0.175 0.178 0.165 0.162 0.166
+Fire HD 10 0.386 0.340 0.000 0.314 0.320 0.297 0.291 0.295
+Fire 7 Kids 0.123 0.105 0.128 0.000 0.100 0.091 0.089 0.091
+iPad 10.2 0.153 0.129 0.148 0.122 0.000 0.113 0.112 0.114
+Fire HD 8 Kids 0.043 0.038 0.049 0.035 0.035 0.000 0.032 0.033
+Dragon Touch 0.022 0.018 0.021 0.017 0.017 0.016 0.000 0.016
+iPad 9.7 0.048 0.041 0.044 0.037 0.037 0.035 0.034 0.000
+Table A2: Estimated diversion ratios for tablets (plain logit).
+Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
+Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
+8 10 Kids 10.2 8 Kids Touch 9.7
+Fire 7 0.000 0.939 0.451 0.144 0.328 0.019 0.347 0.073
+Fire HD 8 0.592 0.000 0.027 0.023 0.068 0.002 0.267 0.008
+Fire HD 10 0.247 0.023 0.000 0.167 0.213 0.328 0.299 0.332
+Fire 7 Kids 0.035 0.006 0.090 0.000 0.055 0.624 0.000 0.016
+iPad 10.2 0.100 0.025 0.147 0.072 0.000 0.018 0.058 0.558
+Fire HD 8 Kids 0.004 0.000 0.174 0.582 0.009 0.000 0.000 0.011
+Dragon Touch 0.003 0.004 0.006 0.000 0.001 0.000 0.000 0.002
+iPad 9.7 0.018 0.002 0.105 0.012 0.326 0.009 0.029 0.000
+Table A3: Estimated diversion ratios for tablets (mixed logit with observed attributes).
+Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
+Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
+8 10 Kids 10.2 8 Kids Touch 9.7
+Fire 7 0.000 0.590 0.162 0.362 0.143 0.033 0.024 0.080
+Fire HD 8 0.753 0.000 0.429 0.336 0.200 0.193 0.108 0.116
+Fire HD 10 0.058 0.144 0.000 0.079 0.142 0.297 0.022 0.077
+Fire 7 Kids 0.127 0.131 0.080 0.000 0.012 0.440 0.293 0.007
+iPad 10.2 0.049 0.079 0.141 0.010 0.000 0.015 0.310 0.715
+Fire HD 8 Kids 0.003 0.034 0.151 0.200 0.004 0.000 0.188 0.001
+Dragon Touch 0.000 0.001 0.000 0.011 0.013 0.019 0.000 0.003
+iPad 9.7 0.011 0.020 0.036 0.002 0.485 0.003 0.055 0.000
+Table A4: Estimated diversion ratios for tablets (selected model: Descriptions ST).
+Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
 Category Data Type Model Type AIC Diversion to Closest Substitute
@@
 The last column shows the increase in the predicted diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$, relative to plain logit (in percentage points).
-[…]
-Category: Tablets
-Mixed Logit with Attributes 4293.2 -54.9
-Mixed Logit with Unstructured Data 4236.6 -111.5
-Category: Monitors
-Mixed Logit with Attributes 1376.4 -12.3
-Mixed Logit with Unstructured Data 1366.4 -22.4
-Category: Memory Cards
-Mixed Logit with Attributes 2709.3 -72.6
-Mixed Logit with Unstructured Data 2685.1 -96.8
-Category: Headphones
-Mixed Logit with Attributes 9882.0 -4.5
-Mixed Logit with Unstructured Data 9875.8 -10.7
-Table A4: AICAIC Comparison of Our Approach and Mixed Logit with Attributes
-Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
-8 10 Kids 10.2 8 Kids Touch 9.7
-Fire 7 0.000 0.330 0.381 0.301 0.312 0.283 0.279 0.286
-Fire HD 8 0.225 0.000 0.229 0.175 0.178 0.165 0.162 0.166
-Fire HD 10 0.386 0.340 0.000 0.314 0.320 0.297 0.291 0.295
-Fire 7 Kids 0.123 0.105 0.128 0.000 0.100 0.091 0.089 0.091
-iPad 10.2 0.153 0.129 0.148 0.122 0.000 0.113 0.112 0.114
-Fire HD 8 Kids 0.043 0.038 0.049 0.035 0.035 0.000 0.032 0.033
-Dragon Touch 0.022 0.018 0.021 0.017 0.017 0.016 0.000 0.016
-iPad 9.7 0.048 0.041 0.044 0.037 0.037 0.035 0.034 0.000
-Table A5: Estimated Diversion Ratios for Tablets (Plain Logit).
-Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
-Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
-8 10 Kids 10.2 8 Kids Touch 9.7
-Fire 7 0.000 0.590 0.162 0.362 0.143 0.033 0.024 0.080
-Fire HD 8 0.753 0.000 0.429 0.336 0.200 0.193 0.108 0.116
-Fire HD 10 0.058 0.144 0.000 0.079 0.142 0.297 0.022 0.077
-Fire 7 Kids 0.127 0.131 0.080 0.000 0.012 0.440 0.293 0.007
-iPad 10.2 0.049 0.079 0.141 0.010 0.000 0.015 0.310 0.715
-Fire HD 8 Kids 0.003 0.034 0.151 0.200 0.004 0.000 0.188 0.001
-Dragon Touch 0.000 0.001 0.000 0.011 0.013 0.019 0.000 0.003
-iPad 9.7 0.011 0.020 0.036 0.002 0.485 0.003 0.055 0.000
-Table A6: Estimated Diversion Ratios for Tablets (Selected Model: Descriptions ST).
-Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.